### v1

# Gröbner Basis Cryptanalysis of Ciminion and Hydra

Abstract
Ciminion and Hydra are two recently introduced symmetric key Pseudo-Random Functions for Multi-Party Computation applications. For efficiency both primitives utilize quadratic permutations at round level. Therefore, polynomial system solving-based attacks pose a serious threat to these primitives. For Ciminion we construct a quadratic degree reverse lexicographic (DRL) Gröbner basis for the iterated polynomial model via affine transformations. For Hydra we provide a computer-aided proof in SageMath that a quadratic DRL Gröbner basis is already contained within the iterated polynomial system for the Hydra heads after affine transformations and a linear change of coordinates.
Our Ciminion DRL Gröbner basis simplifies cryptanalysis, since one does not need to impose genericity assumptions, like being regular or semi-regular, anymore to derive complexity estimates on key recovery attacks.
In the Hydra proposal it was claimed that $r_{\mathcal{H}}=31$ rounds for the heads are sufficient to achieve $128$ bits of security against Gröbner basis attacks for key recovery. However, for $r_{\mathcal{H}}=31$ standard term order conversion to a lexicographic (LEX) Gröbner basis for our Hydra DRL Gröbner basis requires just $126$ bits. Moreover, via the Eigenvalue Method up to $r_{\mathcal{H}}=33$ rounds can be attacked below $128$ bits.
keywords:
Gröbner basis Eigenvalue Method Ciminion Hydra
1 Introduction
With secure Multi-Party Computation (MPC) several parties can jointly compute a function without revealing their private inputs. In practice, data often has to be securely transferred from and to third parties before it can be used in MPC protocols. Also, one may need to store intermediate results securely in a database. For efficiency, it is recommended to perform encryption via MPC-friendly Pseudo-Random Functions (PRF) [GRR+16] which utilize a secret-shared symmetric key. Moreover, one can utilize MPC-friendly PRFs to speed up MPC applications, e.g. [IKOS07, LTW13, CDG+17]. Compared to classical bit-based PRFs their MPC counterparts have two novel design criteria:
  • They must be native over large finite fields $\mathbb{F}_{q}$, typically $\log_{2}\left(q\right)\geq 64$ and $q$ is prime.
  • They must admit a low number of multiplications for evaluation.
Two such recently introduced PRFs are Ciminion [DGGK21] and Hydra [GØSW23]. These constructions are based on the iteration of degree $2$ Feistel and Lai–Massey permutations. Due to the low degree at round level, both designs consider Gröbner basis attacks [ACG+19] as one of the most threatening attack vectors.
Let us shortly review a Gröbner basis attack on an iterated design.
  1. (1)
    Model the function with an iterated system of polynomials
    $\mathcal{R}(\mathbf{x}_{i-1},\mathbf{y})-\mathbf{x}_{i}=\mathbf{0},$
    where $\mathcal{R}:\mathbb{F}_{q}^{n}\to\mathbb{F}_{q}^{n}$ is a round function. The $\mathbf{x}_{i}$’s are intermediate state variables with $\mathbf{x}_{0}$ the input state and $\mathbf{x}_{r}$ the output state of the function and $\mathbf{y}$ the cryptographic secret variables.
  2. (2)
    Compute a Gröbner basis with respect to an efficient term order, e.g. the degree reverse lexicographic (DRL) term order.
  3. (3)
    Convert to an elimination order, e.g. the lexicographic (LEX) term order.
  4. (4)
    Solve the univariate equation.
Let $I\subset P=\mathbb{F}_{q}[x_{1},\dots,x_{n}]$ be a zero-dimensional ideal, let $d=\dim_{\mathbb{F}_{q}}\left(P/I\right)$ denote the $\mathbb{F}_{q}$-vector space dimension of the quotient space, and assume that a DRL Gröbner basis for $I$ is known. The original FGLM algorithm [FGLM93] for term order conversion runs in $\mathcal{O}\left(n\cdot d^{3}\right)$. An improved probabilistic FGLM variant [FGHR14] runs in $\tilde{\mathcal{O}}\left(d^{\omega}\right)$, where $2.372862\leq\omega<$ 2.37286, if the ideal is in shape position, and in $\tilde{\mathcal{O}}\left(n\cdot d^{\omega}\right)$ if a LEX Gröbner basis of the radical ideal is computed. Further, a sparse linear algebra FGLM variant [FM17] achieves $\mathcal{O}\left(\frac{d^{2+\frac{n-1}{n}}}{\sqrt{n}}\right)$. To factor the univariate LEX polynomial efficiently Bariant et al. [BBLP22, §3.1] proposed to compute its GCD with the field equation. To improve from the generic GCD complexity they propose to perform an auxiliary division by remainder to reduce both polynomials to the same degree. Provided that just a view solutions stem from the base field $\mathbb{F}_{q}$, the complexity of $\mathbb{F}_{q}$-valued root extraction via the GCD is given by
$\mathcal{O}\Big{(}d\cdot\log(d)\cdot\log\big{(}\log(d)\big{)}\cdot\big{(}\log(d)+\log(q)\big{)}\Big{)},$ (1)
if $d\leq q$ else one has to replace the roles of $d$ and $q$ in the complexity estimate.
To estimate the complexity of Gröbner basis computations the Ciminion designers assumed that their polynomial systems are regular [Eis95, Chapter 17] and the Hydra designers assumed that their polynomial systems are semi-regular [Frö85, Par10]. Under these assumptions it is well-known that the complexity of Gröbner basis computations can be estimated via the degree of regularity [BFS04], but both designs lack a formal proof for their genericity assumptions.
Recently, Caminata & Gorla [CG21] introduced the notion of being in generic coordinates. Essentially, if a polynomial system is in generic coordinates, then the complexity of DRL Gröbner basis computations can always be expressed in terms of the Macaulay bound. We also note that for a regular polynomial system the degree of regularity is equal to the Macaulay bound.
In another recent work, Steiner [Ste24] analyzed several cryptographic polynomial systems of MiMC [AGR+16], Hades [GLR+20] and GMiMC [AGP+19] for being in generic coordinates. In particular, he derived an efficient criterion [Ste24, Theorem 3.2] to prove that a polynomial system is in generic coordinates. Moreover, in his analysis of GMiMC he revealed that iterated polynomial systems for Feistel polynomial systems can be verified to be in generic coordinates by computing the rank of an associated linear system. Our work can be seen as generalization of his work since Ciminion and Hydra essentially boil down to Feistel-like polynomial systems.
1.1 Our Results
For the iterated Ciminion polynomial system we prove in LABEL:Th:_Ciminion_Gr\"obner_basis that a DRL Gröbner basis can be constructed via affine transformations. This trivializes Step (2) in a Gröbner basis attack, and one can directly proceed with term order conversion or other polynomial system solving techniques to solve for the key.
It is also worthwhile mentioning that Bariant [Bar23] recently described that the univariate LEX polynomial of Ciminion can be constructed via univariate polynomial multiplication. We review this construction in more detail in Section 3.2, most importantly his attack can be prevented with additional key additions in the rounds of Ciminion. Moreover, if the keys are only added to the non-linear components, then we can maintain the iterated DRL Ciminion Gröbner basis (LABEL:Th:_Ciminion_Gr\"obner_basis). We call this slightly modified variant Ciminion2 (Section 3.3.1).
For Hydra, in Theorem 4.1 we derive a linear system to verify generic coordinates for the iterated Hydra polynomial system. In case the Hydra polynomial system is in generic coordinates, then we can also use the linear system to perform a change of coordinates and produce a quadratic polynomial system which contains a zero-dimensional DRL Gröbner basis. For Hydra this gives us to options to solve for the key.
  1. (I)
    We ignore the additional equations and just use the DRL Gröbner basis to proceed with polynomial system solving.
  2. (II)
    We exploit the quadratic DRL Gröbner basis to reduce the standard Macaulay matrix for the overdetermined Hydra polynomial system to a submatrix whose columns are only indexed by square-free monomials.
For (I), to compute the solutions of our polynomial systems we take a different approach than the classical term order conversion to LEX. We propose to compute the zeros via the eigenvalues of the so-called multiplication matrices, see Section 2.5. Then, the complexity of the Eigenvalue Method is dominated by the complexity of determinant computations. We also note that the probabilistic FGLM algorithm [FM17] boils down to matrix multiplications with the multiplication matrices, hence the generic Eigenvalue Method and the probabilistic FGLM algorithm share the same complexity estimate up to multiplicative constants. In particular, in Section 5 we discuss how the multiplicative structure of quadratic DRL Gröbner bases can be exploited to improve the complexity of the determinant for the characteristic polynomial. Consequently, this will yield the best currently known attacks on Hydra.
For (II), computing a DRL Gröbner basis for multivariate quadratic polynomial systems is an omnipresent problem in cryptography. However, for the full Hydra polynomial system we have an advantage over the general problem, because the system already contains a zero-dimensional DRL Gröbner basis. Therefore, to decrease the size of the Macaulay matrices in a Gröbner basis computation, we propose to perform an auxiliary division by remainder computation with respect to the Gröbner basis. Then, the remainders will only consist of square-free monomials, so we can reduce Gaussian elimination to a submatrix of the Macaulay matrix whose columns are only indexed by square-free monomials. We discuss this approach and its complexity in Section 6. We also note that the reduction to square-free Macaulay matrices for polynomial systems $\mathcal{F}\subset\mathbb{F}_{2}[x_{1},\dots,x_{n}]$ was already discussed by Bardet et al. [BFSS13]. Over $\mathbb{F}_{2}$ one can always include the field equations $F=\left(x_{1}^{2}-x_{1},\dots,x_{n}^{2}-x_{n}\right)$ to restrict to the $\mathbb{F}_{2}$-valued solutions of $\mathcal{F}$. Obviously, the complexity of division by remainder with respect to $F$ is negligible compared to the cost of Gaussian elimination. This is not the case for our Hydra polynomial system, in particular a simple inequality reveals that for Hydra the auxiliary division by remainder step, i.e. the construction of square-free Macaulay matrices, dominates the complexity of DRL Gröbner basis computations (Equation 66). In addition, to highlight the analogy from polynomial system solving over $\mathbb{F}_{2}$, over an arbitrary field $K$ we call square-free Macaulay matrices $K$-Boolean Macaulay matrices.
We implemented Ciminion and Hydra polynomial systems in the computer algebra systems SageMath [Sag24] and OSCAR [OSC24]. [^1] https://github.com/sca-research/Groebner-Basis-Cryptanalysis-of-Ciminion-and-Hydra.git In particular, for Hydra we have implemented the linear system for the generic coordinates verification as well as the linear change of coordinates to produce a quadratic DRL Gröbner basis. We stress that our SageMath implementations is able to handle practical parameters like $128$ bit prime numbers and full rounds Hydra with $r_{\mathcal{H}}=39$.
Finally, in Section 7 we discuss how our results can be applied for cryptanalysis of Ciminion and Hydra.
  • For Ciminion2, utilizing the DRL Gröbner basis the cost of solving for the first key pair is proportional to
    $\mathcal{O}\left(d^{\omega\cdot(r_{C}+r_{E}-2)+1}\right).$ (2)
    Let $64\leq s\leq\log_{2}\left(q\right)$ denote the target security level of Ciminion, then one requires that $r_{E}\in\left\{s+6,\left\lceil\frac{2\cdot(s+6)}{3}\right\rceil\right\}$ [DGGK21, Table 1]. Since $\omega\geq 2$, this complexity will always exceed the target security level.
  • For Hydra, extracting a zero-dimensional DRL Gröbner basis and solving via the Eigenvalue Method turns out to be the most competitive attack vector. In [GØSW23, §7.4] the Hydra designers claim that $r_{\mathcal{H}}=31$ rounds are sufficient to achieve at least $128$ bits of security against Gröbner basis attacks. Note that this claim is already invalidated if we perform term order conversion to LEX via the probabilistic FGLM algorithm. Via the Eigenvalue Method we can attack up to $r_{\mathcal{H}}=33$ rounds below $128$ bit complexity. However, Hydra with full rounds $r_{\mathcal{H}}=39$ achieves $128$ bits security against all strategies analyzed in this paper. Our complexities for various attack strategies on Hydra are summarized in Table 1.
Table 1: Summary of Hydra complexity estimations for various Gröbner basis attacks, see Tables 4, 5 and 6. All computations are performed over the prime $q=2^{127}+45$ with $\omega=2$.
$r_{\mathcal{H}}$ Eigenvalue Method Term Order Conversion Proven Solving Degree Boolean Semi-Regular Estimate Semi-Regular Estimate [GØSW23]
$20$ $\mathbf{75}$ $82$ $105$ $91$ $89$
$29$ $\mathbf{111}$ $118$ $151$ $130$ $131$
$31$ $\mathbf{119}$ $126$ $161$ $140$ $142$
$33$ $\mathbf{127}$ $135$ $172$ $149$ $153$
$34$ $\mathbf{131}$ $139$ $177$ $154$ $159$
$39$ $\mathbf{151}$ $159$ $203$ $176$ $183$
2 Preliminaries
We denote fields with $K$ and their algebraic closure by $\bar{K}$. If the field and the number of variables are clear from context, then we abbreviate the polynomial ring as $P=K[x_{1},\dots,x_{n}]$. Matrices $\mathbf{M}\in k^{m\times n}$ are denoted with upper bold letters and vectors $\mathbf{v}\in\mathbb{F}_{q}^{n}$ by lower bold letters. Matrix-vector and matrix-matrix products are written as $\mathbf{M}\mathbf{v}$.
For polynomial ideals $I\subset K[x_{1},\dots,x_{n}]$, we denote the zero locus, i.e. its set of zeros, over $\bar{k}$ as
$\mathcal{Z}\left(I\right)=\left\{\mathbf{p}\in\bar{k}^{n}\mid f(\mathbf{p})=0,\ \forall f\in I\right\}\subset\mathbb{A}^{n}_{\bar{k}}.$ (3)
For homogeneous ideals $I$ we denote the projective zero locus over $\bar{K}$ by $\mathcal{Z}_{+}\left(I\right)\subset\mathbb{P}^{n-1}_{\bar{k}}$.
Let $f\in K[x_{1},\dots,x_{n}]$ be a polynomial, and let $x_{0}$ be an additional variable, the homogenization of $f$ with respect to $x_{0}$ is defined as
$f^{\text{hom}}(x_{0},\dots,x_{n})=x_{0}^{\deg\left(f\right)}\cdot f\left(\frac{x_{1}}{x_{0}},\dots,\frac{x_{n}}{x_{0}}\right)\in K[x_{0},\dots,x_{n}].$ (4)
Analog, we denote the homogenization of ideals $I^{\text{hom}}=\left\{f^{\text{hom}}\mid f\in I\right\}$ and of finite systems of polynomials $\mathcal{F}^{\text{hom}}=\left\{f_{1}^{\text{hom}},\dots,f_{m}^{\text{hom}}\right\}$.
For a homogeneous ideal $I\subset P$ and an integer $d\geq 0$ we denote
$I_{d}=\left\{f\in I\mid\deg\left(f\right)=d,\ f\text{ homogeneous}\right\}.$ (5)
For a term order $>$ and an ideal $I\subset P$ we denote with $\operatorname{in}_{>}(I)=\{\operatorname{LT}_{>}(f)\mid f\in I\}$ the initial ideal of $I$, i.e. the ideal of leading terms of $I$, with respect to $>$.
Every polynomial $f\in K[x_{1},\dots,x_{n}]$ of degree $d$ can be written as $f=f_{d}+f_{d-1}+\ldots+f_{0}$, where $f_{i}$ is homogeneous of degree $i$. We denote its highest degree component $f_{d}$ of $f$ with $f^{\text{top}}$, and analog we denote $\mathcal{F}^{\text{top}}=\left\{f_{1}^{\text{top}},\dots,f_{m}^{\text{top}}\right\}$. Note that
$f^{\text{top}}=f^{\text{hom}}\mod(x_{0}).$ (6)
Let $I,J\subset K[x_{1},\dots,x_{n}]$ be ideals, then we denote with
$I:J=\left\{f\in K[x_{1},\dots,x_{n}]\mid\forall g\in J\colon f\cdot g\in I\right\}$ (7)
the usual ideal quotient, and with
$I:J^{\infty}=\bigcup_{i\geq 1}I:J^{i}$ (8)
the saturation of $I$ with respect to $J$.
Let $I,\mathfrak{m}\in K[x_{0},\dots,x_{n}]$ be homogeneous ideals where $\mathfrak{m}=(x_{0},\dots,x_{n})$, then we call $I^{\text{sat}}=I:\mathfrak{m}^{\infty}$ the saturation of $I$.
Since we analyze vector-valued functions $F:\mathbb{F}_{q}^{n}\to\mathbb{F}_{q}^{n}$ we also denote vectors of polynomial ring variables as $\mathbf{x}=(x_{1},\dots,x_{n})^{\intercal}$. Moreover, for a term order $>$ on $P$, if not specified otherwise we always assume that the variables within a vector $\mathbf{x}$ are naturally ordered with respect to $>$, i.e. $x_{1}>\ldots>x_{n}$.
We will often encounter the lexicographic and the degree reverse lexicographic term order which we will abbreviate as LEX and DRL respectively.
2.1 Ciminion
Let $\mathbb{F}_{q}$ be a finite field, the Ciminion [DGGK21] PRF over $\mathbb{F}_{q}$ is most easily described via an illustration, see Figure 1. A nonce $\aleph\in\mathbb{F}_{q}$ and the first key pair $(K_{1},K_{2})$ are fed into the permutation $p_{C}:\mathbb{F}_{q}^{3}\to\mathbb{F}_{q}^{3}$. Then the outputs are fed into the permutation $p_{E}:\mathbb{F}_{q}^{3}\to\mathbb{F}_{q}^{3}$. The first two outputs of $p_{E}$ can then be added to messages $P_{1},P_{2}\in\mathbb{F}_{q}$ to encrypt them, and the third output remains private. If one wishes to encrypt an additional message pair $P_{3},P_{4}\in\mathbb{F}_{q}$ one feeds the outputs of $p_{C}$ after a key addition with $(K_{3},K_{4})$ into the rolling function $rol:\mathbb{F}_{q}^{3}\to\mathbb{F}_{q}^{3}$ and then again into $p_{E}$.
The permutations $p_{C}$ and $p_{E}$ are based on an iterated Feistel network
$\begin{split}\mathcal{R}^{(l)}:\mathbb{F}_{q}^{3}&\to\mathbb{F}_{q}^{3},\\ \begin{pmatrix}x\\ y\\ z\end{pmatrix}&\mapsto\begin{pmatrix}0&0&1\\ 1&c_{4}^{(l)}&c_{4}^{(l)}\\ 0&1&1\end{pmatrix}\begin{pmatrix}x\\ y\\ z+x\cdot y\end{pmatrix}+\begin{pmatrix}c_{1}^{(l)}\\ c_{2}^{(l)}\\ c_{3}^{(l)}\end{pmatrix},\end{split}$ (9)
where $c_{1}^{(l)},c_{2}^{(l)},c_{3}^{(l)}\in\mathbb{F}_{q}$ and $c_{4}^{(l)}\in\mathbb{F}_{q}\setminus\{0,1\}$. Note that the Feistel network $(x,y,z)^{\intercal}\mapsto(x,y,z+x\cdot y)^{\intercal}$ is also known as Toffoli gate which plays a special role in reversible computing [Tof80].
Figure 1: Encryption with Ciminion over a finite field $\mathbb{F}_{q}$.
The permutation $p_{C}$ is defined as the $r_{C}$-fold composition of $\mathcal{R}^{(l)}$, and the function $p_{E}$ is defined as the $r_{E}$-fold composition of $\mathcal{R}^{(l)}$. Finally, the rolling function is defined as
$\begin{split}rol:\mathbb{F}_{q}^{3}&\to\mathbb{F}_{q}^{3},\\ \begin{pmatrix}x\\ y\\ z\end{pmatrix}&\mapsto\begin{pmatrix}z+x\cdot y\\ x\\ y\end{pmatrix}.\end{split}$ (10)
Formulae to compute the round numbers for Ciminion are given in Table 2.
Table 2: Proposed round numbers for Ciminion [DGGK21, Table 1]. The security level $s$ must satisfy $64\leq s\leq\log_{2}\left(q\right)$, and $q\geq 2^{64}$.
Instance $p_{C}$ $p_{E}$
Standard $s+6$ $\max\left\{\left\lceil\frac{s+37}{12}\right\rceil,6\right\}$
Data limit $2^{\frac{s}{2}}$ elements $\frac{2\cdot(s+6)}{3}$ $\max\left\{\left\lceil\frac{s+37}{12}\right\rceil,6\right\}$
Conservative $s+6$ $\max\left\{\left\lceil\frac{3}{2}\cdot\frac{s+37}{12}\right\rceil,9\right\}$
E.g., for $s=128$ and data limit $2^{\frac{s}{2}}$ we have the round numbers $r_{C}=90$ and $r_{E}=14$.
It is now straight-forward to set up the iterated polynomial system for the first key pair $(K_{1},K_{2})$ given the first plain/ciphertext pair. For ease of writing we denote the $i$th intermediate state variables as $\mathbf{x}_{j}^{(i)}=\left(x_{j,1}^{(i)},x_{j,2}^{(i)},x_{j,3}^{(i)}\right)^{\intercal}$ where $j\in\{C,E\}$.
Definition 2.1 (Iterated polynomial system for Ciminion).
Let $\mathbb{F}_{q}$ be a finite field, and let $\big{(}\aleph,(p_{1},p_{2}),(c_{1},c_{2})\big{)}\in\mathbb{F}_{q}\times\mathbb{F}_{q}^{2}\times\mathbb{F}_{q}^{2}$ be a nonce/plain/ciphertext sample given by a Ciminion encryption function. We call the polynomial system $\mathcal{F}_{\texttt{Ciminion}}=\left\{\mathbf{f}^{(i)}\right\}_{1\leq i\leq e_{C}+r_{E}}\subset\mathbb{F}_{q}\Big{[}y_{1},y_{2},\mathbf{x}_{C}^{(1)},\dots,\mathbf{x}_{C}^{(r_{C}-1)},\allowbreak\mathbf{x}_{E}^{(1)},\dots,\mathbf{x}_{E}^{(r_{E})},x\Big{]}$ the iterated Ciminion polynomial system
$\mathbf{f}^{(i)}=\begin{dcases}\mathcal{R}^{(1)}\left(\aleph,y_{1},y_{2}\right)-\mathbf{x}_{C}^{(1)},&i=1,\\ \mathcal{R}^{(i)}\left(\mathbf{x}_{C}^{(i-1)}\right)-\mathbf{x}_{C}^{(i)},&2\leq i\leq r_{C}-1,\\ \mathcal{R}^{(i)}\left(\mathbf{x}_{C}^{(r_{C}-1)}\right)-\mathbf{x}_{E}^{(1)},&i=r_{C},\\ \mathcal{R}^{(i-r_{C})}\left(\mathbf{x}_{E}^{(i-r_{C})}\right)-\mathbf{x}_{E}^{(i-r_{C}+1)},&r_{C}+1\leq i\leq r_{C}+r_{E}-1,\\ \mathcal{R}^{(r_{E})}\left(\mathbf{x}_{E}^{(r_{E})}\right)-\begin{pmatrix}c_{1}-p_{1}\\ c_{2}-p_{2}\\ x\end{pmatrix},&i=r_{C}+r_{E}.\end{dcases}$
2.2 Hydra
Let $\mathbb{F}_{q}$ be a finite field, the Hydra [GØSW23] PRF over $\mathbb{F}_{q}$ is most also easily described via an illustration, see Figure 2. A nonce $\aleph\in\mathbb{F}_{q}$ and an initial value $\texttt{IV}\in\mathbb{F}_{q}^{3}$ are added to a key $\texttt{K}\in\mathbb{F}_{q}^{4}$, and then fed into the keyed body function $\mathcal{B}:\mathbb{F}_{q}^{4}\times\mathbb{F}_{q}^{4}\to\mathbb{F}_{q}^{8}$ which produces the state $\mathbf{y}||\mathbf{z}=(\mathbf{y},\mathbf{z})^{\intercal}\in\mathbb{F}_{q}^{4}\times\mathbb{F}_{q}^{4}$. Then $\mathbf{y}||\mathbf{z}$ is fed into the first keyed head function $\mathcal{H}_{\texttt{K}}:\mathbb{F}_{q}^{8}\times\mathbb{F}_{q}^{4}\to\mathbb{F}_{q}^{8}$. After adding $\mathbf{y}||\mathbf{z}$ to the output of the head the result is released as the first Hydra sample. Say one has already requested $(i-1)$ Hydra samples, for an additional sample $\mathbf{y}||\mathbf{z}$ is fed into the $i$th rolling function $\mathcal{R}_{i}:\mathbb{F}_{q}^{8}\to\mathbb{F}_{q}^{8}$, and then again into the head $\mathcal{H}_{\texttt{K}}$. After adding the output of $\mathcal{R}_{i}$ to the output of $\mathcal{H}_{\texttt{K}}$ the result is released as $i$th Hydra sample.
Figure 2: The Hydra PRF with hidden body.
For our polynomial model we are only going to consider the heads of the Hydra. Thus, we do not describe the details of the body function $\mathcal{B}$, interested readers can find the description in [GØSW23, §5.2]. Let $\mathbf{M}_{\mathcal{E}}\in\operatorname{GL}_{4}\left(\mathbb{F}_{q}\right)$ be an MDS matrix, e.g. the Hydra designers recommend $\operatorname{circ}\left(2,3,1,1\right)$ or $\operatorname{circ}\left(3,2,1,1\right)$222 We consider circulant matrices with respect to a right shift, i.e.  $\operatorname{circ}(a_{1},\dots,a_{n})=\begin{pmatrix}a_{1}&a_{2}&\dots&a_{n-1}&a_{n}\\ a_{n}&a_{1}&\dots&a_{n-2}&a_{n-1}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ a_{2}&a_{3}&\dots&a_{n}&a_{1}\end{pmatrix}.$ For the keyed permutation $\mathcal{H}_{\texttt{K}}$, let
$\texttt{K}^{\prime}$ $=\begin{pmatrix}\texttt{K}\\ \mathbf{M}_{\mathcal{E}}\texttt{K}\end{pmatrix}\in\mathbb{F}_{q}^{8},$ (11)
$f(\mathbf{x})$ $=\left(\sum_{i=1}^{8}(-1)^{\left\lfloor\frac{i-1}{4}\right\rfloor}\cdot x_{i}\right)^{2}.$ (12)
Moreover, let $\mathbf{M}_{\mathcal{J}}\in\operatorname{GL}_{8}\left(\mathbb{F}_{q}\right)$ be an invertible matrix. Then the heads round function is defined as
$\begin{split}\mathcal{J}_{i}:\mathbb{F}_{q}^{8}\times\mathbb{F}_{q}^{4}&\to\mathbb{F}_{q}^{8},\\ \left(\mathbf{x},\texttt{K}\right)&\mapsto\mathbf{M}_{\mathcal{J}}\big{(}x_{i}+f(\mathbf{x})\big{)}_{1\leq i\leq 8}+\texttt{K}^{\prime}+\mathbf{c}^{(i)},\end{split}$ (13)
where $\mathbf{c}^{(i)}\in\mathbb{F}_{q}^{8}$ is a constant. The heads function is now defined as the following $r_{\mathcal{H}}$-fold composition
$\begin{split}\mathcal{H}_{\texttt{K}}:\mathbb{F}_{q}^{4}\times\mathbb{F}_{q}^{4}\times\mathbb{F}_{q}^{4}&\to\mathbb{F}_{q}^{8},\\ (\mathbf{y},\mathbf{z},\texttt{K})&\mapsto\mathcal{J}_{r_{\mathcal{H}}}\circ\cdots\circ\mathcal{J}_{1}\left(\begin{pmatrix}\mathbf{y}\\ \mathbf{z}\end{pmatrix},\texttt{K}\right).\end{split}$ (14)
Let $\mathbf{M}_{\mathcal{I}}\in\operatorname{GL}_{4}\left(\mathbb{F}_{q}\right)$ be a matrix of the form
$\mathbf{M}_{\mathcal{I}}=\begin{pmatrix}\mu_{1,1}&1&1&1\\ \mu_{2,1}&\mu_{2,2}&1&1\\ \mu_{3,1}&1&\mu_{3,3}&1\\ \mu_{4,1}&1&1&\mu_{4,4}\end{pmatrix}.$ (15)
Finally, for the rolling function let
$g(\mathbf{y},\mathbf{z})$ $=\left(\sum_{i=1}^{4}(-1)^{i-1}\cdot y_{i}\right)\cdot\left(\sum_{i=1}^{4}(-1)^{\left\lfloor\frac{i-1}{2}\right\rfloor}\cdot z_{i}\right),$ (16)
$\mathbf{M}_{\mathcal{R}}$ $=\operatorname{diag}\left(\mathbf{M}_{\mathcal{I}},\mathbf{M}_{\mathcal{I}}\right).$ (17)
Then, the rolling function is defined as
$\begin{split}\mathcal{R}:\mathbb{F}_{q}^{4}\times\mathbb{F}_{q}^{4}&\to\mathbb{F}_{q}^{8},\\ \begin{pmatrix}\mathbf{y}\\ \mathbf{z}\end{pmatrix}&\mapsto\mathbf{M}_{\mathcal{R}}\begin{pmatrix}\{y_{i}+g(\mathbf{y},\mathbf{z})\}_{1\leq i\leq 4}\\ \{z_{i}+g(\mathbf{z},\mathbf{y})\}_{1\leq i\leq 4}\end{pmatrix},\end{split}$ (18)
and the $i$th rolling function is simply the $i$-fold composition $\mathcal{R}_{i}=\mathcal{R}\circ\cdots\circ\mathcal{R}$. Note that the degree increasing parts in $\mathcal{J}_{i}$ and $\mathcal{R}$ are generalized Lai–Massey permutations [RS22].
For $\kappa$ bits of security, the number of rounds for the Hydra heads with data limit $2^{\frac{\kappa}{2}}$ are computed via, see [GØSW23, §5.5],
$r_{\mathcal{H}}=\left\lceil 1.25\cdot\max\left\{24,2+r_{\mathcal{H}}^{\ast}\right\}\right\rceil,$ (19)
where $r_{\mathcal{H}}^{\ast}$ is the smallest integer that satisfies [GØSW23, Equation 12] and [GØSW22, Equation 15]. Note that [GØSW23, Equation 12] is derived from the Gröbner basis cryptanalysis of Hydra. E.g., for $q=2^{127}+45$ and $\kappa=128$ the Hydra designers propose $r_{\mathcal{H}}=39$.
Given two Hydra outputs it is straight-forward to set up an overdetermined polynomial system for Hydra. For ease of writing we denote the $i$th intermediate state variables for the $j$th output as $\mathbf{x}^{(i)}_{j}=\left(x_{j,1}^{(i)},\dots,x_{j,8}^{(i)}\right)^{\intercal}$ and analog for $\mathbf{y}=(y_{1},\dots,y_{4})^{\intercal}$, $\mathbf{z}=(z_{1},\dots,z_{4})^{\intercal}$ and $\mathbf{k}=(k_{1},\dots,k_{4})^{\intercal}$.
Definition 2.2 (Iterated polynomial system for Hydra).
Let $\mathbb{F}_{q}$ be a finite field, and let $\mathbf{c}_{1},\mathbf{c}_{2}\in\mathbb{F}_{q}^{8}$ be the first two outputs of a Hydra function. We call the following polynomial system $\mathcal{F}_{\textsf{Hydra}}=\left\{\mathbf{f}_{1}^{(i)},\mathbf{f}_{\mathcal{R}},\mathbf{f}_{2}^{(i)}\right\}_{1\leq i\leq r_{\mathcal{H}}}\subset\mathbb{F}_{q}\Big{[}\mathbf{y},\mathbf{z},\mathbf{x}_{1}^{(1)},\dots,\mathbf{x}_{1}^{(r_{\mathcal{H}}-1)},\mathbf{x}_{2}^{(0)},\dots,\mathbf{x}_{2}^{(r_{\mathcal{H}}-1)},\mathbf{k}\Big{]}$ the iterated Hydra polynomial system
$\mathbf{f}_{1}^{(i)}$ =$=\begin{dcases}\mathcal{J}_{1}\left(\begin{pmatrix}\mathbf{y}\\ \mathbf{z}\end{pmatrix},\mathbf{k}^{\prime}\right)-\mathbf{x}_{1}^{(1)},&i=1,\\ \mathcal{J}_{i}\left(\mathbf{x}_{1}^{(i-1)},\mathbf{k}^{\prime}\right)-\mathbf{x}_{1}^{(i)},&2\leq i\leq r_{\mathcal{H}}-1,\\ \mathcal{J}_{r_{\mathcal{H}}}\left(\mathbf{x}_{1}^{(r_{\mathcal{H}}-1)},\mathbf{k}^{\prime}\right)+\begin{pmatrix}\mathbf{y}\\ \mathbf{z}\end{pmatrix}-\mathbf{c}_{1},&i=r_{\mathcal{H}},\end{dcases}$
$\mathbf{f}_{\mathcal{R}}$ $=\mathcal{R}\left(\begin{pmatrix}\mathbf{y}\\ \mathbf{z}\end{pmatrix}\right)-\mathbf{x}_{2}^{(0)},$
$\mathbf{f}_{2}^{(i)}$ =$=\begin{dcases}\mathcal{J}_{1}\left(\mathbf{x}_{2}^{(0)},\mathbf{k}^{\prime}\right)-\mathbf{x}_{2}^{(1)},&i=1,\\ \mathcal{J}_{i}\left(\mathbf{x}_{2}^{(i-1)},\mathbf{k}^{\prime}\right)-\mathbf{x}_{2}^{(i)},&2\leq i\leq r_{\mathcal{H}}-1,\\ \mathcal{J}_{r_{\mathcal{H}}}\left(\mathbf{x}_{2}^{(r_{\mathcal{H}})},\mathbf{k}^{\prime}\right)+\mathbf{x}_{2}^{(1)}-\mathbf{c}_{2},&i=r_{\mathcal{H}},\end{dcases}$
where $\mathbf{k}^{\prime}=\begin{pmatrix}\mathbf{k}\\ \mathbf{M}_{\mathcal{E}}\mathbf{k}\end{pmatrix}.$
This polynomial system consists of $816\cdot r_{\mathcal{H}}+$ 8 many polynomials in $416\cdot r_{\mathcal{H}}+$ 4 many variables. It can also easily be expanded to an arbitrary number of output samples.
2.3 Gröbner Bases
Gröbner bases are a fundamental tool in computational algebra. They were first introduced by Bruno Buchberger in his PhD thesis [Buc65]. For their definition one needs to induce an order on the monomials of a polynomial ring $P=K[x_{1},\dots,x_{n}]$. First notice that a monomial $m=\prod_{i=1}^{n}x_{i}^{a_{i}}\in P$ can be identified with its exponent vector $\mathbf{a}=\left(a_{1},\dots,a_{n}\right)^{\intercal}\in\mathbb{Z}_{\geq 0}^{n}$. Therefore, a term order $>$ on $P$ is a binary relation on $\mathbb{Z}_{\geq 0}^{n}$ such that, see [CLO15, Chapter 2 §2 Definition 1]:
  1. (i)
    $>$ is a total ordering in $\mathbb{Z}_{\geq 0}^{n}$.
  2. (ii)
    If $\mathbf{a}>\mathbf{b}$ and $\mathbf{c}\in\mathbb{Z}_{\geq 0}^{n}$, then $\mathbf{a}+\mathbf{c}>\mathbf{b}+\mathbf{c}$.
  3. (iii)
    Every non-empty subset of $\mathbb{Z}_{\geq 0}^{n}$ has a smallest element under $>$, i.e. $>$ is a well-ordering on $\mathbb{Z}_{\geq 0}^{n}$.
Now let $I\subset P=K[x_{1},\dots,x_{n}]$ be an ideal, and let $f\in P$. By $\operatorname{LM}_{>}\left(f\right)$ we denote the leading monomial of $f$, i.e. the largest monomial with non-zero coefficient in $f$ under $>$. A finite ideal basis $\mathcal{G}=\{g_{1},\dots,g_{m}\}\subset I$ such that
$\big{(}\operatorname{LM}_{>}(g_{1}),\dots,\operatorname{LM}_{>}(g_{m})\big{)}=\operatorname{in}_{>}(I)$ (20)
is called a $>$-Gröbner basis of $I$. I.e., the leading monomials of $\mathcal{G}$ are minimal under $>$. With Gröbner bases one can solve many important computational problems like the membership problem or the computation of the zero locus. For a general introduction to Gröbner bases we refer to [KR00, KR05, CLO15].
2.3.1 Linear Algebra-Based Gröbner Basis Algorithms
The first algorithm to compute Gröbner bases was Buchberger’s algorithm, see e.g. [CLO15, Chapter 2 §7]. Given an input polynomial system $\mathcal{F}\subset P=k[x_{1},\dots,x_{n}]$ and a term order $>$ on $P$, Buchberger’s algorithms performs pairwise cancellation of the leading monomials under $>$ of two elements in $\mathcal{F}$ followed by multivariate division by remainder. If a non-zero element is produced, then it is added to $\mathcal{F}$ and the process is repeated. Buchberger proved that this procedure terminates and produces a $>$-Gröbner basis.
Lazard [Laz83] generalized this construction by performing multiple leading term cancellations and reductions at the same time via a matrix representation of the polynomial system $\mathcal{F}$. Assume that =$\mathcal{F}=\{f_{1},\dots,f_{m}\}\subset P$ is homogeneous and fix a term order $>$ on $P$. The homogeneous Macaulay matrix $M_{d}$ for $\mathcal{F}$ has columns indexed by the monomials of degree $d$ in $P$ sorted via $>$, and its rows are indexed by $s\cdot f$, where $f\in\mathcal{F}$ and $s\in P$ is a monomial such that $\deg\left(s\cdot f\right)=d$. The entry of row $s\cdot f$ at column $t$ is the coefficient of $t$ in the polynomial $s\cdot f$. If $\mathcal{F}$ is inhomogeneous, then one replaces the degree equalities by inequalities to obtain the inhomogeneous Macaulay matrix $M_{\leq d}$.
Obviously, if we perform Gaussian elimination on $M_{0},\dots,M_{d}$ respectively $M_{\leq d}$ for a large enough value of $d$, then we construct a $>$-Gröbner basis for $\mathcal{F}$. The least such $d$ is called the solving degree, to the best of our knowledge this notion was first introduced in [DS13].
Definition 2.3 (Solving degree, [CG21, Definition 6]).
Let =$\mathcal{F}=\{f_{1},\dots,f_{m}\}\subset K[x_{1},\dots,\allowbreak x_{n}]$ and let $>$ be a term order. The solving degree of $\mathcal{F}$ is the least degree $d$ such that Gaussian elimination on the Macaulay matrix $M_{\leq d}$ produces a Gröbner basis of $\mathcal{F}$ with respect to $>$. We denote it by $\operatorname{sd}_{>}(\mathcal{F})$.
If $\mathcal{F}$ is homogeneous, we consider the homogeneous Macaulay matrix $M_{d}$ and let the solving degree of $\mathcal{F}$ be the least degree $d$ such that Gaussian elimination on $M_{0},\dots,M_{d}$ produces a Gröbner basis of $\mathcal{F}$ with respect to $>$.
By [CG21, Theorem 7] the homogeneous solving degree always upper bounds the inhomogeneous one, i.e.
$\operatorname{sd}_{DRL}\left(\mathcal{F}\right)\leq\operatorname{sd}_{DRL}\left(\mathcal{F}^{\text{hom}}\right).$ (21)
Algorithms which perform Gaussian elimination on the Macaulay matrices are generally referred to as linear algebra-based Gröbner basis algorithms. Modern examples of such algorithms are Faugère’s F4 [Fau99] and Matrix-F5 [Fau02] algorithms. Since the solving degree is a priori unknown for most polynomial system, F4/5 construct the Macaulay matrices and perform Gaussian elimination for increasing values of $d$. Therefore, they require a stopping criterion to decide whether a Gröbner basis has already been found. On the other hand, we can always stop F4/5 after the final Gaussian elimination in the solving degree, since the row space basis must already contain a Gröbner basis. For this reason, we also consider the solving degree as complexity measure for advanced linear algebra-based algorithms.
2.3.2 Complexity Estimates via the Solving Degree
It is well-known that the complexity of Gaussian elimination of a matrix $\mathbf{A}\in K^{n\times m}$, where $K$ is a field and $r=\operatorname{rank}\left(\mathbf{A}\right)$, is given by [Sto00, §2.2] $\mathcal{O}\left(n\cdot m\cdot r^{\omega-2}\right)$ field operations, where $2.372862\leq\omega<$ 2.37286 is a linear algebra constant [AW21].
Given a homogeneous polynomial system =$\mathcal{F}=\{f_{1},\dots,f_{m}\}\subset P=K[x_{1},\dots,x_{n}]$, it is well-known that the number of monomials in $P$ of degree $d$ is $\binom{n+d-1}{d}$. Moreover, for $f\in\mathcal{F}$ the number of rows $\hat{m}\cdot f$, where $\hat{m}\in P$ a monomial such that $\deg\left(\hat{m}\cdot f\right)=d$, is also bounded by $\binom{n+d-1}{d}$. Thus, the complexity of Gaussian elimination on the homogeneous Macaulay matrices $M_{0},\dots,M_{d}$ is bounded by
$\mathcal{O}\left(m\cdot d\cdot\binom{n+d-1}{d}^{\omega}\right).$ (22)
Also, note that dehomogenization of a Gröbner basis yields an inhomogeneous Gröbner basis [KR05, Proposition 4.3.18] (if the term order is compatible with homogenization). Therefore, we can also regard Equation 22 as complexity estimate for Gaussian elimination on the inhomogeneous Macaulay matrix $M_{\leq d}$.
2.4 Generic Coordinates
Caminata & Gorla [CG23] have proven that for a special class of polynomial systems the DRL solving degree is always upper bounded by the so-called Macaulay bound. They called these polynomial to be in generic coordinates. For the formal definition of this notion, we use the shorthand notation “$r\nmid 0\mod I$” for expressing that $r$ is not a zero-divisor in $P/I$. Moreover, we always assume that the homogenization variable $x_{0}$ is the least variable under DRL.
Definition 2.4 ([CG21, CG22, Definition 5]).
Let $K$ be an infinite field. Let $I\subset K[x_{0},\dots,\allowbreak x_{n}]$ be a homogeneous ideal with $|\mathcal{Z}_{+}(I)|<\infty$. We say that $I$ is in generic coordinates if either $|\mathcal{Z}_{+}(I)|=0$ or $x_{0}\nmid 0\mod I^{\text{sat}}$.
Let $K$ be any field, and let $K\subset L$ be an infinite field extension. $I$ is in generic coordinates over $K$ if $I\otimes_{K}L[x_{0},\dots,x_{n}]\subset L[x_{0},\dots,x_{n}]$ is in generic coordinates.
As already mentioned, polynomial systems in generic coordinates have a proven DRL solving degree upper bound.
Theorem 2.5 ([CG21, Theorem 9, 10, Corollary 2]).
Let $K$ be an algebraically closed field, and let =$\mathcal{F}=\{f_{1},\dots,f_{m}\}\subset K[x_{1},\dots,x_{n}]$ be an inhomogeneous polynomial system such that $\left(\mathcal{F}^{\text{hom}}\right)$ is in generic coordinates. Let $d_{i}=\deg\left(f_{i}\right)$ and assume that $d_{1}\geq\ldots\geq d_{m}$, then for $l\in\min\{n+1,m\}$
$\operatorname{sd}_{DRL}\left(\mathcal{F}\right)\leq d_{1}+\ldots+d_{l}-l+1.$
In the Gröbner basis and algebraic cryptanalysis literature the latter inequality is also usually referred as Macaulay bound. Moreover, Lazard [Laz83] first observed that DRL Gröbner bases for “generic” polynomial systems can be found within the Macaulay bound.
Finally, being in generic coordinates was recently characterized by Steiner [Ste24].
Theorem 2.6 ([Ste24, Theorem 3.2]).
Let $K$ be an algebraically closed field, and let =$\mathcal{F}=\{f_{1},\dots,f_{m}\}\subset K[x_{1},\dots,x_{n}]$ be a polynomial system such that
  1. (i)
    $(\mathcal{F})\neq(1)$, and
  2. (ii)
    $\dim\left(\mathcal{F}\right)=0$.
Then the following are equivalent.
  1. (1)
    $\left(\mathcal{F}^{\text{hom}}\right)$ is in generic coordinates.
  2. (2)
    $\sqrt{\mathcal{F}^{\text{top}}}=\left(x_{1},\dots,x_{n}\right)$.
  3. (3)
    $\left(\mathcal{F}^{\text{top}}\right)$ is zero-dimensional in $K[x_{1},\dots,x_{n}]$.
  4. (4)
    For every $1\leq i\leq n$ there exists $d_{i}\in\mathbb{Z}_{\geq 1}$ such that $x_{i}^{d_{i}}\in\operatorname{in}_{DRL}\left(\mathcal{F}^{\text{hom}}\right)$.
Let $\mathcal{F}\subset K[x_{1},\dots,x_{n}]$ be an inhomogeneous polynomial system, to prove that $\mathcal{F}^{\text{hom}}$ is in generic coordinates one can use the following procedure based on (2):
  1. (1)
    Compute $\left(\mathcal{F}^{\text{top}}\right)=\left(\mathcal{F}^{\text{hom}}\right)\mod(x_{0})$.
  2. (2)
    Construct a monomial $x_{i}^{d}\in\left(\mathcal{F}^{\text{top}}\right)$ for some $1\leq i\leq n$ and $d\geq 1$.
  3. (3)
    Replace $\left(\mathcal{F}^{\text{top}}\right)=\left(\mathcal{F}^{\text{top}}\right)\mod(x_{i})$ and return to step (2).
If Step (2) is successful for all $1\leq i\leq n$, then the polynomial system $\mathcal{F}^{\text{hom}}$ is in generic coordinates.
Another quantity in the literature that is often used to express the complexity of finding a DRL Gröbner basis is the so-called degree of regularity.
Definition 2.7 (Degree of regularity, [BFS04, Definition 4]).
Let $K$ be a field, and let $\mathcal{F}\subset P=K[x_{1},\dots,x_{n}]$. Assume that $\left(\mathcal{F}^{\text{top}}\right)_{d}=P_{d}$ for some integer $d\geq 0$. The degree of regularity is defined as
$d_{\operatorname{reg}}\left(\mathcal{F}\right)=\min\left\{d\geq 0\;\middle|\;\left(\mathcal{F}^{\text{top}}\right)_{d}=P_{d}\right\}.$
Note that by Theorem 2.6 a polynomial system is in generic coordinates if and only if it admits a finite degree of regularity.
2.5 The Eigenvalue Method
As outlined in the introduction, in the algebraic cryptanalysis literature the standard approach to polynomial system solving is
  1. (1)
    DRL Gröbner basis computation,
  2. (2)
    term order conversion to LEX, and
  3. (3)
    univariate factoring.
In this section we shortly discuss that there is an alternative to term order conversion via linear algebra-based techniques [KR16, Chapter 6]. Our motivation to change the strategy is quite simple. Via the Eigenvalue Method we will be able to exploit the multiplicative structure of Ciminion and Hydra DRL Gröbner bases. Consequently, this will provide a slight improvement compared to the generic complexity of the probabilistic FGLM algorithm [FGHR14].
Let $I\subset P=K[x_{1},\dots,x_{n}]$ be a zero-dimensional ideal, i.e. $P/I$ is a finite dimensional $K$-vector space [KR00, Proposition 3.7.1], and let $d=\dim_{K}\left(P/I\right)$. If $>$ is a term order on $P$ and $\mathcal{G}\subset I$ is $>$-Gröbner basis, then a $K$-vector space basis $\mathcal{B}$ of $P/I$ is given by the monomials not contained in $\operatorname{in}_{>}\left(\mathcal{G}\right)=\operatorname{in}_{>}\left(I\right)$. Now let $f\in P$, then the multiplication map
$\begin{split}\theta_{f}:P&\to P/I,\\ x&\mapsto x\cdot f,\end{split}$ (23)
can be represented via a matrix. This matrix is called the multiplication matrix $\mathbf{M}_{f}$ for $f$. Given a $>$-Gröbner basis its computation is straight-forward: Index the columns of $\mathbf{M}_{f}$ by the $K$-vector space basis $\mathcal{B}$, and the rows by $b\cdot f$, where $b\in\mathcal{B}$. The row $b\cdot f$ is then simply the coefficient vector of $b\cdot f$ in $P/I$ which can be computed via $b\cdot f\mod(\mathcal{G})$ with respect to $>$.
Now assume in addition that $K$ is algebraically closed, then for every point $\mathbf{x}\in\mathcal{Z}\left(I\right)$ the point $f(\mathbf{x})$ is an eigenvalue of $\mathbf{M}_{f}$, see [KR16, Proposition 6.2.1]. Therefore, we can compute the zero locus $\mathcal{Z}\left(I\right)$ as follows: Compute the eigenvalues of $\mathbf{M}_{x_{1}},\dots,\mathbf{M}_{x_{n}}$, iterate over all possible combinations, and verify whether a combination is indeed a solution to $I$. This approach is known as Eigenvalue Method [KR16, §6.2.A].
In his seminal work Strassen showed that matrix inversion, determinant and Gaussian elimination have the same complexity as matrix multiplication up to some constant [Str69]. For the determinant this observation follows form LU-decomposition of block square matrices. We recall this decomposition since we will exploit it at a later point in this paper. Let $\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D}\in K^{2\cdot N\times 2\cdot N}$ be matrices, then we have the block matrix decomposition
$\begin{pmatrix}\mathbf{A}&\mathbf{B}\\ \mathbf{C}&\mathbf{D}\end{pmatrix}=\begin{pmatrix}\mathbf{I}_{N\times N}&\mathbf{0}_{N\times N}\\ \mathbf{C}\mathbf{A}^{-1}&\mathbf{I}_{N\times N}\end{pmatrix}\begin{pmatrix}\mathbf{A}&\mathbf{B}\\ \mathbf{0}_{N\times N}&\mathbf{D}-\mathbf{C}\mathbf{A}^{-1}\mathbf{B}\end{pmatrix},$ (24)
provided that $\mathbf{A}$ is invertible. With well-known formulae for block matrix determinants, see e.g. [Sil00], one then has that
$\det\begin{pmatrix}\mathbf{A}&\mathbf{B}\\ \mathbf{C}&\mathbf{D}\end{pmatrix}=\det\left(\mathbf{A}\right)\cdot\det\left(\mathbf{D}-\mathbf{C}\mathbf{A}^{-1}\mathbf{B}\right).$ (25)
So the determinant of the block matrix has been reduced to two determinants of smaller matrices, one matrix inversion and two matrix multiplications. By Strassen [Str69] the complexity of multiplying two $N\times N$ matrices is in $\mathcal{O}\left(N^{\omega+o(1)}\right)$, where $2.372862\leq\omega<$ 2.37286 is again the linear algebra constant [AW21]. Note that the term $N^{o(1)}$ is usually ignored in complexity estimations.
Lastly, it is well known that the eigenvalues of a matrix $\mathbf{M}\in K^{N\times N}$ are roots of its characteristic polynomial
$\chi_{\mathbf{M}}(x)=\det\left(\mathbf{I}_{N\times N}\cdot x-\mathbf{M}\right)\in K[x].$ (26)
Therefore, construction of the characteristic polynomials in the Eigenvalue Method can be performed in
$\mathcal{O}\left(n\cdot d^{\omega}\right),$ (27)
where $d=\dim_{K}\left(P/I\right)$. In addition, if we are only interested in solutions over a finite field $\mathbb{F}_{q}$, then factorization of the characteristic polynomials can be done via the GCD with the field equation whose complexity was given in Equation 1.
Note that the probabilistic FGLM algorithm [FGHR14, Proposition 3] has polylogarithmic complexity $\tilde{\mathcal{O}}\left(d^{\omega}\right)$ for the construction of a LEX Gröbner basis of $I$ if $I$ is in shape position, and polylogarithmic complexity $\tilde{\mathcal{O}}\left(n\cdot d^{\omega}\right)$ for the construction of a LEX Gröbner basis of $\sqrt{I}$, which applies to any ideal not in shape position. This is no coincidence, the probabilistic FGLM algorithm constructs a LEX Gröbner basis via the multiplication matrix for the last variable. Hence, for both methods the complexity is dominated by $\mathcal{O}\left(d^{\omega}\right)$.
3 Analysis of Ciminion
At round level Ciminion is a three branch Feistel network. In [Ste24, §4.1] a DRL Gröbner basis for Feistel-$2n/n$ polynomial systems has been computed by substituting the linear branches into the non-linear ones. In principle, we are going to replicate this strategy to construct a DRL Gröbner basis for Ciminion.
3.1 A DRL Gröbner Basis for Ciminion
For simplicity let us denote the matrix in the $i$th Ciminion round by $\mathbf{A}_{i}\in\mathbb{F}_{q}^{3\times 3}$, then
$\mathbf{A}_{i}^{-1}=\begin{pmatrix}0&1&-c_{4}^{(i)}\\ -1&0&1\\ 1&0&0\end{pmatrix},$ (28)
where $c_{4}^{(i)}\notin\mathbb{F}_{q}\setminus\{0,1\}$.
Theorem 3.1.
Let $\mathbb{F}_{q}$ be a finite field, and let
$\mathcal{F}_{\texttt{Ciminion}}=\{\mathbf{f}^{(1)},\dots,\mathbf{f}^{(r_{C}+r_{E})}\}\subset\mathbb{F}_{q}\Big{[}y_{1},y_{2},\mathbf{x}_{C}^{(1)},\allowbreak\dots,\mathbf{x}_{C}^{(r_{C}-1)},\allowbreak\mathbf{x}_{E}^{(1)},\dots,\mathbf{x}_{E}^{(r_{E})},x\Big{]}$
be a Ciminion polynomial system. For the DRL term order $y_{1}>y_{2}>\mathbf{x}_{C}^{(1)}>\ldots>\mathbf{x}_{C}^{(r_{C}-1)}>\mathbf{x}_{E}^{(1)}>\mathbf{x}_{E}^{(r_{E})}>x$, the following procedure computes a DRL Gröbner basis for Ciminion:
  1. (1)
    For every round $1\leq i\leq r_{C}+r_{E}$ compute $\mathbf{g}^{(i)}=\mathbf{A}_{i}^{-1}\mathbf{g}_{i}$.
  2. (2)
    Collect the linear polynomials from the $\mathbf{g}^{(i)}$’s in $\mathcal{L}$, and perform Gaussian elimination on $\mathcal{L}$.
  3. (3)
    Reduce the non-linear polynomials in the $\mathbf{g}^{(i)}$’s modulo $\mathcal{L}$.
Proof.
By inspection of the round function (Equation 9) we have that
$\mathbf{g}^{(1)}$ =($\ y_{1}\\ y_{2}+\aleph\cdot y_{1}\end{pmatrix}+\mathbf{A}_{1}^{-1}\left(\mathbf{c}^{(1)}-\mathbf{x}_{C}^{(1)}\right),$ (29)
$\mathbf{g}^{(i)}$ $=\begin{pmatrix}x_{C,1}^{(i-1)}\\[5.0pt] x_{C,2}^{(i-1)}\\[5.0pt] x_{C,3}^{(i-1)}+x_{C,1}^{(i-1)}\cdot x_{C,2}^{(i-1)}\end{pmatrix}+\mathbf{A}_{i}^{-1}\left(\mathbf{c}^{(i)}-\begin{pmatrix}x_{C,1}^{(i)}\\ x_{C,2}^{(i)}\\ x_{C,3}^{(i)}\end{pmatrix}\right),$ (30)
$\mathbf{g}^{(r_{C}+r_{E})}$ $=\begin{pmatrix}x_{E,1}^{(r_{E})}\\[5.0pt] x_{E,2}^{(r_{E})}\\[5.0pt] x_{E,3}^{(r_{E})}+x_{E,1}^{(r_{E})}\cdot x_{C,2}^{(r_{E})}\end{pmatrix}+\mathbf{A}_{i}^{-1}\left(\mathbf{c}^{(r_{E})}-\begin{pmatrix}c_{1}-p_{1}\\ c_{2}-p_{2}\\ x\end{pmatrix}\right),$ (31)
and analog for the equations involving the $\mathbf{x}_{E}^{(i)}$’s. Hence, $\mathcal{L}$ contains $\mathbf{g}^{(1)}$ and the first two components of all remaining $\mathbf{g}^{(i)}$’s.
For Gaussian elimination on $\mathcal{L}$ let us start in the last round $i=r_{C}+r_{E}$. By construction and Equation 28 the linear polynomials with leading monomials $x_{E,1}^{(r_{E})}$, $x_{E,2}^{(r_{E})}$ contain a linear term in $x$ and a constant term. Moving to $i=r_{C}+r_{E}-1$ next, then the linear polynomials from $\mathbf{g}^{(r_{C}+r_{E}-1)}$ depend on $\mathbf{x}_{E}^{(r_{C}+r_{E}-1)}$ and $\mathbf{x}_{E}^{(r_{C}+r_{E})}$, via Gaussian elimination the variables $x_{E,1}^{(r_{E})}$ and $x_{E,2}^{(r_{E})}$ get eliminated. Thus, the linear polynomials with leading monomials in $x_{E,1}^{(r_{E}-1)}$, $x_{E,2}^{(r_{E}-1)}$ can only depend on $x_{E,3}^{(r_{E})}$ and $x$. Moreover, by Equation 28 the coefficients for $x_{E,3}^{(r_{E})}$ are non-zero. Via a downwards induction we can then conclude that the linear polynomials with leading monomials $x_{\ast,1}^{(i)}$, $x_{\ast,2}^{(i)}$ only depend on the variables $x_{\ast,3}^{(i+1)},\dots,x_{E,3}^{(r_{E})},x$ and that the coefficients of $x_{\ast,3}^{(i+1)}$ are non-zero. The first round needs special care since the first row of $\mathbf{A}_{1}^{-1}$ is non-zero on the variable $x_{C,2}^{(1)}$. (If it were zero on $x_{C,1}^{(1)}$ and $x_{C,2}^{(1)}$, the first polynomial in $\mathbf{g}^{(1)}$ would have leading monomial $x_{C,3}^{(1)}$ and the proof would be finished.) Expansion of Equation 29 yields
$\mathbf{g}^{(1)}=\begin{pmatrix}\aleph\\ y_{1}\\ y_{2}+\aleph\cdot y_{1}\end{pmatrix}+\mathbf{A}_{1}^{-1}\mathbf{c}_{1}-\begin{pmatrix}x_{C,2}^{(1)}-c_{4}^{(1)}\cdot x_{C,3}^{(1)}\\ -x_{C,1}^{(1)}+x_{C,3}^{(1)}\\ x_{C,1}^{(1)}\end{pmatrix}.$
Producing a polynomial with leading monomial $y_{2}$ is straight-forward, we simply have to take $g_{3}^{(1)}-\aleph\cdot g_{2}^{(1)}$. Note that there is second polynomial with leading monomial $x_{C,2}^{(1)}$ after inversion of the matrices
$g_{2}^{(2)}=x_{C,2}^{(1)}+\left(\mathbf{A}_{2}^{-1}\mathbf{c}^{(2)}\right)_{2}-\left(-x_{C,1}^{(2)}+x_{C,3}^{(2)}\right).$
Then
$g_{1}^{(1)}+g_{2}^{(2)}=\aleph+\left(\mathbf{A}_{1}^{-1}\mathbf{c}^{(1)}\right)_{1}+c_{4}^{(1)}\cdot x_{C,3}^{(1)}+\left(\mathbf{A}_{2}^{-1}\mathbf{c}^{(2)}\right)_{2}+x_{C,1}^{(2)}-x_{C,3}^{(2)},$
and by the choice of our DRL term order and $c_{4}^{(1)}\neq 0$ this polynomial has leading monomial $x_{C,3}^{(1)}$. Obviously, after Gaussian elimination of $\mathcal{L}$ the terms of $g_{1}^{(1)}+g_{2}^{(2)},g_{2}^{(1)},g_{3}^{(1)}-\aleph\cdot g_{2}^{(1)}$ after the leading monomials will also only depend on the variables $x_{C,3}^{(2)},\dots,x_{E,3}^{(r_{E})},x$.
Lastly, let us consider the reduction modulo $\mathcal{L}$. By our previous observations all linear equations are of the form
$y_{1}+\mathcal{A}_{1}\left(x_{C,3}^{(2)},\dots,x_{E,3}^{(r_{C}+r_{E}-1)},x\right)$ $=0,$
$y_{2}+\mathcal{A}_{2}\left(x_{C,3}^{(2)},\dots,x_{E,3}^{(r_{C}+r_{E}-1)},x\right)$ $=0,$
$x_{E,3}^{(1)}+\mathcal{A}_{3}\left(x_{C,3}^{(2)},\dots,x_{E,3}^{(r_{C}+r_{E}-1)},x\right)$ $=0,$
$x_{\ast,1}^{(i)}+c_{4}^{(i)}\cdot x_{\ast,3}^{(i+1)}+\ldots$ $=0,$
$x_{\ast,2}^{(i)}-x_{\ast,3}^{(i+1)}+\ldots$ $=0,$
$x_{E,1}^{(r_{E})}+c_{4}^{(r_{C}+r_{E})}\cdot x+\ldots$ $=0,$
$x_{E,2}^{(r_{E})}-x+\ldots$ $=0,$
where $\mathcal{A}_{1}$, $\mathcal{A}_{2}$ and $\mathcal{A}_{3}$ are affine polynomials. So, we can directly conclude from Equation 30 that
$\hat{g}_{3}^{(i)}=\left(g_{3}^{(i)}\mod(\mathcal{L})\right)\in\mathbb{F}_{q}\Big{[}x_{C,3}^{(2)},\dots,x_{E,3}^{(r_{C}+r_{E}-1)},x\Big{]},$
where $2\leq i\leq r_{C}+r_{E}$, has leading monomial ${x_{\ast,3}^{(i)}}^{2}$ and respectively $x^{2}$ for the last round.
Thus, we have constructed a polynomial system with pairwise coprime leading monomials, so by [CLO15, Chapter 2 §9 Theorem 3, Proposition 4] we have constructed a Gröbner basis. ∎
Note that if we only keep the $r_{E}+r_{C}-1$ quadratic polynomials of $\mathcal{G}$, then we still have a fully determined polynomial system in $r_{E}+r_{C}-1$ variables as well as a Gröbner basis. Therefore, we call the quadratic polynomials of $\mathcal{G}$ the downsized Ciminion polynomial system.
For the next corollary, recall that a LEX Gröbner basis =$\mathcal{F}=\{f_{1},\dots,f_{n}\}\subset K[x_{1},\dots,\allowbreak x_{n}]$ is said to be in $x_{n}$-shape position if it is of the form
$\begin{split}f_{1}&=x_{1}-\tilde{f}_{n}(x_{n}),\\ &\dots\\ f_{n-1}&=x_{n-1}-\tilde{f}_{n-1}(x_{n}),\\ f_{n}&=\tilde{f}_{n}(x_{n}).\end{split}$ (32)
Note that the LEX Gröbner basis of a radical ideal can be brought into shape position for some ordering of the variables, see [KR00, Theore 3.7.25].
Corollary 3.2.
In the situation of LABEL:Th:_Ciminion_Gr\"obner_basis, let $\mathcal{G}_{\texttt{Ciminion}}$ denote the Ciminion DRL Gröbner basis.
  1. (1)
    $\dim_{\mathbb{F}_{q}}\left(\mathcal{F}_{\texttt{Ciminion}}\right)=2^{r_{C}+r_{E}-1}$.
  2. (2)
    Let $I\subset\overline{\mathbb{F}_{q}}\Big{[}y_{1},y_{2},\mathbf{x}_{C}^{(1)},\allowbreak\dots,\mathbf{x}_{C}^{(r_{C}-1)},\allowbreak\mathbf{x}_{E}^{(1)},\dots,\mathbf{x}_{E}^{(r_{E})},x\Big{]}[x_{0}]$ be a homogeneous ideal such that $\mathcal{Z}_{+}(I)\neq\emptyset$ and $\mathcal{G}^{\text{hom}}\subset I$. Then $I$ is in generic coordinates.
  3. (3)
    $\operatorname{sd}_{DRL}\left(\mathcal{F}_{\texttt{Ciminion}}\right)\leq r_{C}+r_{E}$.
  4. (4)
    The LEX Gröbner basis for Ciminion under $y_{1}>y_{2}>\mathbf{x}_{C}^{(1)}>\ldots>\mathbf{x}_{C}^{(r_{C}-1)}>\mathbf{x}_{E}^{(1)}>\mathbf{x}_{E}^{(r_{E})}>x$ is in $x$-shape position.
Proof.
(1), (2) and (3) are consequences of the DRL Gröbner basis from LABEL:Th:_Ciminion_Gr\"obner_basis and application of Theorems 2.5 and 2.6. For the last claim we take a look at the shape of the DRL Gröbner basis
$g_{3}^{(2)}$ $=\mathcal{Q}_{2}\left(x_{C,3}^{(2)},\dots,x_{E,3}^{(i)},x\right)-\mathcal{A}_{2}\left(x_{C,3}^{(2)},\dots,x_{E,3}^{(r_{E})},x\right),$
$\dots$
$g_{3}^{(i)}$ $=\mathcal{Q}_{i}\left(x_{\ast,3}^{(i)},\dots,x_{E,3}^{(r_{E})},x\right)-\mathcal{A}_{i}\left(x_{\ast,3}^{(i-1)},\dots,x_{E,3}^{(r_{E})},x\right),$
$\dots$
$g_{3}^{(r_{C}+r_{E})}$ $=\mathcal{Q}\left(x\right)-\mathcal{A}_{r_{C}+r_{E}}\left(x_{E,3}^{(r_{E})},x\right),$
where the $\mathcal{Q}_{i}$’s are quadratic polynomials which contain the monomial ${x_{\ast,3}^{(i)}}^{2}$ and respectively $x^{2}$ for the last round, and the $\mathcal{A}_{i}$’s are affine polynomials that contain the monomial $x_{E,3}^{(i-1)}$ for all $i\geq 3$ and without any restrictions on $\mathcal{A}_{2}$. Then by iterative substitution of $g_{3}^{(r_{C}+r_{E})}$ into $g_{3}^{(r_{C}+r_{E}-1)}$, $g_{3}^{(r_{C}+r_{E}-1)}$ into $g_{3}^{(r_{C}+r_{E}-2)}$, etc. we produce a LEX Gröbner basis in $x$-shape position. ∎
3.2 Bariant’s Attack
Recently Bariant introduced an attack on Ciminion via univariate greatest common divisor computations [Bar23]. As in our polynomial model he introduced an auxiliary variable $X$ for the truncated component. If one inverts $p_{E}$ and $p_{C}$ using $(c_{1},c_{2},X)$ as output state one then obtains three polynomials, see [Bar23, Fig. 3],
$\aleph=f_{1}(X),\qquad K_{1}=f_{2}(X),\qquad K_{2}=f_{3}(X).$ (33)
Since the first polynomial is univariate in $X$ one can compute its greatest common divisor with $X^{q}-X$ to solve for $X$ and henceforth also solve for $K_{1}$ and $K_{2}$.
We note that this attack can also be interpreted in terms of term order conversion of our DRL Gröbner basis to LEX, since the univariate polynomial $\aleph=f_{1}(X)$ is simply the univariate polynomial of the Ciminion LEX Gröbner basis in $x$-shape position. Moreover, the construction of the univariate LEX polynomial can be effectively expressed in the cost of univariate polynomial multiplication over finite fields [HvdH22] which is quasi-linear in the degree.
As countermeasure Bariant proposes to add a linear combination of $K_{1}$ and $K_{2}$ to each component after $p_{E}$. We note that this modification breaks our Gröbner basis from LABEL:Th:_Ciminion_Gr\"obner_basis. As settlement, we propose to add a linear combination of $K_{1}$ and $K_{2}$ only to the non-linear parts of the round functions, e.g.
$\begin{split}\mathcal{R}^{(l)}:\mathbb{F}_{q}^{3}\times\mathbb{F}_{q}^{2}&\to\mathbb{F}_{q}^{3},\\ \begin{pmatrix}x\\ y\\ z\end{pmatrix},(K_{1},K_{2})&\mapsto\begin{pmatrix}0&0&1\\ 1&c_{4}^{(l)}&c_{4}^{(l)}\\ 0&0&1\end{pmatrix}\begin{pmatrix}x\\ y\\ z+x\cdot y+\alpha\cdot K_{1}+\beta\cdot K_{2}\end{pmatrix}+\begin{pmatrix}c_{1}^{(l)}\\ c_{2}^{(l)}\\ c_{3}^{(l)}\end{pmatrix},\end{split}$ (34)
where $\alpha,\beta\in\mathbb{F}_{q}$ and $\alpha\cdot\beta\neq 0$. This modification does not affect the proof of LABEL:Th:_Ciminion_Gr\"obner_basis, hence we maintain the DRL Gröbner basis for the modified Ciminion polynomial system. However, this modification breaks the construction of the univariate LEX polynomial via univariate polynomial multiplication, because the DRL Gröbner basis polynomial for the $r_{C}$th round will now also contain linear terms for the variables $x_{C,3}^{(2)},\dots,x_{C,3}^{(r_{C}-2)}$ instead of only $x_{C,3}^{(r_{C}-1)},x_{E,3}^{(1)},\dots,x_{E,3}^{(r_{E})},x$. While the LEX Gröbner basis of the modified Ciminion might still be in $x$-shape position, we now can only construct it by running a FGLM algorithm.
Note if the modification of Equation 34 is performed in all rounds of $p_{E}$, then the Middle State-Output Relation polynomial system of Bariant et al. [BBLP22, §5] becomes underdetermined, so this attack connecting two outputs of two nonce samples would not be applicable anymore.
3.3 Implementation
We implemented Ciminion in the computer algebra systems SageMath [Sag24] and OSCAR [OSC24]. [^3] https://github.com/sca-research/Groebner-Basis-Cryptanalysis-of-Ciminion-and-Hydra.git In principle, the SageMath implementation can handle $128$ bit prime numbers, though we do not recommend using primes of that size for experimental Gröbner basis computations, because SageMath uses Python implementations for primes of that size. The OSCAR computer algebra system offers an efficient F4 implementation that can handle primes of size $q<2^{31}$, therefore we recommend to use OSCAR for experimental computations over small prime fields.
We implemented the Gröbner basis computation from LABEL:Th:_Ciminion_Gr\"obner_basis in SageMath. For practical parameters $p=2^{127}+45$, $r_{C}=90$ and $r_{E}=14$ a DRL Gröbner basis can be computed in negligible time.
3.3.1 Ciminion2
We also implemented a modified version of Ciminion to protect against Bariant’s attack, see Equation 34, called Ciminion2, which adds $K_{1}+K_{2}$ in the last round of $p_{C}$ and every round of $p_{E}$ to the non-linear component. I.e., the round function modifies to
$\mathcal{R}^{(l)}:\begin{pmatrix}x\\ y\\ z\end{pmatrix},\begin{pmatrix}K_{1}\\ K_{2}\end{pmatrix}\mapsto\begin{pmatrix}0&0&1\\ 1&c_{4}^{(l)}&c_{4}^{(l)}\\ 0&1&1\end{pmatrix}\begin{pmatrix}x\\ y\\ z+x\cdot y+K_{1}+K_{2}\end{pmatrix}+\begin{pmatrix}c_{1}^{(l)}\\ c_{2}^{(l)}\\ c_{3}^{(l)}\end{pmatrix},$ (35)
where $r_{C}\leq l\leq r_{C}+r_{E}$. For Ciminion2 a DRL Gröbner basis is also computed in negligible time in SageMath for practical parameters.
4 Analysis of Hydra
The Hydra heads are based on the Lai–Massey permutation. For the Gröbner basis analysis we recall that the Lai–Massey permutation
$\mathcal{L}:\begin{pmatrix}x\\ y\end{pmatrix}\mapsto\begin{pmatrix}x+F(x-y)\\ y+F(x-y)\end{pmatrix}$ (36)
is affine equivalent to the Feistel permutation
$\mathcal{F}:\begin{pmatrix}x\\ y\end{pmatrix}\mapsto\begin{pmatrix}x\\ y+F(x)\end{pmatrix}$ (37)
via the matrices $\mathbf{A}_{1}=\begin{pmatrix}1&-1\\ 0&1\end{pmatrix}$ and $\mathbf{A}_{2}=\begin{pmatrix}1&1\\ 0&1\\ \end{pmatrix}$, i.e. $\mathcal{L}=\mathbf{A}_{2}\circ\mathcal{F}\circ\mathbf{A}_{1}$. Inspired by this equivalence one can reduce the Hydra polynomial system to only one non-linear equation per round, i.e. the Hydra polynomial system transforms to a Feistel-like polynomial system.
In [Ste24, §6.2] the iterated polynomial systems representing block ciphers based on generalized Feistel networks were analyzed. For these polynomial systems linear systems were derived, and in case the systems are of full rank the iterated polynomial systems are in generic coordinates, see [Ste24, Theorem 6.5].
Based on the aforementioned affine equivalence, we are going to replicate this strategy to construct such a linear system for Hydra.
4.1 Generic Coordinates for Hydra
Recall that the Hydra iterated polynomial system (Definition 2.2) consists of $816\cdot r_{\mathcal{H}}+$ 8 many polynomials in $416\cdot r_{\mathcal{H}}+$ 4 many variables. Let
$\mathbf{A}_{n}=\begin{pmatrix}1&0&\ldots&0&-1\\ 0&1&\ldots&0&-1\\ &&\ddots&&\\ 0&0&\ldots&1&-1\\ 0&0&\ldots&0&1\end{pmatrix}\in\operatorname{GL}_{n}\left(\mathbb{F}_{q}\right),$ (38)
and let $\mathbf{B}=\operatorname{diag}\left(\mathbf{A}_{4},\mathbf{A}_{4}\right)\in\operatorname{GL}_{8}\left(\mathbb{F}_{q}\right)$. Then the first seven components of $\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{j}^{(i)}$, where $1\leq i\leq r_{\mathcal{H}}$ and $j\in\{1,2\}$, are affine and the eighth component is quadratic. Moreover, the forth and the eighth components of $\mathbf{B}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{f}_{\mathcal{R}}$ are quadratic while the other ones are affine. In particular, by extracting the $2\cdot 7\cdot r_{\mathcal{H}}$ affine components of the $\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{j}^{(i)}$’s we can eliminate the variables
$y_{1},\dots,y_{4},z_{1},\dots,z_{3},x_{j,1}^{(i)},\dots,x_{j,7}^{(i)},$ (39)
where $10\leq i\leq r_{\mathcal{H}}-$ 1 and $j\in\{1,2\}$, in the quadratic equations. (This can be explicitly seen in the proof of Theorem 4.1.) Thus, after the elimination we reduce to $82\cdot r_{\mathcal{H}}+$ 8 polynomials, of whom $6$ are also affine, in $42\cdot r_{\mathcal{H}}+$ 4 many variables. We call this polynomial system the downsized Hydra polynomial system.
Analog to the analysis of Feistel networks in [Ste24, §6.2], we now derive a linear system for Hydra with which we can verify whether the Hydra polynomial system is in generic coordinates.
Theorem 4.1.
Let $\mathbb{F}_{q}$ be a finite field, let $\overline{\mathbb{F}_{q}}$ be its algebraic closure, let
$\mathcal{F}_{\textsf{Hydra}}=\left\{\mathbf{f}_{1}^{(i)},\mathbf{f}_{\mathcal{R}},\mathbf{f}_{2}^{(i)}\right\}_{1\leq i\leq r_{\mathcal{R}}}\subset P=\overline{\mathbb{F}_{q}}\Big{[}\mathbf{y},\mathbf{z},\mathbf{x}_{1}^{(1)},\dots,\mathbf{x}_{1}^{(r_{\mathcal{H}}-1)},\mathbf{x}_{2}^{(0)},\dots,\mathbf{x}_{2}^{(r_{\mathcal{H}}-1)},\mathbf{k}\Big{]}$
be a Hydra polynomial system, and let
$\mathcal{G}_{\textsf{Hydra}}=\left\{\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{1}^{(i)},\mathbf{B}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{f}_{\mathcal{R}},\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{2}^{(i)}\right\}_{1\leq i\leq r_{\mathcal{H}}}.$
Then every homogeneous ideal $I\subset P[x_{0}]$ such that $\mathcal{Z}_{+}(I)\neq\emptyset$ and $\mathcal{G}_{\textsf{Hydra}}^{\text{hom}}\subset I$ is in generic coordinates if the following linear system has rank $416\cdot r_{\mathcal{H}}+$ 4
$y_{1}-z_{4}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{1}^{(1)}\Big{)}\bigg{)}_{1}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{1}^{(1)}-\mathbf{k}^{\prime}\Big{)}\bigg{)}_{8}$ $=0,$
$\vdots$
$z_{3}-z_{4}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{1}^{(1)}\Big{)}\bigg{)}_{7}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{1}^{(1)}-\mathbf{k}^{\prime}\Big{)}\bigg{)}_{8}$ $=0,$
$\sum_{l=1}^{4}y_{l}-z_{l}$ $=0,$
$x_{2,k}^{(0)}-x_{2,8}^{(0)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{2}^{(1)}\Big{)}\bigg{)}_{k}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{2}^{(1)}-\mathbf{k}^{\prime}\Big{)}\bigg{)}_{8}$ $=0,$
$\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{2,l}^{(0)}$ $=0,$
$x_{j,k}^{(i)}-x_{j,8}^{(i)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{j}^{(i+1)}\Big{)}\bigg{)}_{k}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{j}^{(i+1)}-\mathbf{k}\Big{)}\bigg{)}_{8}$ $=0,$
$\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{j,l}^{(i)}$ $=0,$
$x_{j,k}^{(r_{\mathcal{H}}-1)}-x_{j,8}^{(r_{\mathcal{H}}-1)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+(\mathbf{y},\mathbf{z})^{\intercal}\Big{)}\bigg{)}_{k}-\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+(\mathbf{y},\mathbf{z})^{\intercal}\Big{)}\bigg{)}_{8}$ $=0,$
$\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{j,l}^{(r_{\mathcal{H}}-1)}$ $=0,$
$x_{j,k}^{(r_{\mathcal{H}}-1)}-x_{j,8}^{(r_{\mathcal{H}}-1)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+\mathbf{x}_{2}^{(0)}\Big{)}\bigg{)}_{k}-\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+\mathbf{x}_{2}^{(0)}\Big{)}\bigg{)}_{8}$ $=0,$
$\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{j,l}^{(r_{\mathcal{H}}-1)}$ $=0,$
$y_{1}-y_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{1}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{8}$ $=0,$
$\vdots$
$y_{3}-y_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{3}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{8}$ $=0,$
$z_{1}-z_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{5}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{8}$ $=0,$
$\vdots$
$z_{3}-z_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{7}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{8}$ $=0,$
where $12\leq i\leq r_{\mathcal{H}}-$ 1, $j\in\{1,2\}$ and $1\leq k\leq 7$.
Proof.
By [Ste24, Theorem 3.2] we have to show that
$\sqrt{\mathcal{G}^{\text{top}}}=\left(\mathbf{y},\mathbf{z},\mathbf{x}_{j}^{(i)},\dots,\mathbf{x}_{1}^{(r_{\mathcal{H}}-1)},\mathbf{x}_{2}^{(0)},\dots,\mathbf{x}_{2}^{(r_{\mathcal{H}}-1)},\mathbf{k}\right)$
To shorten some writing let
$\mathbf{g}_{j}^{(i)}$ $=\left(\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{j}^{(i)}\right)^{\text{hom}}\mod x_{0},$
$\mathbf{g}_{\mathcal{R}}$ $=\left(\mathbf{B}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{f}_{\mathcal{R}}\right)^{\text{hom}}\mod x_{0}.$
Then for $i=1$ we have that
$\mathbf{g}_{1}^{(1)}$ =($=\begin{pmatrix}\left\{y_{k}-z_{4}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{1}^{(1)}\Big{)}\bigg{)}_{k}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{1}^{(1)}-\mathbf{k}^{\prime}\Big{)}\bigg{)}_{8}\right\}_{1\leq k\leq 7}\\[10.0pt] \left(\sum_{l=1}^{4}y_{l}-z_{l}\right)^{2}\end{pmatrix},$
$\mathbf{g}_{2}^{(1)}$ =($=\begin{pmatrix}\left\{x_{2,k}^{(0)}-x_{2,8}^{(0)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{2}^{(1)}\Big{)}\bigg{)}_{k}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{2}^{(1)}-\mathbf{k}^{\prime}\Big{)}\bigg{)}_{8}\right\}_{1\leq k\leq 7}\\[10.0pt] \left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{2,l}^{(0)}\right)^{2}\end{pmatrix}.$
For $12\leq i\leq r_{\mathcal{H}}-$ 1 and $j\in\{1,2\}$ we have that
$\mathbf{g}_{j}^{(i)}=\begin{pmatrix}\left\{x_{j,k}^{(i)}-x_{j,8}^{(i)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{j}^{(i+1)}\Big{)}\bigg{)}_{k}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{j}^{(i+1)}-\mathbf{k}\Big{)}\bigg{)}_{8}\right\}_{1\leq k\leq 7}\\[10.0pt] \left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{j,l}^{(i)}\right)^{2}\end{pmatrix}.$
For $i=r_{\mathcal{H}}$ we have that
$\mathbf{g}_{1}^{(r_{\mathcal{H}})}=$
($\begin{pmatrix}\left\{x_{1,k}^{(r_{\mathcal{H}}-1)}-x_{1,8}^{(r_{\mathcal{H}}-1)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+(\mathbf{y},\mathbf{z})^{\intercal}\Big{)}\bigg{)}_{k}-\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+(\mathbf{y},\mathbf{z})^{\intercal}\Big{)}\bigg{)}_{8}\right\}_{1\leq k\leq 7}\\[10.0pt] \left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{1,l}^{(r_{\mathcal{H}}-1)}\right)^{2}\end{pmatrix},$
$\mathbf{g}_{2}^{(r_{\mathcal{H}})}=$
($\begin{pmatrix}\left\{x_{2,k}^{(r_{\mathcal{H}}-1)}-x_{2,8}^{(r_{\mathcal{H}}-1)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+\mathbf{x}_{2}^{(0)}\Big{)}\bigg{)}_{k}-\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+\mathbf{x}_{2}^{(0)}\Big{)}\bigg{)}_{8}\right\}_{1\leq k\leq 7}\\[10.0pt] \left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{2,l}^{(r_{\mathcal{H}}-1)}\right)^{2}\end{pmatrix}.$
Finally, for the rolling function we have that
$\mathbf{B}\mathbf{g}_{\mathcal{R}}=\begin{pmatrix}\left\{y_{k}-y_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{k}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{4}\right\}_{1\leq k\leq 3}\\[10.0pt] g(\mathbf{y},\mathbf{z})\\ \left\{z_{k}-z_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{4+k}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{8}\right\}_{1\leq k\leq 3}\\[10.0pt] g(\mathbf{z},\mathbf{y})\end{pmatrix}.$
Except for the rolling function, all quadratic functions are powers of affine equations, so by Theorem 2.6 (2) we can replace the powers by the affine equations. This yields the linear system from the assertion. ∎
Naturally, this theorem can be extended to an arbitrary number of Hydra samples. It also follows by a simple contraposition that the downsized Hydra polynomial system in $42\cdot r_{\mathcal{H}}+$ 4 variables is in generic coordinates if the full polynomial system is in generic coordinates. Now it is straight-forward to compute the Macaulay bound (Theorem 2.5) for Hydra.
Corollary 4.2.
Let $\mathbb{F}_{q}$ be a finite field, let $\overline{\mathbb{F}_{q}}$ be its algebraic closure, and let
~$\widetilde{\mathcal{F}}_{\textsf{Hydra}}\subset\overline{\mathbb{F}_{q}}\Big{[}z_{4},x_{1,8}^{(1)},\dots,x_{1,8}^{(r_{\mathcal{H}}-1)},x_{2,8}^{(0)},\dots,x_{2,8}^{(r_{\mathcal{H}}-1)},\mathbf{k}\Big{]}$
be the downsized Hydra polynomial system. If ~$\widetilde{\mathcal{F}}_{\textsf{Hydra}}^{\text{hom}}$ is in generic coordinates, then
$\operatorname{sd}_{DRL}\left(\widetilde{\mathcal{F}}_{\textsf{Hydra}}\right)\leq 2\cdot r_{\mathcal{H}}+3.$
Remark 4.3.
This bound can be further improved by utilizing the $6$ affine equations from the rolling function. In case that these $6$ equations are linearly independent of each other and all other affine equations, we can eliminate further $6$ variables. Thus, the Hydra polynomial system reduces to $22\cdot r_{\mathcal{H}}+$ 2 quadratic polynomials in $22\cdot r_{\mathcal{H}}-$ 2 many variables. Then
$\operatorname{sd}_{DRL}\left(\widetilde{\mathcal{F}}_{\textsf{Hydra}}\right)\leq 2\cdot r_{\mathcal{H}}.$
The reason why we did not incorporate this improvement in Theorem 2.5 is that elimination via the affine equations from the rolling function depends non-trivially on $\mathbf{M}_{\mathcal{R}}$ and $\mathbf{M}_{\mathcal{J}}$. Thus, a generic treatment is not possible.
4.2 Extracting a Zero-Dimensional DRL Gröbner Basis for Hydra
Under the assumption of Remark 4.3, we can extract a zero-dimensional Gröbner basis from the downsized Hydra polynomial system via a linear change of coordinates. Then, $2\cdot r_{\mathcal{H}}$ quadratic polynomials stem from Hydra heads and the remaining two from the rolling function.
Recall that the quadratic terms in the heads are of the form
$g_{1}^{(1)}$ $=\left(\sum_{l=1}^{4}y_{l}-z_{l}\right)^{2},\qquad$ $g_{2}^{(1)}=\left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{2,l}^{(0)}\right)^{2},$ (40)
$g_{j}^{(i)}$ $=\left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{j,l}^{(i)}\right)^{2},$ $2\leq i\leq r_{\mathcal{H}}-1,\ j\in\{1,2\}$ (41)
$g_{1}^{(r_{\mathcal{H}})}$ $=\left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{1,l}^{(r_{\mathcal{H}}-1)}\right)^{2},\qquad$ $g_{2}^{(r_{\mathcal{H}})}=\left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{2,l}^{(r_{\mathcal{H}}-1)}\right)^{2},$ (42)
i.e. they are the squares of linear combinations of the variables. Therefore, we collect these linear combinations in a matrix $\mathbf{M}\in\mathbb{F}_{q}^{2\cdot r_{\mathcal{H}}\times(2\cdot r_{\mathcal{H}}-2)}$ analog to the linear system in Theorem 4.1 to verify generic coordinates for Hydra. If $\operatorname{rank}\left(\mathbf{M}\right)=2\cdot r_{\mathcal{H}}-2$, then we can perform a linear change of coordinates to new variables $\hat{x}_{1},\dots,\hat{x}_{2\cdot r_{\mathcal{H}}-2}$. After the change of coordinates the Hydra polynomial system contains $22\cdot r_{\mathcal{H}}-$ 2 polynomials of the form
$g_{i}=\hat{x}_{i}^{2}+\sum_{j=1}^{2\cdot r_{\mathcal{H}}-2}\alpha_{i,j}\cdot\hat{x}_{j}+\alpha_{i,0}.$ (43)
Trivially, $\mathcal{G}_{\textsf{Hydra}}=\{g_{i}\}_{1\leq i\leq 2\cdot r_{\mathcal{H}}-2}\subset\mathbb{F}_{q}\big{[}\hat{x}_{1},\dots,\hat{x}_{2\cdot r_{\mathcal{H}}-2}\big{]}$ forms a zero-dimensional DRL Gröbner basis, see [CLO15, Chapter 2 §9 Theorem 3, Proposition 4]. In particular, for the $\mathbb{F}_{q}$-vector space dimension of the quotient ring we then have that
$\dim_{\mathbb{F}_{q}}\left(\mathcal{G}_{\textsf{Hydra}}\right)=2^{2\cdot r_{\mathcal{H}}-2}.$ (44)
We also note that after an auxiliary division by remainder computation, the additional four polynomials are of the form
$h_{j}=\sum_{1\leq k<l\leq 2\cdot r_{\mathcal{H}}-2}\alpha_{i,k,l}\cdot\hat{x}_{k}\cdot\hat{x}_{l}+\sum_{k=1}^{2\cdot r_{\mathcal{H}}-2}\beta_{i,k}\cdot\hat{x}_{k}+\beta_{i,0}.$ (45)
4.3 Implementation
We have implemented Hydra and its iterated polynomial system in the computer algebra system SageMath [Sag24] and OSCAR [OSC24]. [^4] https://github.com/sca-research/Groebner-Basis-Cryptanalysis-of-Ciminion-and-Hydra.git In principle, the SageMath implementation can handle $128$ bit prime numbers, though we do not recommend using primes of that size for experimental Gröbner basis computations, because SageMath uses Python implementations for primes of that size. The OSCAR computer algebra system offers an efficient F4 implementation that can handle primes of size $q<2^{31}$, therefore we recommend to use OSCAR for experimental computations over small prime fields.
Moreover, we implemented the linear system from Theorem 4.1 in SageMath to verify whether a Hydra instance is in generic coordinates. We stress that our generic coordinates test can handle practical parameters like $128$ bit prime numbers. Also, our verification function can handle an arbitrary number of Hydra samples.
In case a Hydra polynomial system is in generic coordinates, we implemented the change of coordinates from Section 4.2 for two samples to extract a zero-dimensional DRL Gröbner basis from the Hydra polynomial system.
4.3.1 A Concrete Instance
In [GØSW22, Appendix C] the Hydra designers propose concrete matrices over the prime field $q=2^{127}+45$
$\mathbf{M}_{\mathcal{E}}$ $=\operatorname{circ}\left(3,2,1,1\right),$ (46)
$\mathbf{M}_{\mathcal{I}}$ $=\begin{pmatrix}1&1&1&1\\ 1&4&1&1\\ 3&1&3&1\\ 4&1&1&2\end{pmatrix},$ (47)
$\mathbf{M}_{\mathcal{J}}$ $=\begin{pmatrix}3&1&1&1&1&1&1&1\\ 7&3&1&1&1&1&1&1\\ 4&1&4&1&1&1&1&1\\ 3&1&1&8&1&1&1&1\\ 7&1&1&1&7&1&1&1\\ 8&1&1&1&1&5&1&1\\ 5&1&1&1&1&1&2&1\\ 4&1&1&1&1&1&1&6\end{pmatrix}.$ (48)
For security level $\kappa=128$ the designers suggest to use $r_{\mathcal{H}}=39$, see [GØSW23, §5.5].
For this instance using our SageMath implementation we verified that the Hydra polynomial system is in generic coordinates for $m\in\{2,3,4,5,10\}$ many samples. For $m=2$ we can compute a Hydra DRL Gröbner basis with the change of coordinates from Section 4.2. In particular, the polynomial system then consists of $22\cdot r_{\mathcal{H}}+$ 2 polynomials in $22\cdot r_{\mathcal{H}}-$ 2 variables, so we can use the slightly improved solving degree bound from Remark 4.3 to estimate the complexity of DRL Gröbner basis computations. On an Intel Core i7-10700 with 64 GB RAM verification of generic coordinates for $m=2$ needs $\leq 10\,\text{s}$ and for $m=10$ it needs $\leq 4\,\text{min}$.
In addition, for $r_{\mathcal{H}}=39$, to extract a Gröbner basis let $\mathbf{M}\in\mathbb{F}_{q}^{2\cdot r_{\mathcal{H}}\times(2\cdot r_{\mathcal{H}}-2)}$ collect the linear combinations in the quadratic terms of the downsized Hydra polynomial system. Then, the rows with indices $3,\dots,2\cdot r_{\mathcal{H}}$ are linearly independent. We also note that we did not find a round number $r_{\mathcal{H}}\geq 2$ for which the Gröbner basis extraction fails.
5 Shaving-Off Bits with the Eigenvalue Method
Let =$\mathcal{F}=\{f_{1},\dots,f_{n}\}\subset\mathbb{F}_{q}[x_{1},\dots,x_{n}]$ be such that $\operatorname{LM}_{DRL}\left(f_{i}\right)=x_{i}^{2}$ for all $i$, then $\mathcal{F}$ is a zero-dimensional Gröbner basis. Moreover, for Ciminion and Hydra we found Gröbner bases that satisfy this assumption. In Section 2.5 we already discussed that one can solve for the zero-locus of $\mathcal{F}$ by computing the eigenvalues of the multiplication matrices while maintaining the complexity estimate of the probabilistic FGLM algorithm [FGHR14]. We will now discuss that we can slightly improve the complexity estimate by taking the multiplicative structure of $\mathbb{F}_{q}[x_{1},\dots,x_{n}]/\left(\mathcal{F}\right)$ into account.
Recall that a $\mathbb{F}_{q}$-vector space basis of $\mathbb{F}_{q}[x_{1},\dots,x_{n}]/\left(\mathcal{F}\right)$ is given by the monomials not contained in $\operatorname{in}_{DRL}\left(\mathcal{F}\right)=\left(x_{1}^{2},\dots,x_{n}^{2}\right)$. Let $\mathcal{B}^{\prime}$ be the monomials not contained in $\left(x_{1}^{2},\dots,x_{n-1}^{2}\right)\subset\mathbb{F}_{q}[x_{1},\dots,x_{n-1}]$, then we sort the quotient space basis as =$\mathcal{B}=\left\{\mathcal{B}^{\prime},x_{n}\cdot\mathcal{B}^{\prime}\right\}$. The multiplication matrix for $x_{n}$ is then of the form
[…]
where $\mathbf{A},\mathbf{B}\in\mathbb{F}_{q}^{2^{n-1}\times 2^{n-1}}$. By definition, the characteristic polynomial is given by the determinant
$\chi_{x_{n}}(x)=\det\left(\mathbf{I}_{2^{n}\times 2^{n}}\cdot x-\mathbf{M}_{x_{n}}\right)=\det\begin{pmatrix}\mathbf{I}_{2^{n-1}\times 2^{n-1}}\cdot x&-\mathbf{I}_{2^{n-1}\times 2^{n-1}}\\ \mathbf{A}&\mathbf{I}_{2^{n-1}\times 2^{n-1}}\cdot x-\mathbf{B}\end{pmatrix}.$ (50)
For the LU-decomposition from Equation 24, we lift the determinant computation from the polynomial ring $\mathbb{F}_{q}[x]$ to its fraction field $\mathbb{F}_{q}(x)$, i.e. we treat $x$ as invertible. Then with Equation 25
$\chi_{x_{n}}(x)$ $=\det\begin{pmatrix}\mathbf{I}_{2^{n-1}\times 2^{n-1}}\cdot x&-\mathbf{I}_{2^{n-1}\times 2^{n-1}}\\ -\mathbf{A}&\mathbf{I}_{2^{n-1}\times 2^{n-1}}\cdot x-\mathbf{B}\end{pmatrix}$ (51)
$=\det\left(\mathbf{I}_{2^{n-1}\times 2^{n-1}}\cdot x\right)\cdot\det\left(\mathbf{I}_{2^{n-1}\times 2^{n-1}}\cdot x-\mathbf{B}-\mathbf{A}\cdot x^{-1}\right)$ (52)
$=x^{2^{n-1}}\cdot\det\left(\mathbf{I}_{2^{n-1}\times 2^{n-1}}\cdot x-\mathbf{B}-\mathbf{A}\cdot x^{-1}\right)$ (53)
$=\det\left(\mathbf{I}_{2^{n-1}\times 2^{n-1}}\cdot x^{2}-\mathbf{B}\cdot x-\mathbf{A}\right),$ (54)
where the latter equality follows from the fact that the determinant is a homogeneous function. We note that a similar determinant trick has been applied in [BBL+24, §3.2] to improve characteristic polynomial computations for specially shaped Gröbner bases.
For the next variable $x_{n-1}$, we extract a $\mathbb{F}_{q}$-valued solution for $x_{n}$ via the GCD method, substitute it into $\mathcal{F}\setminus\left\{f_{n}\right\}$, and repeat the trick for the multiplication matrix, see Equation 49, with one variable less.
Let $D(d,q)$ denote the complexity of the GCD with the field equation, see Equation 1. Provided that every characteristic polynomial has at most $N$ roots in $\mathbb{F}_{q}$, then the complexity of the Eigenvalue method with standard determinant complexity is given by
$\mathcal{O}\Bigg{(}\sum_{i=1}^{n}N^{n-i}\cdot\Big{(}2^{\omega\cdot i}+D\big{(}2^{i},q\big{)}\Big{)}\Bigg{)}=\mathcal{O}\Bigg{(}2^{\omega}\cdot\frac{2^{\omega\cdot n}-N^{n}}{2^{\omega}-N}+\sum_{i=1}^{n}N^{n-i}\cdot D\big{(}2^{i},q\big{)}\bigg{)}\Bigg{)}.$ (55)
With the determinant trick from Equation 54 the complexity for the Eigenvalue Method then improves to
$\mathcal{O}\Bigg{(}\sum_{i=1}^{n}N^{n-i}\cdot\Big{(}2^{\omega\cdot(i-1)}+D\big{(}2^{i},q\big{)}\Big{)}\Bigg{)}=\mathcal{O}\Bigg{(}\frac{2^{\omega\cdot n}-N^{n}}{2^{\omega}-N}+\sum_{i=1}^{n}N^{n-i}\cdot D\big{(}2^{i},q\big{)}\bigg{)}\Bigg{)}.$ (56)
In addition, for $N=1$, this implies by $\omega\geq 2$ that
$\log_{2}\left(\frac{2^{\omega\cdot n}-1}{2^{\omega}-1}\right)\leq\omega\cdot n-\log_{2}\left(3\right).$ (57)
I.e., we expect an improvement of at least $\log_{2}\left(3\right)\approx 1.5849$ compared to generic probabilistic term order conversion for an ideal in shape position.
By imposing more structure on $\mathcal{F}$ we can refine Equation 56 even further. Suppose that every $f_{i}\in\mathcal{F}$ is of the form
$f_{i}=x_{i}^{2}+\mathcal{A}_{i}(x_{1},\dots,x_{n}),$ (58)
where $\mathcal{A}_{i}$ is an affine polynomial. Assume we have found a solution for $x_{n}$, say $x_{n}=\alpha_{n}$, then we can rearrange
$\beta_{n-1}\cdot x_{n-1}=\alpha_{n}^{2}+\tilde{\mathcal{A}}_{i}(x_{1},\dots,x_{n-2},\alpha_{n}).$ (59)
(If $\beta_{n-1}=0$, then we simply rename the variables.) We can then further eliminate the variable $x_{n-1}$ in the polynomial system $\mathcal{F}\setminus\{f_{n-1},f_{n}\}\subset\mathbb{F}_{q}[x_{1},\dots,x_{n}]$ by substituting the previous affine equation for $x_{n-1}$ and $x_{n}=\alpha_{n}$. I.e., by finding one eigenvalue we can eliminate two variables. We also note that the Hydra DRL Gröbner basis (Section 4.2) satisfies the shape of Equation 58. Then, by iterating this variable elimination Equation 56 improves to
$\mathcal{O}\left(\begin{rcases}\begin{dcases}\sum_{i=1}^{\frac{n}{2}}N^{n-2\cdot i}\cdot\Big{(}2^{\omega\cdot(2\cdot i-1)}+D\big{(}2^{2\cdot i},q\big{)}\Big{)},&2\mid n,\\ \sum_{i=1}^{\frac{n-1}{2}}N^{n-1-2\cdot i}\cdot\Big{(}2^{\omega\cdot 2\cdot i}+D\big{(}2^{2\cdot i+1},q\big{)}\Big{)},&2\nmid n\end{dcases}\end{rcases}\right)$ (60)
$=\mathcal{O}\left(\begin{rcases}\begin{dcases}2^{\omega}\cdot\frac{2^{\omega\cdot n}-N^{n}}{2^{2\cdot\omega}-N^{2}}+\sum_{i=1}^{\frac{n}{2}}N^{n-2\cdot i}\cdot D\big{(}2^{2\cdot i},q\big{)},&2\mid n,\\ 2^{2\cdot\omega}\cdot\frac{2^{\omega\cdot(n-1)}-N^{n}}{2^{2\cdot\omega}-N^{2}}+\sum_{i=1}^{\frac{n-1}{2}}N^{n-1-2\cdot i}\cdot D\big{(}2^{2\cdot i+1},q\big{)},&2\nmid n\end{dcases}\end{rcases}\right),$ (61)
where we again assume that every characteristic polynomial has at most $N$ roots in $\mathbb{F}_{q}$.
Again for $N=1$, this implies that
$\log_{2}\left(\begin{rcases}\begin{dcases}2^{\omega}\cdot\frac{2^{\omega\cdot n}-1}{2^{2\cdot\omega}-1},&2\mid n,\\ 2^{2\cdot\omega}\cdot\frac{2^{\omega\cdot(n-1)}-1}{2^{2\cdot\omega}-1},&2\nmid n\end{dcases}\end{rcases}\right)$ $\leq\omega\left(n+1\right)-\log_{2}\left(2^{2\cdot\omega}-1\right)$ (62)
$\approx\begin{dcases}2\cdot n-1.9069,&\omega=2,\\ 3\cdot n-2.9773,&\omega=3.\end{dcases}$ (63)
I.e., we expect an improvement of $\approx\omega$ compared to generic probabilistic term order conversion.
6 Boolean Macaulay Matrices
Let $\mathcal{F}\subset\mathbb{F}_{2}[x_{1},\dots,x_{n}]$ be a polynomial system. If we are only interested in $\mathbb{F}_{2}$-valued solutions, then it is sufficient to compute a Gröbner basis of the ideal $\left(\mathcal{F},x_{1}^{2}-x_{1},\dots,x_{n}^{2}-x_{n}\right)$. On the other hand, via the field equations $F=\left(x_{1}^{2}-x_{1},\dots,x_{n}^{2}-x_{n}\right)$ we can reduce any monomial in the Macaulay matrix to a square-free monomial. Effectively, this reduces the full Macaulay matrix to a square-free submatrix, the so-called Boolean Macaulay matrix. Gaussian elimination for Boolean DRL Macaulay matrices over $\mathbb{F}_{2}$ was analyzed in [BFSS13] by Bardet et al.
Interestingly, for Ciminion & Hydra we found zero-dimensional DRL Gröbner bases with leading monomials $x_{i}^{2}$. Hence, for Gröbner basis computations we can use these bases to reduce the Macaulay matrix to a “Boolean” one over an arbitrary finite field $\mathbb{F}_{q}$. This motivates the following definitions.
Definition 6.1.
Let $K$ be a field, let $\mathcal{F}\subset P=K[x_{1},\dots,x_{n}]$, and let $>$ be a degree compatible term order on $P$. $\mathcal{F}$ is said to be a $K$-Boolean polynomial system if
  1. (i)
    for every $1\leq i\leq n$ there exists $f\in\mathcal{F}$ such that $\operatorname{LM}_{>}\left(f\right)=x_{i}^{2}$, and
  2. (ii)
    $\mathcal{F}$ is a $>$-Gröbner basis.
Definition 6.2 (cf. [BFSS13, Definition 1]).
Let $K$ be a field, let $\mathcal{F}_{\text{Bool}},\mathcal{F}\subset P=K[x_{1},\dots,x_{n}]$ be polynomial systems, and let $>$ be a degree compatible term order on $P$. Assume that $\mathcal{F}_{\text{Bool}}$ is $K$-Boolean with respect to $>$. For the (inhomogeneous) $K$-Boolean Macaulay matrix $M_{\leq d}^{\text{Bool}}$:
  • Columns are indexed by square-free monomials $s\in P\setminus\operatorname{in}_{>}\left(\mathcal{F}_{\text{Bool}}\right)$ with $\deg\left(m\right)\leq d$ sorted via $>$.
  • Rows are index by polynomials $t\cdot f$ where $t\in P\setminus\operatorname{in}_{>}\left(\mathcal{F}_{\text{Bool}}\right)$ is a square-free monomial and $f\in\mathcal{F}$ such that $\deg\left(t\cdot f\right)\leq d$.
  • In the row $t\cdot f$ one writes the coefficients of the remainder $r_{t\cdot f}=t\cdot f\mod\left(\mathcal{F}_{\text{Bool}}\right)$ with respect to $>$.
Note that any monomial $t\in P$ can be reduced to a square-free monomial $\hat{t}\equiv t\mod\left(\mathcal{F}_{\text{Bool}}\right)$ with $\deg\left(\hat{t}\right)\leq\deg\left(t\right)$ due to degree compatibility of $>$. Since remainders with respect to a Gröbner basis are unique, see [CLO15, Chapter 2 §6 Proposition 1], for the remainder of $t\cdot f$ it does not matter whether we reduce $t$ first and then $\hat{t}\cdot f$ or $t\cdot f$ at once. So it is indeed sufficient to consider only products with square-free monomials.
Obviously, the remainder $r_{t\cdot f}=t\cdot f\mod\left(\mathcal{F}_{\text{Bool}}\right)$ is polynomial consisting of square-free terms. Moreover, by elementary properties of multivariate division by remainder with respect to a degree compatible term order, see [CLO15, Chapter 2 §3 Theorem 3], we have that
$t\cdot f=\sum_{i=1}^{m}g_{i}\cdot f_{i}^{\text{Bool}}+r_{t\cdot f},$ (64)
where $\deg\left(g_{i}\cdot f_{i}^{\text{Bool}}\right),\deg\left(r_{t\cdot f}\right)\leq\deg\left(t\cdot f\right)$. I.e., the $K$-Boolean Macaulay matrix can be constructed via row operations from the Macaulay matrix $M_{\leq d}$ of $\mathcal{F}\cup\mathcal{F}_{\text{Bool}}$. Alternatively, we can obtain the $K$-Boolean Macaulay matrix by pivoting non-square-free monomials in $M_{\leq d}$ via rows coming from $\mathcal{F}_{\text{Bool}}$, and then performing an auxiliary Gaussian elimination step that stops after all redundant pivots coming from $\mathcal{F}_{\text{Bool}}$ have been eliminated.
Let $M_{\leq d}$ be a Macaulay matrix, we denote with $\operatorname{rowsp}\left(M_{\leq d}\right)$ the row space basis, i.e. the polynomial system obtained via Gaussian elimination on $M_{\leq d}$. Analog, we use this notation for Boolean Macaulay matrices.
Definition 6.3 (Boolean Solving Degree).
Let $K$ be a field, let $\mathcal{F}_{\text{Bool}},\mathcal{F}\subset P=K[x_{1},\dots,\allowbreak x_{n}]$ be polynomial systems, and let $>$ be a degree compatible term order on $P$. Assume that $\mathcal{F}_{\text{Bool}}$ is $K$-Boolean with respect to $>$. The Boolean solving degree of $\mathcal{F}\cup\mathcal{F}_{\text{Bool}}$ is the least degree $d$ such that $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)\cup\mathcal{F}_{\text{Bool}}$ contains a Gröbner basis with respect to $>$. We denote it by $\operatorname{sd}^{\text{Bool}}_{>}\left(\mathcal{F}\cup\mathcal{F}_{\text{Bool}}\right)$.
Note that $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)$ might not contain zero-dimensional Gröbner basis of the full system. A trivial example is =$\mathcal{F}=\left\{x_{1}^{2},x_{1}\cdot x_{2},x_{2}^{2}\right\}$. Therefore, to extract the Gröbner basis from $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)\cup\mathcal{F}_{\text{Bool}}$ we set up the ideal of $>$-leading terms and then compute its minimal generating set. With the minimal generating set we can then filter the Gröbner basis. (It is also obvious that all non-square-free monomials are generated by the leading terms of $\mathcal{F}_{\text{Bool}}$, hence it is indeed sufficient to just consider $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)\cup\mathcal{F}_{\text{Bool}}$.)
The next theorem follows straight-forward from $M_{\leq d}^{\text{Bool}}\subset M_{\leq d}$ (as $K$-vector spaces).
Theorem 6.4.
Let $K$ be a field, let $\mathcal{F}_{\text{Bool}},\mathcal{F}\subset P=K[x_{1},\dots,x_{n}]$ be polynomial systems, and let $>$ be a degree compatible term order on $P$. Assume that $\mathcal{F}_{\text{Bool}}$ is $K$-Boolean with respect to $>$. Then
  1. (1)
    $\operatorname{sd}^{\text{Bool}}_{>}\left(\mathcal{F}\cup\mathcal{F}_{\text{Bool}}\right)\leq\operatorname{sd}_{>}\left(\mathcal{F}\cup\mathcal{F}_{\text{Bool}}\right)$.
  2. (2)
    If $d=\operatorname{sd}^{\text{Bool}}_{>}\left(\mathcal{F}\cup\mathcal{F}_{\text{Bool}}\right)<n$, then Gaussian elimination on $M_{\leq d}^{\text{Bool}}$ requires
    $\mathcal{O}\left(\left|\mathcal{F}\right|\cdot\left(\sum_{i=1}^{d}\binom{n}{i}\right)^{\omega}\right)$
    field operations.
  3. (3)
    If $d=\operatorname{sd}_{>}\left(\mathcal{F}\cup\mathcal{F}_{\text{Bool}}\right)\geq n$, then Gaussian elimination on $M_{\leq d}^{\text{Bool}}$ requires at most
    $\mathcal{O}\left(\left|\mathcal{F}\right|\cdot 2^{\omega\cdot n}\right)$
    field operations to produce a $>$-Gröbner basis.
Proof.
For (1), we pivot non-square-free monomials via polynomials from $\mathcal{F}_{\text{Bool}}$, then we perform the auxiliary Gaussian elimination step that produces $M_{\leq d}^{\text{Bool}}$ as submatrix, and finally we perform Gaussian elimination on $M_{\leq d}^{\text{Bool}}$. Since Gaussian elimination does not depend on the order of reductions $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)\cup\mathcal{F}_{\text{Bool}}$ has to contain the $>$-Gröbner basis.
For (2), trivially in degree $i$ there exist $\binom{n}{i}$ square-free monomials, and every $f\in\mathcal{F}$ can contribute to at most $\binom{n}{i}$ rows. I.e., the number of columns is bounded by $D=\sum_{i=1}^{d}\binom{n}{i}$, and the number of rows by $\left|\mathcal{F}\right|\cdot D$. The complexity of Gaussian elimination then follows from [Sto00, §2.2].
For (3), let $s\in P$ be a non-square-free monomial with $\deg\left(s\right)>n$, then $\hat{s}\equiv s\mod\left(\mathcal{F}_{\text{Bool}}\right)$ where $\hat{s}\in P$ is square-free, in particular $\deg\left(\hat{s}\right)\leq n$. Therefore, is suffices to perform Gaussian elimination on the full $K$-Boolean Macaulay matrix. Moreover, it is well known that $\sum_{i=0}^{n}\binom{n}{i}=2^{n}$. ∎
Remark 6.5.
This theorem can be refined by replacing the binomial coefficient $\binom{n}{i}$ by the number of monomials in $P\setminus\operatorname{in}_{>}\left(\mathcal{F}_{\text{Bool}}\right)$ of degree $i$.
So far our analysis does not differ from the one over $\mathbb{F}_{2}$, however the binary field has a special feature. Division by remainder with respect to $\left(x_{1}^{2}-x_{1},\dots,x_{n}^{2}-x_{n}\right)$ is “for free”, i.e. its complexity is negligible compared to Gaussian elimination. For general $K$-Boolean polynomial systems this is not necessarily the case. Luckily, the complexity of general division by remainder with respect to quadratic $K$-Boolean polynomial systems follows as corollary to [Ten19, Proposition 3.10].
Corollary 6.6.
Let $K$ be a field, let $\mathcal{F}_{\text{Bool}}\subset P=K[x_{1},\dots,x_{n}]$ be a quadratic polynomial system, let $f\in P$ be a polynomial, and let $>$ be a degree compatible term order on $P$. Assume that $\mathcal{F}_{\text{Bool}}$ is $K$-Boolean with respect to $>$. Then computation of $r_{f}=f\mod\left(\mathcal{F}_{\text{Bool}}\right)$ with respect to $>$ requires
$\mathcal{O}\left(\frac{n^{2}+3\cdot n+2}{2}\cdot\sum_{i=0}^{\deg\left(f\right)}\binom{n+i-1}{i}\right)$
$\in\mathcal{O}\left(\frac{n^{2}+3\cdot n+2}{2}\cdot\deg\left(f\right)\cdot\binom{n+\deg\left(f\right)-1}{\deg\left(f\right)}\right)$
field operations.
Proof.
There exist $\sum_{i=0}^{\deg\left(f\right)}\binom{n+i-1}{i}$ monomials of degree $\leq\deg\left(f\right)$ and $\binom{n+1}{2}+\binom{n}{1}+1$ monomials of degree $\leq 2$. Now the claim follows from [Ten19, Proposition 3.10]. ∎
Thus, the complexity of constructing the $K$-Boolean Macaulay matrix $M_{\leq d}^{\text{Bool}}$ with respect to quadratic $K$-Boolean system $\mathcal{F}_{\text{Bool}}$ is bounded by
$\mathcal{O}\left(\left|\mathcal{F}\right|\cdot\frac{n^{2}+3\cdot n+2}{2}\cdot\left(\sum_{i=0}^{d}\binom{n+i-1}{i}\right)\cdot\left(\sum_{i=0}^{d}\binom{n}{i}\right)\right).$ (65)
Moreover, we have the inequality
[…]
on M_≤d^Bool
[…]
on M_≤d
[…]
on M_≤d^Bool
[…]
on M_≤d
[…]
I.e. for $\omega=2$, if we ignore implied constants and prefactors, then we expect that construction of the $K$-Boolean Macaulay matrix dominates the Gröbner basis computation for general quadratic $K$-Boolean polynomial systems.
Of course, the $K$-Boolean Macaulay matrix generalizes to more advanced linear algebra-based Gröbner basis algorithms. Current state-of-the-art are Faugère’s F4 [Fau99] and Matrix-F5 [Fau02] algorithms. These algorithms perform Gaussian elimination on the Macaulay matrix $M_{\leq d}$ for increasing values of $d$. Moreover, F4 and F5 avoid redundant rows via efficient selection criteria. We illustrate a simplified variant of F4. Let $\mathcal{F}\subset P=K[x_{1},\dots,x_{n}]$ be a polynomial system, let $>$ be a term order, and let $M_{\leq d}$ be the Macaulay matrix for $\mathcal{F}$ in degree $d$. Then, we perform Gaussian elimination on $M_{\leq d}$ which yields the row space basis $\mathcal{G}_{1}$, now we set up a new Macaulay matrix $M_{\leq d}$ for $\mathcal{G}_{1}$ and again compute the row space basis $\mathcal{G}_{2}$ via Gaussian elimination. We iterate this procedure until the row space basis becomes stationary, i.e. $\mathcal{G}_{i}=\mathcal{G}_{i+1}$, then we verify whether $\mathcal{G}_{i}$ is already a $>$-Gröbner basis. If not then we pass to degree $d+1$ and repeat the procedure. Obviously, if the input system contains a $K$-Boolean polynomial system, then we can replace every full Macaulay matrix by its $K$-Boolean submatrix.
7 Cryptanalysis of Ciminion & Hydra
In this section we discuss how our results from the previous sections can be applied to cryptanalysis of Ciminion and Hydra.
7.1 Ciminion
For Gröbner basis cryptanalysis of Ciminion [DGGK21, § 4.4] the designers considered a weaker scheme where the keys are added after the permutations $p_{E}$ instead of before the rolling functions, see [DGGK21, Fig. 7]. The designer claim that the weaker scheme is simpler to analyze than Ciminion. We reject this claim for the iterated Ciminion polynomial system, after all we have computed a DRL Gröbner basis for Ciminion with rather simple affine variable eliminations, see LABEL:Th:_Ciminion_Gr\"obner_basis.
The designers assumed that weakened Ciminion polynomial systems are regular, moreover for regular systems it is well-known that the degree of regularity is equal to the Macaulay bound. Then, the DRL Gröbner basis complexity is bounded by [BFS04, Theorem 7]
$\mathcal{O}\left(m\cdot d_{\operatorname{reg}}\cdot\binom{n+d_{\operatorname{reg}}}{d_{\operatorname{reg}}}^{\omega}\right).$ (67)
For Ciminion cryptanalysis, it is then assumed that the DRL Gröbner basis computation is the most costly part in an attack. The Ciminion designers did not provide a dedicated analysis for the iterated polynomial model from Definition 2.1. They mostly analyze polynomial models where $p_{E}$ and $p_{C}$ or respectively $p_{E}\circ p_{C}$ are modeled without intermediate variables, i.e. they used a single polynomial vector for these permutations. The designers further claim that polynomial models with many intermediate variables do not improve upon the few variables but high degree polynomial models.
On the other hand, let $I,J\subset P$ be zero-dimensional ideals, then elementary we have that
$I\subset J\qquad\Rightarrow\qquad\mathcal{Z}\left(J\right)\subset\mathcal{Z}\left(I\right),$ (68)
and by the isomorphism theorem for rings
$I\subset J\qquad\Rightarrow\qquad\big{(}P/I\big{)}\big{/}(J/I)\cong P/J.$ (69)
In particular, the latter isomorphism also induces a surjection of finite dimensional $K$-vector spaces $P/I\twoheadrightarrow P/J$ which implies that $\dim_{K}\left(I\right)\geq\dim_{K}\left(J\right)$. In cryptographic terms this means that if $>$-Gröbner bases for $I$ and $J$ are known, then solving for the zero locus of $I$ is at least as hard as solving for the zero locus of $J$ if one uses only generic algorithms.
Back to Ciminion, the designers proposed to study the model in two variables for the keys for the full function, we call this model the fully substituted model. Trivially, we have that
(fully substituted Ciminion model)(iterated Ciminion $(\text{fully substituted {Ciminion} model})\subset(\text{iterated {Ciminion} model}).$ (70)
Since we have a DRL Gröbner basis for the iterated model at hand, we reject the claim by the Ciminion designers that introducing intermediate state variables does not improve upon the fully substituted model. So, the only non-trivial step in a fully determined Ciminion polynomial system is the construction of a univariate polynomial.
As outlined in Corollary 3.2 and Section 3.2, the univariate LEX polynomial can be constructed by iterated substitution of a univariate polynomial into a quadratic polynomial. By [HvdH22] the complexity of univariate polynomial multiplication over a finite field is
$\mathcal{O}\Big{(}d\cdot\log(q)\cdot\log\big{(}d\cdot\log(q)\big{)}\Big{)},$ (71)
where $d$ is the maximum of the degrees of the polynomials. Then, the complexity of constructing the univariate LEX polynomial of Ciminion is bounded by
$\mathcal{O}\Big{(}2^{r_{C}+r_{E}-2}\cdot\log(q)\cdot\log\big{(}2^{r_{C}+r_{E}-2}\cdot\log(q)\big{)}\Big{)}.$ (72)
In Table 3 we present the complexity of term order conversion via fast polynomial multiplication and the GCD with the field equation (Equation 1) for Ciminion. Since the complexities of polynomial multiplication and the GCD are quasi-linear in the degree, one requires more than $110$ rounds for standard Ciminion to achieve $128$ bits of security.
Recall that we proposed a slightly modified variant Ciminion2 (Section 3.3.1) with a key addition to the non-linear component in the last rounds of $p_{C}$ every round of $p_{E}$. As discussed in Section 3.2, this modification breaks the construction of the univariate LEX polynomial via fast polynomial multiplication. For Ciminion2 we therefore either have to do term order conversion to LEX or use the Eigenvalue method. For term order conversion via the probabilistic FGLM algorithm [FGHR14], since we cannot prove $x$-shape position we have to compute the LEX Gröbner basis for the radical ideal which has complexity
$\tilde{\mathcal{O}}\left((r_{C}+r_{E}-1)\cdot 2^{\omega\cdot(r_{C}+r_{E}-1)}\right).$
On the other hand, for Ciminion2 it suffices to solve for the key variables via the Eigenvalue method. Since the Ciminion2 DRL Gröbner basis satisfies the assumption of Section 5, the complexity of characteristic polynomial construction is
$\mathcal{O}\left(2\cdot 2^{\omega\cdot(r_{C}+r_{E}-2)}\right).$ (73)
In Table 3 we present the complexity of the Eigenvalue Method and the GCD with the field equation for Ciminion2. Since the Eigenvalue Method adds a factor $\omega$ (in logarithmic terms), more than $66$ rounds for Ciminion2 already achieve $128$ bits of security.
If we add additional polynomials to the Ciminion or Ciminion2 polynomial systems, then we can recompute the DRL Gröbner bases. It is clear from the Macaulay bound (Theorem 2.5) that a polynomial $f\in P$ with $\deg\left(f\right)\geq 2$ can only increase the Macaulay bound. On the other hand, if $\deg\left(f\right)=1$, then we can eliminate one variable for free. Note that this does not affect the property of being in generic coordinates, but we would have to recompute the Ciminion Macaulay bound (Corollary 3.2). For the former case, we use
$\mathcal{O}\left(\binom{2\cdot(r_{C}+r_{E}-1)}{r_{C}+r_{E}}^{\omega}\right)$ (74)
as baseline for the complexity of DRL Gröbner basis computations. Based on our proven DRL complexity estimate, more than $36$ rounds for Ciminion and Ciminion2 polynomial systems achieve $128$ bits of security.
Lastly, in Table 3 we provide an additional column to compare our estimations to the rational of the Ciminion designers. In [DGGK21, §4.4] it is proposed to model the Ciminion function via two polynomials in two variables, i.e. the fully substituted model, of degree $2^{r_{C}+r_{E}-1}$. Under the assumption that these polynomials are regular we can then use the binomial term in Equation 67 as baseline for DRL Gröbner basis computations
$\mathcal{O}\left(\binom{2+2\cdot\left(2^{r_{C}+r_{E}-1}-1\right)+1-1}{2\cdot\left(2^{r_{C}+r_{E}-1}-1\right)+1}^{\omega}\right)=\mathcal{O}\left(2^{\omega\cdot(r_{C}+r_{E})}\right),$ (75)
For completeness, this complexity should not be compared to Equation 74. Since our DRL Gröbner basis complexity assumes that we have a Ciminion DRL Gröbner basis at hand plus some additional information that can be modeled via a non-linear polynomial. The complexity of Equation 75 quantifies the complexity of finding a DRL Gröbner basis in the first place without any additional information. Now ignorance of the iterated polynomial model in [DGGK21] becomes even more damaging. After all, we have constructed an iterated DRL Gröbner basis in $\mathcal{O}\left(r_{C}+r_{E}\right)$ while for the fully substituted model we obtain an exponential complexity estimate for finding the DRL basis. Moreover, the strategies for direct polynomial system solving outperform the complexity of finding the DRL basis for the fully substituted model.
Table 3: Complexity estimations for various Gröbner basis attacks on Ciminion and Ciminion2 with $q=2^{127}+45$ and $\omega=2$.
Complexity (bits)
$r_{C}$ $r_{E}$ Fast Polynomial Multiplication Eigenvalue Method DRL Computation Iterated Model DRL Computation Fully Substituted Model [DGGK21, §4.4]
$30$ $6$ $49$ $68$ $134$ $72$
$60$ $6$ $80$ $128$ $253$ $132$
$90$ $6$ $111$ $188$ $372$ $192$
$90$ $14$ $119$ $204$ $404$ $208$
$90$ $23$ $128$ $222$ $440$ $226$
Also, recall that $r_{C}\in\left\{s+6,\left\lceil\frac{2\cdot(s+6)}{3}\right\rceil\right\}$, where $s$ is the security level, for two Ciminion instantiations, see Table 2. In particular, for Ciminion2 this already implies that for the quotient space dimension $\dim_{\mathbb{F}_{q}}\left(\mathcal{F}_{\texttt{Ciminion}}\right)\geq 2^{\frac{s}{2}}$. Therefore, round numbers based on Table 2 should be sufficient to achieve the security level in Gröbner basis attacks via the iterated polynomial system, provided of course a fix against Bariant’s attack [Bar23] is implemented.
7.2 Hydra
For Gröbner basis cryptanalysis of Hydra the designers considered the downsized iterated Hydra polynomial system, see [GØSW23, §7.4] and [GØSW22, Appendix I.2]. The designers assumed that the downsized system of $22\cdot r_{\mathcal{H}}+$ 2 equations in $22\cdot r_{\mathcal{H}}-$ 2 variables is semi-regular [Frö85, Par10]. Under the semi-regularity assumption one can extract the degree of regularity ([BFS04, Definition 4]) $d_{\operatorname{reg}}$ of the Hydra polynomial system from the Hilbert series [KR05, Chapter 5] which then yields the complexity estimate ([BFS04, Theorem 7])
$\mathcal{O}\left((2\cdot r_{\mathcal{H}}+2)\cdot d_{\operatorname{reg}}\cdot\binom{2\cdot r_{\mathcal{H}}-2+d_{\operatorname{reg}}}{d_{\operatorname{reg}}}^{\omega}\right)$ (76)
for DRL Gröbner basis computations. For $r_{\mathcal{H}}^{\ast}=29$ the designers estimate $D_{\operatorname{reg}}=23$, with $\omega=2$ this yields a complexity estimate of $\approx 2^{130.8}$ bits. Moreover, it is claimed that $r_{\mathcal{H}}=r_{\mathcal{H}}^{\ast}+2=31$ rounds are sufficient to provide $128$ bits of security against Gröbner basis attacks.
7.2.1 Term Order Conversion & Eigenvalue Method Complexity Estimate
In Section 4.2 we have outlined how one can extract a zero-dimensional DRL Gröbner basis from the Hydra polynomial system via a linear change of coordinates. We can directly use this Gröbner basis to perform term order conversion to LEX, or apply the dedicated Eigenvalue Method from Section 5 combined with $\mathbb{F}_{q}$-valued root extraction via the GCD. Since we could not prove that a LEX Hydra Gröbner basis is in shape position, we have to use the complexity for the probabilistic FGLM [FGHR14] of radical ideals which becomes
$\tilde{\mathcal{O}}\left((r_{\mathcal{H}}-1)\cdot 2^{\omega\cdot 2\cdot(r_{\mathcal{H}}-1)+1}\right).$ (77)
For the Eigenvalue Method, the Hydra DRL Gröbner basis satisfies the assumptions of Section 5, and in addition it satisfies the shape condition of Equation 58. Therefore, we can use the complexity estimate from Equation 61. Since we do not know the number of $\mathbb{F}_{q}$-valued eigenvalues for the Hydra multiplication matrices we opt for a better safe than sorry approach. We assume that the first eigenvalue that an adversary recovers for each variable leads to the true key, i.e. we assume $N=1$. Then, the Eigenvalue Method complexity becomes
$\mathcal{O}\left(2^{\omega}\cdot\frac{2^{\omega\cdot(2\cdot r_{\mathcal{H}}-2)}-1}{2^{2\cdot\omega}-1}+\sum_{i=1}^{r_{\mathcal{H}}-1}D\big{(}2^{2\cdot i},q\big{)}\right),$ (78)
where $D\big{(}2^{i},q\big{)}$ again denotes the complexity of the GCD with the field equation (Equation 1).
In Table 4 we have computed the complexities for term order conversion and efficient $\mathbb{F}_{q}$-valued root extraction via the GCD. From the formulae, we can directly see that for $\omega=2$ and $r_{\mathcal{H}}\geq 34$ the complexities will always exceed $128$ bits. On the other, hand for $r_{\mathcal{H}}^{\ast}=29$ term order conversion requires only $118$ bits, and for $r_{\mathcal{H}}=r_{\mathcal{H}}^{\ast}+2=31$ it requires $126$ bits. This invalidates the claim by the Hydra designers that $31$ rounds are sufficient to protect against Gröbner basis attacks. Moreover, term order conversion always outperforms DRL Gröbner basis computation under the semi-regularity assumption. Lastly, for the round numbers in Table 4 the Eigenvalue Method shaves-off at least $6$ additional bits compared to term order conversion.
Table 4: Complexity estimate of the Eigenvalue Method and term order conversion to LEX with efficient $\mathbb{F}_{q}$-valued root extraction for Hydra with $q=2^{127}+45$ and $\omega=2$.
Complexity (Bits)
$r_{\mathcal{H}}$ Eigenvalue Method Term Order Conversion Semi-Regular Estimate [GØSW23]
$20$ $75$ $82$ $89$
$29$ $111$ $118$ $131$
$31$ $119$ $126$ $142$
$33$ $127$ $135$ $153$
$34$ $131$ $139$ $159$
$39$ $151$ $159$ $183$
7.2.2 Proven DRL Complexity Estimate
Via Theorem 4.1 and our SageMath implementation (Section 4.3) we obtain the proven DRL solving degree estimate for the Hydra polynomial instance from Section 4.3.1
$\operatorname{sd}_{DRL}\left(\mathcal{F}_{\textsf{Hydra}}\right)\leq 2\cdot r_{\mathcal{H}}.$ (79)
Since the Hydra polynomial system contains a zero-dimensional DRL Gröbner basis after a linear change of coordinates (Section 4.2), we can reduce Gaussian elimination on the full Macaulay matrix to Gaussian elimination on the $\mathbb{F}_{q}$-Boolean Macaulay matrix. With Theorem 6.4 and Equation 65 the complexity of constructing the $\mathbb{F}_{q}$-Boolean Macaulay matrix and Gaussian elimination for Hydra becomes
$\mathcal{O}\left(\left(8\cdot r_{\mathcal{H}}^{2}-4\cdot r_{\mathcal{H}}\right)\cdot\left(\sum_{i=0}^{2\cdot r_{\mathcal{H}}-2}\binom{2\cdot r_{\mathcal{H}}+i-3}{i}\cdot\binom{2\cdot r_{\mathcal{H}}-2}{i}\right)+2^{\omega\cdot 2\cdot(r_{\mathcal{H}}-1)+2}\right).$ (80)
In Table 5 we evaluate our complexity bound.
Table 5: DRL Gröbner basis complexity estimate for Hydra via Equation 80 with $\omega=2$.
Complexity (Bits)
$r_{\mathcal{H}}$ Proven Estimate Semi-Regular Estimate [GØSW23]
$20$ $105$ $89$
$29$ $151$ $131$
$31$ $161$ $142$
$33$ $172$ $153$
$34$ $177$ $159$
$39$ $203$ $183$
Therefore, we have derived a computer aided proof for the complexity of Gröbner basis attacks on Hydra without resorting to unproven assumptions like semi-regularity. Our proven complexity estimate is significantly larger than the hypothesized one for Hydra, since the degree of regularity estimate is always less than the Macaulay bound.
7.2.3 Semi-Regular Complexity Estimate for the Boolean Macaulay Matrices
As already mentioned, in the Hydra cryptanalysis it is assumed that the polynomial system of $22\cdot r_{\mathcal{H}}+$ 2 equations in $22\cdot r_{\mathcal{H}}-$ 2 variables is semi-regular. If we assume that the substituted Hydra polynomial system that contains the DRL Gröbner basis (Section 4.2) is also semi-regular, then we can again improve from the generic semi-regular complexity estimate (Equation 76) by passing to the $\mathbb{F}_{q}$-Boolean Macaulay matrix. Again, we can then use Equation 65 to obtain the complexity of construction and Gaussian elimination on the $\mathbb{F}_{q}$-Boolean Macaulay matrices
$\mathcal{O}\left(\begin{aligned} \left(8\cdot r_{\mathcal{H}}^{2}-4\cdot r_{\mathcal{H}}\right)&\cdot\left(\sum_{i=0}^{d_{\operatorname{reg}}}\binom{2\cdot r_{\mathcal{H}}+i-3}{i}\cdot\binom{2\cdot r_{\mathcal{H}}-2}{i}\right)\\ &+4\cdot\left(\sum_{i=0}^{d_{\operatorname{reg}}}\binom{2\cdot r_{\mathcal{H}}-2}{i}\right)^{\omega}\end{aligned}\right).$ (81)
In Table 6 we evaluate this complexity estimate. In addition, we show the complexity just for Gaussian elimination on the $\mathbb{F}_{q}$-Boolean Macaulay matrices, i.e. the complexity under the assumption that the Boolean Macaulay matrices are constructed “for free”. If we account for the cost of constructing the Boolean Macaulay matrices, then this method performs worse than the optimized Eigenvalue Method. However, it still beats the cryptanalysis of the Hydra designers for larger round numbers. Though, if we just consider the Gaussian elimination step, then this method slightly improves upon the Eigenvalue method.
Table 6: DRL Gröbner basis complexity estimate for Hydra via Equation 80 with $\omega=2$.
Complexity (Bits)
$r_{\mathcal{H}}$ $d_{\operatorname{reg}}$ Boolean Gaussian Elimination Boolean Semi-Regular Estimate Semi-Regular Estimate [GØSW23]
$20$ $16$ $74$ $91$ $89$
$29$ $23$ $108$ $130$ $131$
$31$ $25$ $116$ $140$ $142$
$33$ $27$ $125$ $149$ $153$
$34$ $28$ $129$ $154$ $159$
$39$ $32$ $148$ $176$ $183$
Empirical F4 Working Degree.
The Hydra designers performed small scale Gröbner basis experiments for Hydra with the F4 [Fau99] algorithm, see [GØSW22, Table 4]. We replicated their experiment with our OSCAR implementation over the prime field $\mathbb{F}_{7741}$. We report our results in Table 7, but as input system we use the substituted system from Section 4.2 which contains a zero-dimensional DRL Gröbner basis. Our experiments have been performed on an AMD EPYC-Rome (48) CPU with 94 GB RAM. Like the Hydra designers we observed that the empirical F4 working degree meets the semi-regular degree of regularity estimate.
Table 7: Highest working degree and running times for F4 in small scale Hydra polynomial systems over the prime field $\mathbb{F}_{7741}$.
Our data Grassi et al.’s data [GØSW22, Table 4]
$r_{\mathcal{H}}$ $3$ $4$ $5$ $6$ $7$ $8$ $9$ $10$ $6$ $7$ $8$ $9$ $10$ $11$
Highest F4 working degree $3$ $3$ $4$ $5$ $5$ $6$ $7$ $8$ $5$ $5$ $6$ $7$ $8$ $8$
Total time (s) $0.01$ $0.01$ $0.01$ $0.03$ $0.21$ $2.65$ $66.85$ $2295.7$ $0.1$ $0.9$ $13.4$ $293.8$ $8799.5$ $229606.8$
Overall, we conclude that currently the most competitive algebraic attacks against Hydra directly apply polynomial system solving techniques to the zero-dimensional DRL Gröbner basis, see Table 4.
Acknowledgments
Matthias Steiner has been supported in part by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 725042) and the enCRYPTON project (grant agreement No. 101079319).

### v4

@@
 Ciminion and Hydra are two recently introduced symmetric key Pseudo-Random Functions for Multi-Party Computation applications.
-For efficiency both primitives utilize quadratic permutations at round level.
+For efficiency, both primitives utilize quadratic permutations at round level.
 Therefore, polynomial system solving-based attacks pose a serious threat to these primitives.
-For Ciminion we construct a quadratic degree reverse lexicographic (DRL) Gröbner basis for the iterated polynomial model via affine transformations.
-For Hydra we provide a computer-aided proof in SageMath that a quadratic DRL Gröbner basis is already contained within the iterated polynomial system for the Hydra heads after affine transformations and a linear change of coordinates.
-Our Ciminion DRL Gröbner basis simplifies cryptanalysis, since one does not need to impose genericity assumptions, like being regular or semi-regular, anymore to derive complexity estimates on key recovery attacks.
-In the Hydra proposal it was claimed that $r_{\mathcal{H}}=31$ rounds for the heads are sufficient to achieve $128$ bits of security against Gröbner basis attacks for key recovery.
-However, for $r_{\mathcal{H}}=31$ standard term order conversion to a lexicographic (LEX) Gröbner basis for our Hydra DRL Gröbner basis requires just $126$ bits.
-Moreover, via the Eigenvalue Method up to $r_{\mathcal{H}}=33$ rounds can be attacked below $128$ bits.
+For Ciminion, we construct a quadratic degree reverse lexicographic (DRL) Gröbner basis for the iterated polynomial model via linear transformations.
+With the Gröbner basis we can simplify cryptanalysis, as we no longer need to impose genericity assumptions to derive complexity estimates.
+For Hydra, with the help of a computer algebra program like SageMath we construct a DRL Gröbner basis for the iterated model via linear transformations and a linear change of coordinates.
+In the Hydra proposal it was claimed that $r_{\mathcal{H}}=31$ rounds are sufficient to provide $128$ bits of security against Gröbner basis attacks for an ideal adversary with $\omega=2$.
+However, via our Hydra Gröbner basis standard term order conversion to a lexicographic (LEX) Gröbner basis requires just $126$ bits with $\omega=2$.
+Moreover, using a dedicated polynomial system solving technique up to $r_{\mathcal{H}}=33$ rounds can be attacked below $128$ bits for an ideal adversary.
 keywords:
-Gröbner basis Eigenvalue Method Ciminion Hydra
+Gröbner basis Ciminion Hydra
 1 Introduction
@@
 Also, one may need to store intermediate results securely in a database.
-For efficiency, it is recommended to perform encryption via MPC-friendly Pseudo-Random Functions (PRF) [GRR+16] which utilize a secret-shared symmetric key.
-Moreover, one can utilize MPC-friendly PRFs to speed up MPC applications, e.g.
-[IKOS07, LTW13, CDG+17].
+For efficiency, it is recommended to perform encryption via an MPC-friendly Pseudo-Random Function (PRF) [GRR+16] which utilizes a secret-shared symmetric key.
 Compared to classical bit-based PRFs their MPC counterparts have two novel design criteria:
@@
   • They must admit a low number of multiplications for evaluation.
-Two such recently introduced PRFs are Ciminion [DGGK21] and Hydra [GØSW23].
+Two such recently introduced MPC-friendly PRFs are Ciminion [DGGK21] and Hydra [GØSW23].
 These constructions are based on the iteration of degree $2$ Feistel and Lai–Massey permutations.
-Due to the low degree at round level, both designs consider Gröbner basis attacks [ACG+19] as one of the most threatening attack vectors.
-Let us shortly review a Gröbner basis attack on an iterated design.
+Due to the low degree at round level, both designs consider Gröbner basis attacks as one of the most threatening attack vectors.
+In this paper we perform dedicated Gröbner basis cryptanalysis for both designs. Let us quickly sketch the main steps of a Gröbner basis attack on a symmetric design:
   1. (1)
-    Model the function with an iterated system of polynomials
-    $\mathcal{R}(\mathbf{x}_{i-1},\mathbf{y})-\mathbf{x}_{i}=\mathbf{0},$
-    where $\mathcal{R}:\mathbb{F}_{q}^{n}\to\mathbb{F}_{q}^{n}$ is a round function.
-The $\mathbf{x}_{i}$’s are intermediate state variables with $\mathbf{x}_{0}$ the input state and $\mathbf{x}_{r}$ the output state of the function and $\mathbf{y}$ the cryptographic secret variables.
+    Model the function with a system of polynomials.
+For Ciminion and Hydra we analyze an iterated polynomial model $\mathcal{R}(\mathbf{x}_{i-1},\mathbf{y})-\mathbf{x}_{i}=\mathbf{0}$, where $\mathcal{R}:\mathbb{F}_{q}^{n}\to\mathbb{F}_{q}^{n}$ is the round function, the $\mathbf{x}_{i}$’s are intermediate state variables with $\mathbf{x}_{0}$ the input state and $\mathbf{x}_{r}$ the output state of the function, and $\mathbf{y}$ the cryptographic secret variables.
   2. (2)
@@
   3. (3)
-    Convert to an elimination order, e.g. the lexicographic (LEX) term order.
+    Construct a univariate polynomial, e.g. via term order conversion to the lexicographic (LEX) term order or via the Eigenvalue Method.
   4. (4)
-    Solve the univariate equation.
-Let $I\subset P=\mathbb{F}_{q}[x_{1},\dots,x_{n}]$ be a zero-dimensional ideal, let $d=\dim_{\mathbb{F}_{q}}\left(P/I\right)$ denote the $\mathbb{F}_{q}$-vector space dimension of the quotient space, and assume that a DRL Gröbner basis for $I$ is known.
-The original FGLM algorithm [FGLM93] for term order conversion runs in $\mathcal{O}\left(n\cdot d^{3}\right)$.
-An improved probabilistic FGLM variant [FGHR14] runs in $\tilde{\mathcal{O}}\left(d^{\omega}\right)$, where $2.372862\leq\omega<$ 2.37286, if the ideal is in shape position, and in $\tilde{\mathcal{O}}\left(n\cdot d^{\omega}\right)$ if a LEX Gröbner basis of the radical ideal is computed.
-Further, a sparse linear algebra FGLM variant [FM17] achieves $\mathcal{O}\left(\frac{d^{2+\frac{n-1}{n}}}{\sqrt{n}}\right)$.
-To factor the univariate LEX polynomial efficiently Bariant et al.
-[BBLP22, §3.1] proposed to compute its GCD with the field equation.
-To improve from the generic GCD complexity they propose to perform an auxiliary division by remainder to reduce both polynomials to the same degree.
-Provided that just a view solutions stem from the base field $\mathbb{F}_{q}$, the complexity of $\mathbb{F}_{q}$-valued root extraction via the GCD is given by
-$\mathcal{O}\Big{(}d\cdot\log(d)\cdot\log\big{(}\log(d)\big{)}\cdot\big{(}\log(d)+\log(q)\big{)}\Big{)},$ (1)
-if $d\leq q$ else one has to replace the roles of $d$ and $q$ in the complexity estimate.
-To estimate the complexity of Gröbner basis computations the Ciminion designers assumed that their polynomial systems are regular [Eis95, Chapter 17] and the Hydra designers assumed that their polynomial systems are semi-regular [Frö85, Par10].
-Under these assumptions it is well-known that the complexity of Gröbner basis computations can be estimated via the degree of regularity [BFS04], but both designs lack a formal proof for their genericity assumptions.
-Recently, Caminata & Gorla [CG21] introduced the notion of being in generic coordinates.
-Essentially, if a polynomial system is in generic coordinates, then the complexity of DRL Gröbner basis computations can always be expressed in terms of the Macaulay bound.
-We also note that for a regular polynomial system the degree of regularity is equal to the Macaulay bound.
-In another recent work, Steiner [Ste24] analyzed several cryptographic polynomial systems of MiMC [AGR+16], Hades [GLR+20] and GMiMC [AGP+19] for being in generic coordinates.
-In particular, he derived an efficient criterion [Ste24, Theorem 3.2] to prove that a polynomial system is in generic coordinates.
-Moreover, in his analysis of GMiMC he revealed that iterated polynomial systems for Feistel polynomial systems can be verified to be in generic coordinates by computing the rank of an associated linear system.
-Our work can be seen as generalization of his work since Ciminion and Hydra essentially boil down to Feistel-like polynomial systems.
-1.1 Our Results
-For the iterated Ciminion polynomial system we prove in LABEL:Th:_Ciminion_Gr\"obner_basis that a DRL Gröbner basis can be constructed via affine transformations.
-This trivializes Step (2) in a Gröbner basis attack, and one can directly proceed with term order conversion or other polynomial system solving techniques to solve for the key.
-It is also worthwhile mentioning that Bariant [Bar23] recently described that the univariate LEX polynomial of Ciminion can be constructed via univariate polynomial multiplication.
-We review this construction in more detail in Section 3.2, most importantly his attack can be prevented with additional key additions in the rounds of Ciminion.
-Moreover, if the keys are only added to the non-linear components, then we can maintain the iterated DRL Ciminion Gröbner basis (LABEL:Th:_Ciminion_Gr\"obner_basis).
-We call this slightly modified variant Ciminion2 (Section 3.3.1).
-For Hydra, in Theorem 4.1 we derive a linear system to verify generic coordinates for the iterated Hydra polynomial system.
-In case the Hydra polynomial system is in generic coordinates, then we can also use the linear system to perform a change of coordinates and produce a quadratic polynomial system which contains a zero-dimensional DRL Gröbner basis.
-For Hydra this gives us to options to solve for the key.
-  1. (I)
-    We ignore the additional equations and just use the DRL Gröbner basis to proceed with polynomial system solving.
-  2. (II)
-    We exploit the quadratic DRL Gröbner basis to reduce the standard Macaulay matrix for the overdetermined Hydra polynomial system to a submatrix whose columns are only indexed by square-free monomials.
-For (I), to compute the solutions of our polynomial systems we take a different approach than the classical term order conversion to LEX.
-We propose to compute the zeros via the eigenvalues of the so-called multiplication matrices, see Section 2.5.
-Then, the complexity of the Eigenvalue Method is dominated by the complexity of determinant computations.
-We also note that the probabilistic FGLM algorithm [FM17] boils down to matrix multiplications with the multiplication matrices, hence the generic Eigenvalue Method and the probabilistic FGLM algorithm share the same complexity estimate up to multiplicative constants.
-In particular, in Section 5 we discuss how the multiplicative structure of quadratic DRL Gröbner bases can be exploited to improve the complexity of the determinant for the characteristic polynomial.
-Consequently, this will yield the best currently known attacks on Hydra.
-For (II), computing a DRL Gröbner basis for multivariate quadratic polynomial systems is an omnipresent problem in cryptography.
-However, for the full Hydra polynomial system we have an advantage over the general problem, because the system already contains a zero-dimensional DRL Gröbner basis.
-Therefore, to decrease the size of the Macaulay matrices in a Gröbner basis computation, we propose to perform an auxiliary division by remainder computation with respect to the Gröbner basis.
-Then, the remainders will only consist of square-free monomials, so we can reduce Gaussian elimination to a submatrix of the Macaulay matrix whose columns are only indexed by square-free monomials.
-We discuss this approach and its complexity in Section 6.
-We also note that the reduction to square-free Macaulay matrices for polynomial systems $\mathcal{F}\subset\mathbb{F}_{2}[x_{1},\dots,x_{n}]$ was already discussed by Bardet et al.
-[BFSS13].
-Over $\mathbb{F}_{2}$ one can always include the field equations $F=\left(x_{1}^{2}-x_{1},\dots,x_{n}^{2}-x_{n}\right)$ to restrict to the $\mathbb{F}_{2}$-valued solutions of $\mathcal{F}$.
-Obviously, the complexity of division by remainder with respect to $F$ is negligible compared to the cost of Gaussian elimination.
-This is not the case for our Hydra polynomial system, in particular a simple inequality reveals that for Hydra the auxiliary division by remainder step, i.e.
-the construction of square-free Macaulay matrices, dominates the complexity of DRL Gröbner basis computations (Equation 66).
-In addition, to highlight the analogy from polynomial system solving over $\mathbb{F}_{2}$, over an arbitrary field $K$ we call square-free Macaulay matrices $K$-Boolean Macaulay matrices.
-We implemented Ciminion and Hydra polynomial systems in the computer algebra systems SageMath [Sag24] and OSCAR [OSC24].
+    Factor the univariate polynomial.
+For Ciminion and Hydra the designers conjectured that Step (2) dominates the complexity of a Gröbner basis attack.
+In addition, to estimate the complexity of this step it was assumed that the polynomial systems are regular or semi-regular [BFS04].
+We on the other hand will construct DRL Gröbner bases for both designs via linear transformations and possibly a linear change of variables.
+Thus, we can directly proceed with Step (3).
+Typically, this step is also computationally cheaper than finding the DRL Gröbner basis for generic polynomial systems, which is the main reason why we yield better complexity estimations than the designers.
+1.1 Our Results & Relationship to Other Works
+The research goals of this work can be summarized in three bullet points:
+  • Increase the understanding of polynomial models for MPC-friendly designs.
+  • Provide clean cryptanalysis which avoids strong genericity assumptions.
+  • Improve the complexity estimations of Ciminion and Hydra from a designer’s point of view.
+Ciminion.
+The PRF Ciminion is an iterated Feistel design based on the Toffoli gate [Tof80].
+For the iterated Ciminion polynomial model we construct a DRL Gröbner basis via linear transformations, see Theorem 3.1.
+Hence, we trivialize the hardness assumption of Step (2).
+It is worthwhile mentioning that in a recent work Bariant [Bar23] constructed the univariate LEX polynomial of Ciminion via iterated univariate polynomial multiplications.
+The complexities of constructing and factoring the LEX polynomial are linear-logarithmic in its degree.
+On the other hand, constructing the univariate LEX polynomial via Gröbner basis methods scales at least quadratic-logarithmic in its degree.
+Therefore, to the best of our knowledge Bariant’s attack currently constitutes the most competitive attack on certain Ciminion parameters.
+We also review this attack in more detail in Section 3.2.
+There are two simple countermeasures against Bariant’s attack: increasing the number of rounds or additional key additions.
+For the latter, we also discuss in Section 3.2 how key additions can be performed while maintaining the Ciminion DRL Gröbner basis from Theorem 3.1.
+Additionally, in Section 3.3.1 we propose Ciminion2, a slightly modified variant which incorporates key additions that are compatible with Theorem 3.1.
+In Section 3.4 we discuss Gröbner basis cryptanalysis of Ciminion and Ciminion2 respectively.
+Due to our DRL Gröbner basis we yield a clean analysis as well as state-of-the-art complexity estimations for Ciminion.
+In particular, we improve upon the designer’s original analysis.
+In Table 1 we present the minimal Ciminion round numbers that achieve $128$ bits of security for an ideal adversary [^1] In our setting an ideal adversary has access to a hypothetical matrix multiplication algorithm which achieves the linear algebra constant $\omega=2$.
+against the designer’s estimation (“Fully Substituted Model”), Bariant’s attack and our dedicated Eigenvalue Method for Ciminion and Ciminion2.
+Table 1: Minimal number of rounds required for Ciminion and Ciminion2 to achieve at least $128$ bits of security against various Gröbner basis attacks.
+All computations are performed over the prime $q=2^{127}+45$ with $\omega=2$.
+Complexity (bits)
+$r_{C}+r_{E}$ Bariant’s Attack [Bar23] Eigenvalue Method Fully Substituted Model [DGGK21, §4.4]
+$33$ $46.67$ $63.09$ $130$
+$66$ $81.22$ $129.09$ $262$
+$112$ $128.47$ $221.09$ $446$
+Hydra.
+The heads of the Hydra can be considered as iterated Lai–Massey ciphers, where the key as well as the input are hidden from the adversary and only the output is released.
+In a recent work Steiner [Ste24] introduced the notion of polynomial systems in generic coordinates [CG21] to cryptographic polynomial models.
+In essence, if a polynomial system is in generic coordinates, then we have a proven complexity for the computation of the DRL Gröbner basis [CG21].
+In [Ste24, §6.2] the iterated polynomial models of generalized Feistel ciphers were analyzed for being in generic coordinates.
+For many instances of Feistel ciphers generic coordinates can be verified via the rank of an associated linear system.
+In Theorem 4.1 we will construct such a linear system for the iterated Hydra model.
+Hence, with the aid of a computer algebra program like SageMath [Sag24] we yield a computer-aided proof for the complexity of DRL Gröbner basis computations.
+Additionally, if the linear system for generic coordinates verification has full rank, then we can perform a linear change of coordinates to transform the Hydra polynomial model into a DRL Gröbner basis, see Section 4.2.
+Via the change of coordinates we produce a quadratic DRL Gröbner basis in $22\cdot r_{\mathcal{H}}-$ 2 polynomials and variables together with four additional quadratic polynomials.
+This leaves us with two possible routes to recover the key:
+  1. (A)
+    We can recompute the DRL Gröbner basis for $22\cdot r_{\mathcal{H}}+$ 2 quadratic polynomials.
+  2. (B)
+    We ignore the additional equations and just use the DRL Gröbner basis to construct a univariate polynomial.
+The first approach was also considered by the Hydra designers [GØSW23, §7.4], although without the linear transformation to the DRL Gröbner basis.
+Under the semi-regularity assumption [BFS04] the designers derived the degree of regularity, which determines the complexity of the computation.
+In principle, we can apply the same strategy, but with one important benefit: After the linear transformation we only have to consider square-free monomials in a DRL Gröbner basis computation.
+This effectively saves both time and space.
+In analogy to Gröbner basis computations over $\mathbb{F}_{2}$ we call this the Boolean Semi-Regular Gröbner basis attack.
+Alternatively, we can simply restrict to the DRL Gröbner basis and proceed with term order conversion via the probabilistic FGLM algorithm [FGHR14].
+It turns out that this approach invalidates the designers’ claim that $r_{\mathcal{H}}=31$ rounds cannot be attacked by an ideal adversary below $128$ bits.
+In addition, via a dedicated Eigenvalue Method an ideal adversary can attack up to $r_{\mathcal{H}}=33$ rounds.
+In Section 4.4 we discuss the complexity estimations for Hydra in more detail, a summary of the considered attacks is given in Table 2.
+Let $r_{\mathcal{H}}^{\ast}$ be the least round number which achieves $128$ bits of security against an ideal adversary.
+The Hydra designers require for the head round number that $r_{\mathcal{H}}=\left\lceil 1.25\cdot\max\left\{24,2+r_{\mathcal{H}}^{\ast}\right\}\right\rceil$ [GØSW23, §5.5].
+In the original analysis the designers found that $r_{\mathcal{H}}^{\ast}=29$ and $r_{\mathcal{H}}=39$, however our analysis yields $r_{\mathcal{H}}^{\ast}=34$ which implies the increase $r_{\mathcal{H}}=45$ to maintain the originally intended security margin.
+Table 2: Complexities of Gröbner basis attacks on Hydra which beat the designers’ analysis, and security margin reduction for full rounds.
+All computations are performed over the prime $q=2^{127}+45$ with $\omega=2$.
+Complexity (Bits)
+$r_{\mathcal{H}}$ Term Order Conversion Eigenvalue Method Boolean Semi-Regular Estimate Semi-Regular Estimate [GØSW23]
+$29$ $117.81$ $111.09$ $123.29$ $130.80$
+$31$ $125.91$ $119.09$ $135.91$ $141.77$
+$33$ $134.00$ $127.09$ $145.75$ $152.75$
+$39$ $158.25$ $151.09$ $172.66$ $182.22$
+Implementation.
+We implemented the Ciminion and Hydra polynomial systems and their DRL Gröbner bases in the computer algebra systems SageMath [Sag24] and OSCAR [OSC24].
 [^2] https://github.com/sca-research/Groebner-Basis-Cryptanalysis-of-Ciminion-and-Hydra.git In particular, for Hydra we have implemented the linear system for the generic coordinates verification as well as the linear change of coordinates to produce a quadratic DRL Gröbner basis.
-We stress that our SageMath implementations is able to handle practical parameters like $128$ bit prime numbers and full rounds Hydra with $r_{\mathcal{H}}=39$.
-Finally, in Section 7 we discuss how our results can be applied for cryptanalysis of Ciminion and Hydra.
-  • For Ciminion2, utilizing the DRL Gröbner basis the cost of solving for the first key pair is proportional to
-    $\mathcal{O}\left(d^{\omega\cdot(r_{C}+r_{E}-2)+1}\right).$ (2)
-    Let $64\leq s\leq\log_{2}\left(q\right)$ denote the target security level of Ciminion, then one requires that $r_{E}\in\left\{s+6,\left\lceil\frac{2\cdot(s+6)}{3}\right\rceil\right\}$ [DGGK21, Table 1].
-Since $\omega\geq 2$, this complexity will always exceed the target security level.
-  • For Hydra, extracting a zero-dimensional DRL Gröbner basis and solving via the Eigenvalue Method turns out to be the most competitive attack vector.
-In [GØSW23, §7.4] the Hydra designers claim that $r_{\mathcal{H}}=31$ rounds are sufficient to achieve at least $128$ bits of security against Gröbner basis attacks.
-Note that this claim is already invalidated if we perform term order conversion to LEX via the probabilistic FGLM algorithm.
-Via the Eigenvalue Method we can attack up to $r_{\mathcal{H}}=33$ rounds below $128$ bit complexity.
-However, Hydra with full rounds $r_{\mathcal{H}}=39$ achieves $128$ bits security against all strategies analyzed in this paper.
-Our complexities for various attack strategies on Hydra are summarized in Table 1.
-Table 1: Summary of Hydra complexity estimations for various Gröbner basis attacks, see Tables 4, 5 and 6. All computations are performed over the prime $q=2^{127}+45$ with $\omega=2$.
-$r_{\mathcal{H}}$ Eigenvalue Method Term Order Conversion Proven Solving Degree Boolean Semi-Regular Estimate Semi-Regular Estimate [GØSW23]
-$20$ $\mathbf{75}$ $82$ $105$ $91$ $89$
-$29$ $\mathbf{111}$ $118$ $151$ $130$ $131$
-$31$ $\mathbf{119}$ $126$ $161$ $140$ $142$
-$33$ $\mathbf{127}$ $135$ $172$ $149$ $153$
-$34$ $\mathbf{131}$ $139$ $177$ $154$ $159$
-$39$ $\mathbf{151}$ $159$ $203$ $176$ $183$
+We stress that our SageMath implementations are able to handle practical parameters like $128$ bit prime numbers and full rounds Hydra with $r_{\mathcal{H}}=39$.
+1.2 Organization of the Paper
+In Section 2 we introduce the technical requirements of this paper.
+Ciminion and Hydra are formally introduced in Sections 2.1 and 2.2 respectively.
+In Section 2.3 we quickly review Gröbner basis algorithms and the notion of generic coordinates.
+As preparation for Hydra cryptanalysis, in Section 2.4 we consider polynomial systems $\mathcal{F}\subset K[x_{1},\dots,x_{n}]$ which contain a subset $\mathcal{F}_{\text{Bool}}=\{g_{1},\dots,g_{n}\}\subset\mathcal{F}$ with $\operatorname{LM}_{DRL}\left(g_{i}\right)=x_{i}^{2}$ for all $1\leq i\leq n$.
+For such polynomial systems we can reduce Gaussian elimination on the Macaulay matrix $M_{\leq d}$ to a submatrix $M_{\leq d}^{\text{Bool}}$ which only represents the square-free monomials.
+The complexity of square-free Gaussian elimination is derived in Theorem 2.10, and the one of the auxiliary division by remainder step is given in Corollary 2.12.
+Finally, in Section 2.5 we develop a dedicated strategy, called the Eigenvalue Method, to find a $\mathbb{F}_{q}$-valued key guess via the quadratic DRL Gröbner bases for Ciminion and Hydra.
+In Section 3 we analyze the iterated Ciminion polynomial system.
+In particular, in Theorem 3.1 we construct a Ciminion DRL Gröbner basis via linear transformations.
+Bariant’s attack and compatibility of countermeasures with the DRL Gröbner basis are discussed in Section 3.2.
+Finally, in Section 3.4 we discuss Gröbner basis cryptanalysis of Ciminion.
+In Section 4 we analyze the iterated polynomial system for the Hydra heads.
+A linear system to verify generic coordinates for Hydra is developed in Theorem 4.1.
+In case the linear system has full rank we discuss in Section 4.2 that a linear change of coordinates yields a quadratic DRL Gröbner basis together with four additional quadratic polynomials.
+Note that we do not perform this change of coordinates by hand, instead we outsource it to our SageMath or OSCAR implementations.
+In Section 4.4 we finish this paper with Gröbner basis cryptanalysis for Hydra.
 2 Preliminaries
-We denote fields with $K$ and their algebraic closure by $\bar{K}$.
+We denote fields with $K$ and their algebraic closure by $\overline{K}$.
+The finite field with $q$ elements is denoted as $\mathbb{F}_{q}$.
 If the field and the number of variables are clear from context, then we abbreviate the polynomial ring as $P=K[x_{1},\dots,x_{n}]$.
-Matrices $\mathbf{M}\in k^{m\times n}$ are denoted with upper bold letters and vectors $\mathbf{v}\in\mathbb{F}_{q}^{n}$ by lower bold letters.
+Matrices $\mathbf{M}\in K^{m\times n}$ are denoted with upper bold letters and vectors $\mathbf{v}\in K^{n}$ by lower bold letters.
 Matrix-vector and matrix-matrix products are written as $\mathbf{M}\mathbf{v}$.
-For polynomial ideals $I\subset K[x_{1},\dots,x_{n}]$, we denote the zero locus, i.e. its set of zeros, over $\bar{k}$ as
-$\mathcal{Z}\left(I\right)=\left\{\mathbf{p}\in\bar{k}^{n}\mid f(\mathbf{p})=0,\ \forall f\in I\right\}\subset\mathbb{A}^{n}_{\bar{k}}.$ (3)
-For homogeneous ideals $I$ we denote the projective zero locus over $\bar{K}$ by $\mathcal{Z}_{+}\left(I\right)\subset\mathbb{P}^{n-1}_{\bar{k}}$.
 Let $f\in K[x_{1},\dots,x_{n}]$ be a polynomial, and let $x_{0}$ be an additional variable, the homogenization of $f$ with respect to $x_{0}$ is defined as
 $f^{\text{hom}}(x_{0},\dots,x_{n})=x_{0}^{\deg\left(f\right)}\cdot f\left(\frac{x_{1}}{x_{0}},\dots,\frac{x_{n}}{x_{0}}\right)\in K[x_{0},\dots,x_{n}].$ (1)
-Analog, we denote the homogenization of ideals $I^{\text{hom}}=\left\{f^{\text{hom}}\mid f\in I\right\}$ and of finite systems of polynomials $\mathcal{F}^{\text{hom}}=\left\{f_{1}^{\text{hom}},\dots,f_{m}^{\text{hom}}\right\}$.
-For a homogeneous ideal $I\subset P$ and an integer $d\geq 0$ we denote
-$I_{d}=\left\{f\in I\mid\deg\left(f\right)=d,\ f\text{ homogeneous}\right\}.$ (5)
-For a term order $>$ and an ideal $I\subset P$ we denote with $\operatorname{in}_{>}(I)=\{\operatorname{LT}_{>}(f)\mid f\in I\}$ the initial ideal of $I$, i.e.
-the ideal of leading terms of $I$, with respect to $>$.
-Every polynomial $f\in K[x_{1},\dots,x_{n}]$ of degree $d$ can be written as $f=f_{d}+f_{d-1}+\ldots+f_{0}$, where $f_{i}$ is homogeneous of degree $i$.
-We denote its highest degree component $f_{d}$ of $f$ with $f^{\text{top}}$, and analog we denote $\mathcal{F}^{\text{top}}=\left\{f_{1}^{\text{top}},\dots,f_{m}^{\text{top}}\right\}$.
-Note that
-$f^{\text{top}}=f^{\text{hom}}\mod(x_{0}).$ (6)
-Let $I,J\subset K[x_{1},\dots,x_{n}]$ be ideals, then we denote with
-$I:J=\left\{f\in K[x_{1},\dots,x_{n}]\mid\forall g\in J\colon f\cdot g\in I\right\}$ (7)
-the usual ideal quotient, and with
-$I:J^{\infty}=\bigcup_{i\geq 1}I:J^{i}$ (8)
-the saturation of $I$ with respect to $J$.
-Let $I,\mathfrak{m}\in K[x_{0},\dots,x_{n}]$ be homogeneous ideals where $\mathfrak{m}=(x_{0},\dots,x_{n})$, then we call $I^{\text{sat}}=I:\mathfrak{m}^{\infty}$ the saturation of $I$.
-Since we analyze vector-valued functions $F:\mathbb{F}_{q}^{n}\to\mathbb{F}_{q}^{n}$ we also denote vectors of polynomial ring variables as $\mathbf{x}=(x_{1},\dots,x_{n})^{\intercal}$.
-Moreover, for a term order $>$ on $P$, if not specified otherwise we always assume that the variables within a vector $\mathbf{x}$ are naturally ordered with respect to $>$, i.e.
-$x_{1}>\ldots>x_{n}$.
-We will often encounter the lexicographic and the degree reverse lexicographic term order which we will abbreviate as LEX and DRL respectively.
+Analogously, we denote the homogenization of ideals $I^{\text{hom}}=\left\{f^{\text{hom}}\mid f\in I\right\}$ and of finite systems of polynomials $\mathcal{F}^{\text{hom}}=\left\{f_{1}^{\text{hom}},\dots,f_{m}^{\text{hom}}\right\}$.
+Let $f\in K[x_{1},\dots,x_{n}]$ be a polynomial, we can always decompose it as sum $f=f_{d}+\ldots+f_{0}$, where $f_{i}$ is homogeneous of degree $i$.
+We call $f_{d}$ the homogeneous highest degree component of $f$ and denote it as $f^{\text{top}}$.
+Analogously, we denote $\mathcal{F}^{\text{top}}=\left\{f_{1}^{\text{top}},\dots,f_{m}^{\text{top}}\right\}$.
+Note that we have $f^{\text{top}}=f^{\text{hom}}\mod\left(x_{0}\right)$.
+Let $I\subset R$ be an ideal in a commutative ring, the radical of $I$ is defined as $\sqrt{I}=\left\{f\in R\mid\exists n\in\mathbb{Z}_{\geq 1}\!:f^{n}\in I\right\}$.
+It is well-known that the radical is also an ideal.
 2.1 Ciminion
@@
 The permutations $p_{C}$ and $p_{E}$ are based on an iterated Feistel network
-$\begin{split}\mathcal{R}^{(l)}:\mathbb{F}_{q}^{3}&\to\mathbb{F}_{q}^{3},\\ \begin{pmatrix}x\\ y\\ z\end{pmatrix}&\mapsto\begin{pmatrix}0&0&1\\ 1&c_{4}^{(l)}&c_{4}^{(l)}\\ 0&1&1\end{pmatrix}\begin{pmatrix}x\\ y\\ z+x\cdot y\end{pmatrix}+\begin{pmatrix}c_{1}^{(l)}\\ c_{2}^{(l)}\\ c_{3}^{(l)}\end{pmatrix},\end{split}$ (9)
-where $c_{1}^{(l)},c_{2}^{(l)},c_{3}^{(l)}\in\mathbb{F}_{q}$ and $c_{4}^{(l)}\in\mathbb{F}_{q}\setminus\{0,1\}$.
+$\begin{split}\mathcal{R}^{(i)}:\mathbb{F}_{q}^{3}&\to\mathbb{F}_{q}^{3},\\ \begin{pmatrix}x\\ y\\ z\end{pmatrix}&\mapsto\begin{pmatrix}0&0&1\\ 1&c_{4}^{(i)}&c_{4}^{(i)}\\ 0&1&1\end{pmatrix}\begin{pmatrix}x\\ y\\ z+x\cdot y\end{pmatrix}+\begin{pmatrix}c_{1}^{(i)}\\ c_{2}^{(i)}\\ c_{3}^{(i)}\end{pmatrix},\end{split}$ (2)
+where $c_{1}^{(i)},c_{2}^{(i)},c_{3}^{(i)}\in\mathbb{F}_{q}$ and $c_{4}^{(i)}\in\mathbb{F}_{q}\setminus\{0,1\}$.
 Note that the Feistel network $(x,y,z)^{\intercal}\mapsto(x,y,z+x\cdot y)^{\intercal}$ is also known as Toffoli gate which plays a special role in reversible computing [Tof80].
-Figure 1: Encryption with Ciminion over a finite field $\mathbb{F}_{q}$.
-The permutation $p_{C}$ is defined as the $r_{C}$-fold composition of $\mathcal{R}^{(l)}$, and the function $p_{E}$ is defined as the $r_{E}$-fold composition of $\mathcal{R}^{(l)}$.
 Finally, the rolling function is defined as
-$\begin{split}rol:\mathbb{F}_{q}^{3}&\to\mathbb{F}_{q}^{3},\\ \begin{pmatrix}x\\ y\\ z\end{pmatrix}&\mapsto\begin{pmatrix}z+x\cdot y\\ x\\ y\end{pmatrix}.\end{split}$ (10)
-Formulae to compute the round numbers for Ciminion are given in Table 2.
-Table 2: Proposed round numbers for Ciminion [DGGK21, Table 1]. The security level $s$ must satisfy $64\leq s\leq\log_{2}\left(q\right)$, and $q\geq 2^{64}$.
-Instance $p_{C}$ $p_{E}$
-Standard $s+6$ $\max\left\{\left\lceil\frac{s+37}{12}\right\rceil,6\right\}$
-Data limit $2^{\frac{s}{2}}$ elements $\frac{2\cdot(s+6)}{3}$ $\max\left\{\left\lceil\frac{s+37}{12}\right\rceil,6\right\}$
-Conservative $s+6$ $\max\left\{\left\lceil\frac{3}{2}\cdot\frac{s+37}{12}\right\rceil,9\right\}$
-E.g., for $s=128$ and data limit $2^{\frac{s}{2}}$ we have the round numbers $r_{C}=90$ and $r_{E}=14$.
-It is now straight-forward to set up the iterated polynomial system for the first key pair $(K_{1},K_{2})$ given the first plain/ciphertext pair.
-For ease of writing we denote the $i$th intermediate state variables as $\mathbf{x}_{j}^{(i)}=\left(x_{j,1}^{(i)},x_{j,2}^{(i)},x_{j,3}^{(i)}\right)^{\intercal}$ where $j\in\{C,E\}$.
+$rol:\mathbb{F}_{q}^{3}\to\mathbb{F}_{q}^{3},\begin{pmatrix}x\\ y\\ z\end{pmatrix}\mapsto\begin{pmatrix}z+x\cdot y\\ x\\ y\end{pmatrix}.$ (3)
+Figure 1: Encryption with Ciminion.
+Formulas to compute the round numbers for Ciminion are given in [DGGK21, Table 1]. E.g., for $s=128$ and data limit $2^{\frac{s}{2}}$ we have the round numbers $r_{C}=90$ and $r_{E}=14$.
+It is now straight-forward to set up the iterated polynomial system for the first key pair $(K_{1},K_{2})$ given the first plain/ciphertext pair. For ease of notation we abbreviate $r=r_{C}+r_{E}$.
 Definition 2.1 (Iterated polynomial system for Ciminion).
 Let $\mathbb{F}_{q}$ be a finite field, and let $\big{(}\aleph,(p_{1},p_{2}),(c_{1},c_{2})\big{)}\in\mathbb{F}_{q}\times\mathbb{F}_{q}^{2}\times\mathbb{F}_{q}^{2}$ be a nonce/plain/ciphertext sample given by a Ciminion encryption function.
-We call the polynomial system $\mathcal{F}_{\texttt{Ciminion}}=\left\{\mathbf{f}^{(i)}\right\}_{1\leq i\leq e_{C}+r_{E}}\subset\mathbb{F}_{q}\Big{[}y_{1},y_{2},\mathbf{x}_{C}^{(1)},\dots,\mathbf{x}_{C}^{(r_{C}-1)},\allowbreak\mathbf{x}_{E}^{(1)},\dots,\mathbf{x}_{E}^{(r_{E})},x\Big{]}$ the iterated Ciminion polynomial system
-$\mathbf{f}^{(i)}=\begin{dcases}\mathcal{R}^{(1)}\left(\aleph,y_{1},y_{2}\right)-\mathbf{x}_{C}^{(1)},&i=1,\\ \mathcal{R}^{(i)}\left(\mathbf{x}_{C}^{(i-1)}\right)-\mathbf{x}_{C}^{(i)},&2\leq i\leq r_{C}-1,\\ \mathcal{R}^{(i)}\left(\mathbf{x}_{C}^{(r_{C}-1)}\right)-\mathbf{x}_{E}^{(1)},&i=r_{C},\\ \mathcal{R}^{(i-r_{C})}\left(\mathbf{x}_{E}^{(i-r_{C})}\right)-\mathbf{x}_{E}^{(i-r_{C}+1)},&r_{C}+1\leq i\leq r_{C}+r_{E}-1,\\ \mathcal{R}^{(r_{E})}\left(\mathbf{x}_{E}^{(r_{E})}\right)-\begin{pmatrix}c_{1}-p_{1}\\ c_{2}-p_{2}\\ x\end{pmatrix},&i=r_{C}+r_{E}.\end{dcases}$
+We call the polynomial system $\mathcal{F}_{\texttt{Ciminion}}=\left\{\mathbf{f}^{(i)}\right\}_{1\leq i\leq r}\subset\mathbb{F}_{q}\Big{[}y_{1},y_{2},\mathbf{x}^{(1)},\dots,\mathbf{x}^{(r-1)},x\Big{]}$ the iterated Ciminion polynomial system
+$\mathbf{f}^{(i)}=\begin{dcases}\mathcal{R}^{(1)}\left(\aleph,y_{1},y_{2}\right)-\mathbf{x}^{(1)},&i=1,\\ \mathcal{R}^{(i)}\left(\mathbf{x}^{(i-1)}\right)-\mathbf{x}^{(i)},&2\leq i\leq r-1,\\ \mathcal{R}^{(r)}\left(\mathbf{x}^{(r-1)}\right)-\begin{pmatrix}c_{1}-p_{1}\\ c_{2}-p_{2}\\ x\end{pmatrix},&i=r.\end{dcases}$
 2.2 Hydra
-Let $\mathbb{F}_{q}$ be a finite field, the Hydra [GØSW23] PRF over $\mathbb{F}_{q}$ is most also easily described via an illustration, see Figure 2.
+Let $\mathbb{F}_{q}$ be a finite field, the Hydra [GØSW23] PRF over $\mathbb{F}_{q}$ is also easily described via an illustration, see Figure 2.
 A nonce $\aleph\in\mathbb{F}_{q}$ and an initial value $\texttt{IV}\in\mathbb{F}_{q}^{3}$ are added to a key $\texttt{K}\in\mathbb{F}_{q}^{4}$, and then fed into the keyed body function $\mathcal{B}:\mathbb{F}_{q}^{4}\times\mathbb{F}_{q}^{4}\to\mathbb{F}_{q}^{8}$ which produces the state $\mathbf{y}||\mathbf{z}=(\mathbf{y},\mathbf{z})^{\intercal}\in\mathbb{F}_{q}^{4}\times\mathbb{F}_{q}^{4}$.
@@
 Let $\mathbf{M}_{\mathcal{E}}\in\operatorname{GL}_{4}\left(\mathbb{F}_{q}\right)$ be an MDS matrix, e.g.
-the Hydra designers recommend $\operatorname{circ}\left(2,3,1,1\right)$ or $\operatorname{circ}\left(3,2,1,1\right)$222 We consider circulant matrices with respect to a right shift, i.e.
-$\operatorname{circ}(a_{1},\dots,a_{n})=\begin{pmatrix}a_{1}&a_{2}&\dots&a_{n-1}&a_{n}\\ a_{n}&a_{1}&\dots&a_{n-2}&a_{n-1}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ a_{2}&a_{3}&\dots&a_{n}&a_{1}\end{pmatrix}.$ For the keyed permutation $\mathcal{H}_{\texttt{K}}$, let
+the Hydra designers recommend $\operatorname{circ}\left(2,3,1,1\right)$ or $\operatorname{circ}\left(3,2,1,1\right)$.333 We consider circulant matrices with respect to a right shift, i.e.
+$\operatorname{circ}(a_{1},\dots,a_{n})=\begin{pmatrix}a_{1}&a_{2}&\ldots&a_{n-1}&a_{n}\\ a_{n}&a_{1}&\ldots&a_{n-2}&a_{n-1}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ a_{2}&a_{3}&\ldots&a_{n}&a_{1}\end{pmatrix}.$ For the keyed permutation $\mathcal{H}_{\texttt{K}}$, let
 $\texttt{K}^{\prime}$ $=\begin{pmatrix}\texttt{K}\\ \mathbf{M}_{\mathcal{E}}\texttt{K}\end{pmatrix}\in\mathbb{F}_{q}^{8},$ (4)
-$f(\mathbf{x})$ $=\left(\sum_{i=1}^{8}(-1)^{\left\lfloor\frac{i-1}{4}\right\rfloor}\cdot x_{i}\right)^{2}.$ (12)
-Moreover, let $\mathbf{M}_{\mathcal{J}}\in\operatorname{GL}_{8}\left(\mathbb{F}_{q}\right)$ be an invertible matrix. Then the heads round function is defined as
+$f(x_{1},\dots,x_{n})$ $=\left(\sum_{i=1}^{8}(-1)^{\left\lfloor\frac{i-1}{4}\right\rfloor}\cdot x_{i}\right)^{2}.$ (5)
+Moreover, let $\mathbf{M}_{\mathcal{J}}\in\operatorname{GL}_{8}\left(\mathbb{F}_{q}\right)$ be an invertible matrix. The heads round function is defined as
 $\begin{split}\mathcal{J}_{i}:\mathbb{F}_{q}^{8}\times\mathbb{F}_{q}^{4}&\to\mathbb{F}_{q}^{8},\\ \left(\mathbf{x},\texttt{K}\right)&\mapsto\mathbf{M}_{\mathcal{J}}\big{(}x_{i}+f(\mathbf{x})\big{)}_{1\leq i\leq 8}+\texttt{K}^{\prime}+\mathbf{c}^{(i)},\end{split}$ (6)
@@
 $\begin{split}\mathcal{H}_{\texttt{K}}:\mathbb{F}_{q}^{4}\times\mathbb{F}_{q}^{4}\times\mathbb{F}_{q}^{4}&\to\mathbb{F}_{q}^{8},\\ (\mathbf{y},\mathbf{z},\texttt{K})&\mapsto\mathcal{J}_{r_{\mathcal{H}}}\circ\cdots\circ\mathcal{J}_{1}\left(\begin{pmatrix}\mathbf{y}\\ \mathbf{z}\end{pmatrix},\texttt{K}\right).\end{split}$ (7)
-Let $\mathbf{M}_{\mathcal{I}}\in\operatorname{GL}_{4}\left(\mathbb{F}_{q}\right)$ be a matrix of the form
+Let $\mathbf{M}_{\mathcal{I}}\in\operatorname{GL}_{4}\left(\mathbb{F}_{q}\right)$ be an invertible matrix of the form
 $\mathbf{M}_{\mathcal{I}}=\begin{pmatrix}\mu_{1,1}&1&1&1\\ \mu_{2,1}&\mu_{2,2}&1&1\\ \mu_{3,1}&1&\mu_{3,3}&1\\ \mu_{4,1}&1&1&\mu_{4,4}\end{pmatrix}.$ (8)
-Finally, for the rolling function let
+Note that the $\mu_{i,j}\in\mathbb{F}_{q}$ have to be chosen under additional constraints to prevent invariant subspaces [GØSW23, §5.2]. Finally, for the rolling function let
 $g(\mathbf{y},\mathbf{z})$ $=\left(\sum_{i=1}^{4}(-1)^{i-1}\cdot y_{i}\right)\cdot\left(\sum_{i=1}^{4}(-1)^{\left\lfloor\frac{i-1}{2}\right\rfloor}\cdot z_{i}\right),$ (9)
@@
 Then, the rolling function is defined as
-$\begin{split}\mathcal{R}:\mathbb{F}_{q}^{4}\times\mathbb{F}_{q}^{4}&\to\mathbb{F}_{q}^{8},\\ \begin{pmatrix}\mathbf{y}\\ \mathbf{z}\end{pmatrix}&\mapsto\mathbf{M}_{\mathcal{R}}\begin{pmatrix}\{y_{i}+g(\mathbf{y},\mathbf{z})\}_{1\leq i\leq 4}\\ \{z_{i}+g(\mathbf{z},\mathbf{y})\}_{1\leq i\leq 4}\end{pmatrix},\end{split}$ (18)
-and the $i$th rolling function is simply the $i$-fold composition $\mathcal{R}_{i}=\mathcal{R}\circ\cdots\circ\mathcal{R}$.
-Note that the degree increasing parts in $\mathcal{J}_{i}$ and $\mathcal{R}$ are generalized Lai–Massey permutations [RS22].
+$\begin{split}\mathcal{R}:\mathbb{F}_{q}^{4}\times\mathbb{F}_{q}^{4}&\to\mathbb{F}_{q}^{8},\\ \begin{pmatrix}\mathbf{y}\\ \mathbf{z}\end{pmatrix}&\mapsto\mathbf{M}_{\mathcal{R}}\begin{pmatrix}\big{(}y_{i}+g(\mathbf{y},\mathbf{z})\big{)}_{1\leq i\leq 4}\\ \big{(}z_{i}+g(\mathbf{z},\mathbf{y})\big{)}_{1\leq i\leq 4}\end{pmatrix},\end{split}$ (11)
+and the $i$th rolling function is simply the composition $\mathcal{R}_{i}=\mathcal{R}\circ\mathcal{R}_{i-1}+\mathbf{c}_{\mathcal{R}}^{(i)}$, where $\mathbf{c}_{\mathcal{R}}^{(i)}\in\mathbb{F}_{q}$ is a random constant.
+Note that the degree increasing parts in $\mathcal{J}_{i}$ and $\mathcal{R}$ are generalized Lai–Massey permutations [GOPS22, RS24].
 For $\kappa$ bits of security, the number of rounds for the Hydra heads with data limit $2^{\frac{\kappa}{2}}$ are computed via, see [GØSW23, §5.5],
@@
 Given two Hydra outputs it is straight-forward to set up an overdetermined polynomial system for Hydra.
-For ease of writing we denote the $i$th intermediate state variables for the $j$th output as $\mathbf{x}^{(i)}_{j}=\left(x_{j,1}^{(i)},\dots,x_{j,8}^{(i)}\right)^{\intercal}$ and analog for $\mathbf{y}=(y_{1},\dots,y_{4})^{\intercal}$, $\mathbf{z}=(z_{1},\dots,z_{4})^{\intercal}$ and $\mathbf{k}=(k_{1},\dots,k_{4})^{\intercal}$.
+For ease of writing we denote the $i$th intermediate state variables for the $j$th output as $\mathbf{x}^{(i)}_{j}=\left(x_{j,1}^{(i)},\dots,x_{j,8}^{(i)}\right)^{\intercal}$ and analogously for $\mathbf{y}=(y_{1},\dots,y_{4})^{\intercal}$, $\mathbf{z}=(z_{1},\dots,z_{4})^{\intercal}$ and $\mathbf{k}=(k_{1},\dots,k_{4})^{\intercal}$.
 Definition 2.2 (Iterated polynomial system for Hydra).
@@
 $\mathbf{f}_{1}^{(i)}$ =$=\begin{dcases}\mathcal{J}_{1}\left(\begin{pmatrix}\mathbf{y}\\ \mathbf{z}\end{pmatrix},\mathbf{k}^{\prime}\right)-\mathbf{x}_{1}^{(1)},&i=1,\\ \mathcal{J}_{i}\left(\mathbf{x}_{1}^{(i-1)},\mathbf{k}^{\prime}\right)-\mathbf{x}_{1}^{(i)},&2\leq i\leq r_{\mathcal{H}}-1,\\ \mathcal{J}_{r_{\mathcal{H}}}\left(\mathbf{x}_{1}^{(r_{\mathcal{H}}-1)},\mathbf{k}^{\prime}\right)+\begin{pmatrix}\mathbf{y}\\ \mathbf{z}\end{pmatrix}-\mathbf{c}_{1},&i=r_{\mathcal{H}},\end{dcases}$
-$\mathbf{f}_{\mathcal{R}}$ $=\mathcal{R}\left(\begin{pmatrix}\mathbf{y}\\ \mathbf{z}\end{pmatrix}\right)-\mathbf{x}_{2}^{(0)},$
-$\mathbf{f}_{2}^{(i)}$ =$=\begin{dcases}\mathcal{J}_{1}\left(\mathbf{x}_{2}^{(0)},\mathbf{k}^{\prime}\right)-\mathbf{x}_{2}^{(1)},&i=1,\\ \mathcal{J}_{i}\left(\mathbf{x}_{2}^{(i-1)},\mathbf{k}^{\prime}\right)-\mathbf{x}_{2}^{(i)},&2\leq i\leq r_{\mathcal{H}}-1,\\ \mathcal{J}_{r_{\mathcal{H}}}\left(\mathbf{x}_{2}^{(r_{\mathcal{H}})},\mathbf{k}^{\prime}\right)+\mathbf{x}_{2}^{(1)}-\mathbf{c}_{2},&i=r_{\mathcal{H}},\end{dcases}$
+$\mathbf{f}_{\mathcal{R}}$ $=\mathcal{R}\left(\begin{pmatrix}\mathbf{y}\\ \mathbf{z}\end{pmatrix}\right)+\mathbf{c}_{\mathcal{R}}^{(1)}-\mathbf{x}_{2}^{(0)},$
+$\mathbf{f}_{2}^{(i)}$ =$=\begin{dcases}\mathcal{J}_{1}\left(\mathbf{x}_{2}^{(0)},\mathbf{k}^{\prime}\right)-\mathbf{x}_{2}^{(1)},&i=1,\\ \mathcal{J}_{i}\left(\mathbf{x}_{2}^{(i-1)},\mathbf{k}^{\prime}\right)-\mathbf{x}_{2}^{(i)},&2\leq i\leq r_{\mathcal{H}}-1,\\ \mathcal{J}_{r_{\mathcal{H}}}\left(\mathbf{x}_{2}^{(r_{\mathcal{H}}-1)},\mathbf{k}^{\prime}\right)+\mathbf{x}_{2}^{(0)}-\mathbf{c}_{2},&i=r_{\mathcal{H}},\end{dcases}$
 where $\mathbf{k}^{\prime}=\begin{pmatrix}\mathbf{k}\\ \mathbf{M}_{\mathcal{E}}\mathbf{k}\end{pmatrix}.$
 This polynomial system consists of $816\cdot r_{\mathcal{H}}+$ 8 many polynomials in $416\cdot r_{\mathcal{H}}+$ 4 many variables.
-It can also easily be expanded to an arbitrary number of output samples.
+It can also easily be extended to an arbitrary number of output samples.
 2.3 Gröbner Bases
@@
     Every non-empty subset of $\mathbb{Z}_{\geq 0}^{n}$ has a smallest element under $>$, i.e. $>$ is a well-ordering on $\mathbb{Z}_{\geq 0}^{n}$.
-Now let $I\subset P=K[x_{1},\dots,x_{n}]$ be an ideal, and let $f\in P$.
-By $\operatorname{LM}_{>}\left(f\right)$ we denote the leading monomial of $f$, i.e.
+Now let $I\subset P=K[x_{1},\dots,x_{n}]$ be an ideal, and let $>$ be a term order on $P$.
+For $f\in P$ we denote with $\operatorname{LM}_{>}\left(f\right)$ the leading monomial of $f$, i.e.
 the largest monomial with non-zero coefficient in $f$ under $>$.
 A finite ideal basis $\mathcal{G}=\{g_{1},\dots,g_{m}\}\subset I$ such that
-$\big{(}\operatorname{LM}_{>}(g_{1}),\dots,\operatorname{LM}_{>}(g_{m})\big{)}=\operatorname{in}_{>}(I)$ (20)
-is called a $>$-Gröbner basis of $I$.
-I.e., the leading monomials of $\mathcal{G}$ are minimal under $>$.
-With Gröbner bases one can solve many important computational problems like the membership problem or the computation of the zero locus.
+$\big{(}\operatorname{LM}_{>}(g_{1}),\dots,\operatorname{LM}_{>}(g_{m})\big{)}=\big{(}\operatorname{LM}_{>}(f)\;\big{|}\;f\in I\big{)}$ (13)
+is called a $>$-Gröbner basis of $I$. I.e., the leading monomials of $\mathcal{G}$ are minimal under $>$.
+We abbreviate the ideal of leading terms of $I$ as $\operatorname{LM}_{>}(I)=\big{(}\operatorname{LM}_{>}(f)\;\big{|}\;f\in I\big{)}$.
+Also, for ease of writing we often denote the variables of $P$ in vector form $\mathbf{x}=(x_{1},\dots,x_{n})^{\intercal}$.
+If not specified otherwise, then we always assume that the variables within a variable vector $\mathbf{x}$ are naturally ordered with respect to $>$, i.e.
+$x_{1}>\ldots>x_{n}$.
+With Gröbner bases one can solve many important computational problems like the membership problem or the computation of the set of zeros of an ideal.
 For a general introduction to Gröbner bases we refer to [KR00, KR05, CLO15].
+In this paper we work with the lexicographic (LEX) and the degree reverse lexicographic (DRL) term order.
+For $\mathbf{a},\mathbf{b}\in\mathbb{Z}_{\geq 0}^{n}$, we have that $\mathbf{a}>_{LEX}\mathbf{b}$ if the first non-zero entry of $\mathbf{a}-\mathbf{b}$ is positive.
+We have that $\mathbf{a}>_{DRL}\mathbf{b}$ if either $\sum_{i=1}^{n}a_{i}>\sum_{i=1}^{n}b_{i}$ or $\sum_{i=1}^{n}a_{i}=\sum_{i=1}^{n}b_{i}$ and the last non-zero entry of $\mathbf{a}-\mathbf{b}$ is negative.
+In more generality, a term order $>$ on $P$ is called degree compatible if for any $f,g\in P$ the degree inequality $\deg\left(f\right)>\deg\left(g\right)$ implies that $f>g$.
+Obviously, the DRL term order is degree compatible.
 2.3.1 Linear Algebra-Based Gröbner Basis Algorithms
-The first algorithm to compute Gröbner bases was Buchberger’s algorithm, see e.g.
-[CLO15, Chapter 2 §7].
-Given an input polynomial system $\mathcal{F}\subset P=k[x_{1},\dots,x_{n}]$ and a term order $>$ on $P$, Buchberger’s algorithms performs pairwise cancellation of the leading monomials under $>$ of two elements in $\mathcal{F}$ followed by multivariate division by remainder.
-If a non-zero element is produced, then it is added to $\mathcal{F}$ and the process is repeated.
-Buchberger proved that this procedure terminates and produces a $>$-Gröbner basis.
-Lazard [Laz83] generalized this construction by performing multiple leading term cancellations and reductions at the same time via a matrix representation of the polynomial system $\mathcal{F}$.
-Assume that =$\mathcal{F}=\{f_{1},\dots,f_{m}\}\subset P$ is homogeneous and fix a term order $>$ on $P$.
+Gröbner bases can be constructed via Gaussian elimination on the so-called Macaulay matrices.
+Let $\mathcal{F}\subset P=K[x_{1},\dots,x_{n}]$ be homogeneous, and let $>$ be a term order on $P$.
 The homogeneous Macaulay matrix $M_{d}$ for $\mathcal{F}$ has columns indexed by the monomials of degree $d$ in $P$ sorted via $>$, and its rows are indexed by $s\cdot f$, where $f\in\mathcal{F}$ and $s\in P$ is a monomial such that $\deg\left(s\cdot f\right)=d$.
@@
 If $\mathcal{F}$ is inhomogeneous, then one replaces the degree equalities by inequalities to obtain the inhomogeneous Macaulay matrix $M_{\leq d}$.
-Obviously, if we perform Gaussian elimination on $M_{0},\dots,M_{d}$ respectively $M_{\leq d}$ for a large enough value of $d$, then we construct a $>$-Gröbner basis for $\mathcal{F}$.
-The least such $d$ is called the solving degree, to the best of our knowledge this notion was first introduced in [DS13].
+Obviously, if we perform Gaussian elimination on $M_{0},\dots,M_{d}$ (respectively $M_{\leq d}$) for $d$ large enough, then we construct a $>$-Gröbner basis for $\mathcal{F}$.
+This idea can be traced back to Lazard [Laz83].
+The least $d$ which yields a Gröbner basis is of special interest.
 Definition 2.3 (Solving degree, [CG21, Definition 6]).
@@
 We denote it by $\operatorname{sd}_{>}(\mathcal{F})$.
-If $\mathcal{F}$ is homogeneous, we consider the homogeneous Macaulay matrix $M_{d}$ and let the solving degree of $\mathcal{F}$ be the least degree $d$ such that Gaussian elimination on $M_{0},\dots,M_{d}$ produces a Gröbner basis of $\mathcal{F}$ with respect to $>$.
-By [CG21, Theorem 7] the homogeneous solving degree always upper bounds the inhomogeneous one, i.e.
-$\operatorname{sd}_{DRL}\left(\mathcal{F}\right)\leq\operatorname{sd}_{DRL}\left(\mathcal{F}^{\text{hom}}\right).$ (21)
+If $\mathcal{F}$ is homogeneous, then we consider the homogeneous Macaulay matrix $M_{d}$ and let the solving degree of $\mathcal{F}$ be the least degree $d$ such that Gaussian elimination on $M_{0},\dots,M_{d}$ produces a Gröbner basis of $\mathcal{F}$ with respect to $>$.
 Algorithms which perform Gaussian elimination on the Macaulay matrices are generally referred to as linear algebra-based Gröbner basis algorithms.
 Modern examples of such algorithms are Faugère’s F4 [Fau99] and Matrix-F5 [Fau02] algorithms.
-Since the solving degree is a priori unknown for most polynomial system, F4/5 construct the Macaulay matrices and perform Gaussian elimination for increasing values of $d$.
+Since the solving degree is a priori unknown for most polynomial systems, F4/5 construct the Macaulay matrices for increasing values of $d$.
 Therefore, they require a stopping criterion to decide whether a Gröbner basis has already been found.
@@
 For this reason, we also consider the solving degree as complexity measure for advanced linear algebra-based algorithms.
-2.3.2 Complexity Estimates via the Solving Degree
-It is well-known that the complexity of Gaussian elimination of a matrix $\mathbf{A}\in K^{n\times m}$, where $K$ is a field and $r=\operatorname{rank}\left(\mathbf{A}\right)$, is given by [Sto00, §2.2] $\mathcal{O}\left(n\cdot m\cdot r^{\omega-2}\right)$ field operations, where $2.372862\leq\omega<$ 2.37286 is a linear algebra constant [AW21].
-Given a homogeneous polynomial system =$\mathcal{F}=\{f_{1},\dots,f_{m}\}\subset P=K[x_{1},\dots,x_{n}]$, it is well-known that the number of monomials in $P$ of degree $d$ is $\binom{n+d-1}{d}$.
-Moreover, for $f\in\mathcal{F}$ the number of rows $\hat{m}\cdot f$, where $\hat{m}\in P$ a monomial such that $\deg\left(\hat{m}\cdot f\right)=d$, is also bounded by $\binom{n+d-1}{d}$.
-Thus, the complexity of Gaussian elimination on the homogeneous Macaulay matrices $M_{0},\dots,M_{d}$ is bounded by
-$\mathcal{O}\left(m\cdot d\cdot\binom{n+d-1}{d}^{\omega}\right).$ (22)
-Also, note that dehomogenization of a Gröbner basis yields an inhomogeneous Gröbner basis [KR05, Proposition 4.3.18] (if the term order is compatible with homogenization).
-Therefore, we can also regard Equation 22 as complexity estimate for Gaussian elimination on the inhomogeneous Macaulay matrix $M_{\leq d}$.
-2.4 Generic Coordinates
-Caminata & Gorla [CG23] have proven that for a special class of polynomial systems the DRL solving degree is always upper bounded by the so-called Macaulay bound.
-They called these polynomial to be in generic coordinates.
-For the formal definition of this notion, we use the shorthand notation “$r\nmid 0\mod I$” for expressing that $r$ is not a zero-divisor in $P/I$.
-Moreover, we always assume that the homogenization variable $x_{0}$ is the least variable under DRL.
-Definition 2.4 ([CG21, CG22, Definition 5]).
-Let $K$ be an infinite field.
-Let $I\subset K[x_{0},\dots,\allowbreak x_{n}]$ be a homogeneous ideal with $|\mathcal{Z}_{+}(I)|<\infty$.
-We say that $I$ is in generic coordinates if either $|\mathcal{Z}_{+}(I)|=0$ or $x_{0}\nmid 0\mod I^{\text{sat}}$.
+2.3.2 Polynomial Systems in Generic Coordinates
+Polynomial systems in generic coordinates have a proven DRL solving degree upper bound. For their definition we recall the degree of regularity of a polynomial system.
+Definition 2.4 (Degree of regularity, [BFS04, Definition 4]).
+Let $K$ be a field, and let $\mathcal{F}\subset K[x_{1},\dots,x_{n}]$. If $\dim\left(\mathcal{F}^{\text{top}}\right)=0$, then the degree of regularity is defined as
+$d_{\operatorname{reg}}\left(\mathcal{F}\right)=\min\left\{d\geq 0\;\middle|\;\dim_{\mathbb{F}_{q}}\Big{(}\big{\{}f\in\left(\mathcal{F}^{\text{top}}\right)\;\big{|}\;\deg\left(f\right)=d\big{\}}\Big{)}=\binom{n+d-1}{d}\right\}.$
+If $\dim\left(\mathcal{F}^{\text{top}}\right)>0$, then $d_{\operatorname{reg}}\left(\mathcal{F}\right)=\infty$.
+We present an equivalent definition for generic coordinates which requires slightly less knowledge about commutative algebra and projective geometry than the original definition from [CG21, CG22, Definition 5].
+Definition 2.5 (cf. [CG21, CG22, Definition 5], [Ste24, Theorem 3.2]).
+Let $K$ be an infinite field, and let $\mathcal{F}\subset P=K[x_{1},\dots,x_{n}]$.
+We say that $\left(\mathcal{F}^{\text{hom}}\right)\subset P[x_{0}]$ is in generic coordinates if either $d_{\operatorname{reg}}\left(\mathcal{F}\right)<\infty$ or $\sqrt{\mathcal{F}^{\text{hom}}}=(x_{0},\dots,x_{n})$.
 Let $K$ be any field, and let $K\subset L$ be an infinite field extension.
 $I$ is in generic coordinates over $K$ if $I\otimes_{K}L[x_{0},\dots,x_{n}]\subset L[x_{0},\dots,x_{n}]$ is in generic coordinates.
-As already mentioned, polynomial systems in generic coordinates have a proven DRL solving degree upper bound.
-Theorem 2.5 ([CG21, Theorem 9, 10, Corollary 2]).
-Let $K$ be an algebraically closed field, and let =$\mathcal{F}=\{f_{1},\dots,f_{m}\}\subset K[x_{1},\dots,x_{n}]$ be an inhomogeneous polynomial system such that $\left(\mathcal{F}^{\text{hom}}\right)$ is in generic coordinates.
-Let $d_{i}=\deg\left(f_{i}\right)$ and assume that $d_{1}\geq\ldots\geq d_{m}$, then for $l\in\min\{n+1,m\}$
-$\operatorname{sd}_{DRL}\left(\mathcal{F}\right)\leq d_{1}+\ldots+d_{l}-l+1.$
-In the Gröbner basis and algebraic cryptanalysis literature the latter inequality is also usually referred as Macaulay bound.
-Moreover, Lazard [Laz83] first observed that DRL Gröbner bases for “generic” polynomial systems can be found within the Macaulay bound.
-Finally, being in generic coordinates was recently characterized by Steiner [Ste24].
-Theorem 2.6 ([Ste24, Theorem 3.2]).
-Let $K$ be an algebraically closed field, and let =$\mathcal{F}=\{f_{1},\dots,f_{m}\}\subset K[x_{1},\dots,x_{n}]$ be a polynomial system such that
-  1. (i)
-    $(\mathcal{F})\neq(1)$, and
-  2. (ii)
-    $\dim\left(\mathcal{F}\right)=0$.
-Then the following are equivalent.
-  1. (1)
-    $\left(\mathcal{F}^{\text{hom}}\right)$ is in generic coordinates.
-  2. (2)
-    $\sqrt{\mathcal{F}^{\text{top}}}=\left(x_{1},\dots,x_{n}\right)$.
-  3. (3)
-    $\left(\mathcal{F}^{\text{top}}\right)$ is zero-dimensional in $K[x_{1},\dots,x_{n}]$.
-  4. (4)
-    For every $1\leq i\leq n$ there exists $d_{i}\in\mathbb{Z}_{\geq 1}$ such that $x_{i}^{d_{i}}\in\operatorname{in}_{DRL}\left(\mathcal{F}^{\text{hom}}\right)$.
-Let $\mathcal{F}\subset K[x_{1},\dots,x_{n}]$ be an inhomogeneous polynomial system, to prove that $\mathcal{F}^{\text{hom}}$ is in generic coordinates one can use the following procedure based on (2):
+Let $\mathcal{F}\subset K[x_{1},\dots,x_{n}]$ be an inhomogeneous polynomial system, to prove that $\left(\mathcal{F}^{\text{hom}}\right)$ is in generic coordinates one can use the following procedure based on [Ste24, Theorem 3.2]:
   1. (1)
@@
 If Step (2) is successful for all $1\leq i\leq n$, then the polynomial system $\mathcal{F}^{\text{hom}}$ is in generic coordinates.
-Another quantity in the literature that is often used to express the complexity of finding a DRL Gröbner basis is the so-called degree of regularity.
-Definition 2.7 (Degree of regularity, [BFS04, Definition 4]).
-Let $K$ be a field, and let $\mathcal{F}\subset P=K[x_{1},\dots,x_{n}]$.
-Assume that $\left(\mathcal{F}^{\text{top}}\right)_{d}=P_{d}$ for some integer $d\geq 0$.
-The degree of regularity is defined as
-$d_{\operatorname{reg}}\left(\mathcal{F}\right)=\min\left\{d\geq 0\;\middle|\;\left(\mathcal{F}^{\text{top}}\right)_{d}=P_{d}\right\}.$
-Note that by Theorem 2.6 a polynomial system is in generic coordinates if and only if it admits a finite degree of regularity.
-2.5 The Eigenvalue Method
-As outlined in the introduction, in the algebraic cryptanalysis literature the standard approach to polynomial system solving is
-  1. (1)
-    DRL Gröbner basis computation,
-  2. (2)
-    term order conversion to LEX, and
-  3. (3)
-    univariate factoring.
-In this section we shortly discuss that there is an alternative to term order conversion via linear algebra-based techniques [KR16, Chapter 6].
-Our motivation to change the strategy is quite simple.
-Via the Eigenvalue Method we will be able to exploit the multiplicative structure of Ciminion and Hydra DRL Gröbner bases.
-Consequently, this will provide a slight improvement compared to the generic complexity of the probabilistic FGLM algorithm [FGHR14].
-Let $I\subset P=K[x_{1},\dots,x_{n}]$ be a zero-dimensional ideal, i.e.
-$P/I$ is a finite dimensional $K$-vector space [KR00, Proposition 3.7.1], and let $d=\dim_{K}\left(P/I\right)$.
-If $>$ is a term order on $P$ and $\mathcal{G}\subset I$ is $>$-Gröbner basis, then a $K$-vector space basis $\mathcal{B}$ of $P/I$ is given by the monomials not contained in $\operatorname{in}_{>}\left(\mathcal{G}\right)=\operatorname{in}_{>}\left(I\right)$.
-Now let $f\in P$, then the multiplication map
-$\begin{split}\theta_{f}:P&\to P/I,\\ x&\mapsto x\cdot f,\end{split}$ (23)
-can be represented via a matrix.
-This matrix is called the multiplication matrix $\mathbf{M}_{f}$ for $f$.
-Given a $>$-Gröbner basis its computation is straight-forward: Index the columns of $\mathbf{M}_{f}$ by the $K$-vector space basis $\mathcal{B}$, and the rows by $b\cdot f$, where $b\in\mathcal{B}$.
-The row $b\cdot f$ is then simply the coefficient vector of $b\cdot f$ in $P/I$ which can be computed via $b\cdot f\mod(\mathcal{G})$ with respect to $>$.
-Now assume in addition that $K$ is algebraically closed, then for every point $\mathbf{x}\in\mathcal{Z}\left(I\right)$ the point $f(\mathbf{x})$ is an eigenvalue of $\mathbf{M}_{f}$, see [KR16, Proposition 6.2.1].
-Therefore, we can compute the zero locus $\mathcal{Z}\left(I\right)$ as follows: Compute the eigenvalues of $\mathbf{M}_{x_{1}},\dots,\mathbf{M}_{x_{n}}$, iterate over all possible combinations, and verify whether a combination is indeed a solution to $I$.
-This approach is known as Eigenvalue Method [KR16, §6.2.A].
-In his seminal work Strassen showed that matrix inversion, determinant and Gaussian elimination have the same complexity as matrix multiplication up to some constant [Str69].
-For the determinant this observation follows form LU-decomposition of block square matrices.
-We recall this decomposition since we will exploit it at a later point in this paper.
-Let $\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D}\in K^{2\cdot N\times 2\cdot N}$ be matrices, then we have the block matrix decomposition
-$\begin{pmatrix}\mathbf{A}&\mathbf{B}\\ \mathbf{C}&\mathbf{D}\end{pmatrix}=\begin{pmatrix}\mathbf{I}_{N\times N}&\mathbf{0}_{N\times N}\\ \mathbf{C}\mathbf{A}^{-1}&\mathbf{I}_{N\times N}\end{pmatrix}\begin{pmatrix}\mathbf{A}&\mathbf{B}\\ \mathbf{0}_{N\times N}&\mathbf{D}-\mathbf{C}\mathbf{A}^{-1}\mathbf{B}\end{pmatrix},$ (24)
-provided that $\mathbf{A}$ is invertible. With well-known formulae for block matrix determinants, see e.g. [Sil00], one then has that
-$\det\begin{pmatrix}\mathbf{A}&\mathbf{B}\\ \mathbf{C}&\mathbf{D}\end{pmatrix}=\det\left(\mathbf{A}\right)\cdot\det\left(\mathbf{D}-\mathbf{C}\mathbf{A}^{-1}\mathbf{B}\right).$ (25)
-So the determinant of the block matrix has been reduced to two determinants of smaller matrices, one matrix inversion and two matrix multiplications.
-By Strassen [Str69] the complexity of multiplying two $N\times N$ matrices is in $\mathcal{O}\left(N^{\omega+o(1)}\right)$, where $2.372862\leq\omega<$ 2.37286 is again the linear algebra constant [AW21].
-Note that the term $N^{o(1)}$ is usually ignored in complexity estimations.
-Lastly, it is well known that the eigenvalues of a matrix $\mathbf{M}\in K^{N\times N}$ are roots of its characteristic polynomial
-$\chi_{\mathbf{M}}(x)=\det\left(\mathbf{I}_{N\times N}\cdot x-\mathbf{M}\right)\in K[x].$ (26)
-Therefore, construction of the characteristic polynomials in the Eigenvalue Method can be performed in
-$\mathcal{O}\left(n\cdot d^{\omega}\right),$ (27)
-where $d=\dim_{K}\left(P/I\right)$.
-In addition, if we are only interested in solutions over a finite field $\mathbb{F}_{q}$, then factorization of the characteristic polynomials can be done via the GCD with the field equation whose complexity was given in Equation 1.
-Note that the probabilistic FGLM algorithm [FGHR14, Proposition 3] has polylogarithmic complexity $\tilde{\mathcal{O}}\left(d^{\omega}\right)$ for the construction of a LEX Gröbner basis of $I$ if $I$ is in shape position, and polylogarithmic complexity $\tilde{\mathcal{O}}\left(n\cdot d^{\omega}\right)$ for the construction of a LEX Gröbner basis of $\sqrt{I}$, which applies to any ideal not in shape position.
-This is no coincidence, the probabilistic FGLM algorithm constructs a LEX Gröbner basis via the multiplication matrix for the last variable.
-Hence, for both methods the complexity is dominated by $\mathcal{O}\left(d^{\omega}\right)$.
-3 Analysis of Ciminion
-At round level Ciminion is a three branch Feistel network.
-In [Ste24, §4.1] a DRL Gröbner basis for Feistel-$2n/n$ polynomial systems has been computed by substituting the linear branches into the non-linear ones.
-In principle, we are going to replicate this strategy to construct a DRL Gröbner basis for Ciminion.
-3.1 A DRL Gröbner Basis for Ciminion
-For simplicity let us denote the matrix in the $i$th Ciminion round by $\mathbf{A}_{i}\in\mathbb{F}_{q}^{3\times 3}$, then
-$\mathbf{A}_{i}^{-1}=\begin{pmatrix}0&1&-c_{4}^{(i)}\\ -1&0&1\\ 1&0&0\end{pmatrix},$ (28)
-where $c_{4}^{(i)}\notin\mathbb{F}_{q}\setminus\{0,1\}$.
-Theorem 3.1.
-Let $\mathbb{F}_{q}$ be a finite field, and let
-$\mathcal{F}_{\texttt{Ciminion}}=\{\mathbf{f}^{(1)},\dots,\mathbf{f}^{(r_{C}+r_{E})}\}\subset\mathbb{F}_{q}\Big{[}y_{1},y_{2},\mathbf{x}_{C}^{(1)},\allowbreak\dots,\mathbf{x}_{C}^{(r_{C}-1)},\allowbreak\mathbf{x}_{E}^{(1)},\dots,\mathbf{x}_{E}^{(r_{E})},x\Big{]}$
-be a Ciminion polynomial system.
-For the DRL term order $y_{1}>y_{2}>\mathbf{x}_{C}^{(1)}>\ldots>\mathbf{x}_{C}^{(r_{C}-1)}>\mathbf{x}_{E}^{(1)}>\mathbf{x}_{E}^{(r_{E})}>x$, the following procedure computes a DRL Gröbner basis for Ciminion:
-  1. (1)
-    For every round $1\leq i\leq r_{C}+r_{E}$ compute $\mathbf{g}^{(i)}=\mathbf{A}_{i}^{-1}\mathbf{g}_{i}$.
-  2. (2)
-    Collect the linear polynomials from the $\mathbf{g}^{(i)}$’s in $\mathcal{L}$, and perform Gaussian elimination on $\mathcal{L}$.
-  3. (3)
-    Reduce the non-linear polynomials in the $\mathbf{g}^{(i)}$’s modulo $\mathcal{L}$.
-Proof.
-By inspection of the round function (Equation 9) we have that
-$\mathbf{g}^{(1)}$ =($\ y_{1}\\ y_{2}+\aleph\cdot y_{1}\end{pmatrix}+\mathbf{A}_{1}^{-1}\left(\mathbf{c}^{(1)}-\mathbf{x}_{C}^{(1)}\right),$ (29)
-$\mathbf{g}^{(i)}$ $=\begin{pmatrix}x_{C,1}^{(i-1)}\\[5.0pt] x_{C,2}^{(i-1)}\\[5.0pt] x_{C,3}^{(i-1)}+x_{C,1}^{(i-1)}\cdot x_{C,2}^{(i-1)}\end{pmatrix}+\mathbf{A}_{i}^{-1}\left(\mathbf{c}^{(i)}-\begin{pmatrix}x_{C,1}^{(i)}\\ x_{C,2}^{(i)}\\ x_{C,3}^{(i)}\end{pmatrix}\right),$ (30)
-$\mathbf{g}^{(r_{C}+r_{E})}$ $=\begin{pmatrix}x_{E,1}^{(r_{E})}\\[5.0pt] x_{E,2}^{(r_{E})}\\[5.0pt] x_{E,3}^{(r_{E})}+x_{E,1}^{(r_{E})}\cdot x_{C,2}^{(r_{E})}\end{pmatrix}+\mathbf{A}_{i}^{-1}\left(\mathbf{c}^{(r_{E})}-\begin{pmatrix}c_{1}-p_{1}\\ c_{2}-p_{2}\\ x\end{pmatrix}\right),$ (31)
-and analog for the equations involving the $\mathbf{x}_{E}^{(i)}$’s. Hence, $\mathcal{L}$ contains $\mathbf{g}^{(1)}$ and the first two components of all remaining $\mathbf{g}^{(i)}$’s.
-For Gaussian elimination on $\mathcal{L}$ let us start in the last round $i=r_{C}+r_{E}$.
-By construction and Equation 28 the linear polynomials with leading monomials $x_{E,1}^{(r_{E})}$, $x_{E,2}^{(r_{E})}$ contain a linear term in $x$ and a constant term.
-Moving to $i=r_{C}+r_{E}-1$ next, then the linear polynomials from $\mathbf{g}^{(r_{C}+r_{E}-1)}$ depend on $\mathbf{x}_{E}^{(r_{C}+r_{E}-1)}$ and $\mathbf{x}_{E}^{(r_{C}+r_{E})}$, via Gaussian elimination the variables $x_{E,1}^{(r_{E})}$ and $x_{E,2}^{(r_{E})}$ get eliminated.
-Thus, the linear polynomials with leading monomials in $x_{E,1}^{(r_{E}-1)}$, $x_{E,2}^{(r_{E}-1)}$ can only depend on $x_{E,3}^{(r_{E})}$ and $x$.
-Moreover, by Equation 28 the coefficients for $x_{E,3}^{(r_{E})}$ are non-zero.
-Via a downwards induction we can then conclude that the linear polynomials with leading monomials $x_{\ast,1}^{(i)}$, $x_{\ast,2}^{(i)}$ only depend on the variables $x_{\ast,3}^{(i+1)},\dots,x_{E,3}^{(r_{E})},x$ and that the coefficients of $x_{\ast,3}^{(i+1)}$ are non-zero.
-The first round needs special care since the first row of $\mathbf{A}_{1}^{-1}$ is non-zero on the variable $x_{C,2}^{(1)}$.
-(If it were zero on $x_{C,1}^{(1)}$ and $x_{C,2}^{(1)}$, the first polynomial in $\mathbf{g}^{(1)}$ would have leading monomial $x_{C,3}^{(1)}$ and the proof would be finished.) Expansion of Equation 29 yields
-$\mathbf{g}^{(1)}=\begin{pmatrix}\aleph\\ y_{1}\\ y_{2}+\aleph\cdot y_{1}\end{pmatrix}+\mathbf{A}_{1}^{-1}\mathbf{c}_{1}-\begin{pmatrix}x_{C,2}^{(1)}-c_{4}^{(1)}\cdot x_{C,3}^{(1)}\\ -x_{C,1}^{(1)}+x_{C,3}^{(1)}\\ x_{C,1}^{(1)}\end{pmatrix}.$
-Producing a polynomial with leading monomial $y_{2}$ is straight-forward, we simply have to take $g_{3}^{(1)}-\aleph\cdot g_{2}^{(1)}$.
-Note that there is second polynomial with leading monomial $x_{C,2}^{(1)}$ after inversion of the matrices
-$g_{2}^{(2)}=x_{C,2}^{(1)}+\left(\mathbf{A}_{2}^{-1}\mathbf{c}^{(2)}\right)_{2}-\left(-x_{C,1}^{(2)}+x_{C,3}^{(2)}\right).$
-Then
-$g_{1}^{(1)}+g_{2}^{(2)}=\aleph+\left(\mathbf{A}_{1}^{-1}\mathbf{c}^{(1)}\right)_{1}+c_{4}^{(1)}\cdot x_{C,3}^{(1)}+\left(\mathbf{A}_{2}^{-1}\mathbf{c}^{(2)}\right)_{2}+x_{C,1}^{(2)}-x_{C,3}^{(2)},$
-and by the choice of our DRL term order and $c_{4}^{(1)}\neq 0$ this polynomial has leading monomial $x_{C,3}^{(1)}$.
-Obviously, after Gaussian elimination of $\mathcal{L}$ the terms of $g_{1}^{(1)}+g_{2}^{(2)},g_{2}^{(1)},g_{3}^{(1)}-\aleph\cdot g_{2}^{(1)}$ after the leading monomials will also only depend on the variables $x_{C,3}^{(2)},\dots,x_{E,3}^{(r_{E})},x$.
-Lastly, let us consider the reduction modulo $\mathcal{L}$. By our previous observations all linear equations are of the form
-$y_{1}+\mathcal{A}_{1}\left(x_{C,3}^{(2)},\dots,x_{E,3}^{(r_{C}+r_{E}-1)},x\right)$ $=0,$
-$y_{2}+\mathcal{A}_{2}\left(x_{C,3}^{(2)},\dots,x_{E,3}^{(r_{C}+r_{E}-1)},x\right)$ $=0,$
-$x_{E,3}^{(1)}+\mathcal{A}_{3}\left(x_{C,3}^{(2)},\dots,x_{E,3}^{(r_{C}+r_{E}-1)},x\right)$ $=0,$
-$x_{\ast,1}^{(i)}+c_{4}^{(i)}\cdot x_{\ast,3}^{(i+1)}+\ldots$ $=0,$
-$x_{\ast,2}^{(i)}-x_{\ast,3}^{(i+1)}+\ldots$ $=0,$
-$x_{E,1}^{(r_{E})}+c_{4}^{(r_{C}+r_{E})}\cdot x+\ldots$ $=0,$
-$x_{E,2}^{(r_{E})}-x+\ldots$ $=0,$
-where $\mathcal{A}_{1}$, $\mathcal{A}_{2}$ and $\mathcal{A}_{3}$ are affine polynomials. So, we can directly conclude from Equation 30 that
-$\hat{g}_{3}^{(i)}=\left(g_{3}^{(i)}\mod(\mathcal{L})\right)\in\mathbb{F}_{q}\Big{[}x_{C,3}^{(2)},\dots,x_{E,3}^{(r_{C}+r_{E}-1)},x\Big{]},$
-where $2\leq i\leq r_{C}+r_{E}$, has leading monomial ${x_{\ast,3}^{(i)}}^{2}$ and respectively $x^{2}$ for the last round.
-Thus, we have constructed a polynomial system with pairwise coprime leading monomials, so by [CLO15, Chapter 2 §9 Theorem 3, Proposition 4] we have constructed a Gröbner basis. ∎
-Note that if we only keep the $r_{E}+r_{C}-1$ quadratic polynomials of $\mathcal{G}$, then we still have a fully determined polynomial system in $r_{E}+r_{C}-1$ variables as well as a Gröbner basis.
-Therefore, we call the quadratic polynomials of $\mathcal{G}$ the downsized Ciminion polynomial system.
-For the next corollary, recall that a LEX Gröbner basis =$\mathcal{F}=\{f_{1},\dots,f_{n}\}\subset K[x_{1},\dots,\allowbreak x_{n}]$ is said to be in $x_{n}$-shape position if it is of the form
-$\begin{split}f_{1}&=x_{1}-\tilde{f}_{n}(x_{n}),\\ &\dots\\ f_{n-1}&=x_{n-1}-\tilde{f}_{n-1}(x_{n}),\\ f_{n}&=\tilde{f}_{n}(x_{n}).\end{split}$ (32)
-Note that the LEX Gröbner basis of a radical ideal can be brought into shape position for some ordering of the variables, see [KR00, Theore 3.7.25].
-Corollary 3.2.
-In the situation of LABEL:Th:_Ciminion_Gr\"obner_basis, let $\mathcal{G}_{\texttt{Ciminion}}$ denote the Ciminion DRL Gröbner basis.
-  1. (1)
-    $\dim_{\mathbb{F}_{q}}\left(\mathcal{F}_{\texttt{Ciminion}}\right)=2^{r_{C}+r_{E}-1}$.
-  2. (2)
-    Let $I\subset\overline{\mathbb{F}_{q}}\Big{[}y_{1},y_{2},\mathbf{x}_{C}^{(1)},\allowbreak\dots,\mathbf{x}_{C}^{(r_{C}-1)},\allowbreak\mathbf{x}_{E}^{(1)},\dots,\mathbf{x}_{E}^{(r_{E})},x\Big{]}[x_{0}]$ be a homogeneous ideal such that $\mathcal{Z}_{+}(I)\neq\emptyset$ and $\mathcal{G}^{\text{hom}}\subset I$.
-Then $I$ is in generic coordinates.
-  3. (3)
-    $\operatorname{sd}_{DRL}\left(\mathcal{F}_{\texttt{Ciminion}}\right)\leq r_{C}+r_{E}$.
-  4. (4)
-    The LEX Gröbner basis for Ciminion under $y_{1}>y_{2}>\mathbf{x}_{C}^{(1)}>\ldots>\mathbf{x}_{C}^{(r_{C}-1)}>\mathbf{x}_{E}^{(1)}>\mathbf{x}_{E}^{(r_{E})}>x$ is in $x$-shape position.
-Proof.
-(1), (2) and (3) are consequences of the DRL Gröbner basis from LABEL:Th:_Ciminion_Gr\"obner_basis and application of Theorems 2.5 and 2.6.
-For the last claim we take a look at the shape of the DRL Gröbner basis
-$g_{3}^{(2)}$ $=\mathcal{Q}_{2}\left(x_{C,3}^{(2)},\dots,x_{E,3}^{(i)},x\right)-\mathcal{A}_{2}\left(x_{C,3}^{(2)},\dots,x_{E,3}^{(r_{E})},x\right),$
-$\dots$
-$g_{3}^{(i)}$ $=\mathcal{Q}_{i}\left(x_{\ast,3}^{(i)},\dots,x_{E,3}^{(r_{E})},x\right)-\mathcal{A}_{i}\left(x_{\ast,3}^{(i-1)},\dots,x_{E,3}^{(r_{E})},x\right),$
-$\dots$
-$g_{3}^{(r_{C}+r_{E})}$ $=\mathcal{Q}\left(x\right)-\mathcal{A}_{r_{C}+r_{E}}\left(x_{E,3}^{(r_{E})},x\right),$
-where the $\mathcal{Q}_{i}$’s are quadratic polynomials which contain the monomial ${x_{\ast,3}^{(i)}}^{2}$ and respectively $x^{2}$ for the last round, and the $\mathcal{A}_{i}$’s are affine polynomials that contain the monomial $x_{E,3}^{(i-1)}$ for all $i\geq 3$ and without any restrictions on $\mathcal{A}_{2}$.
-Then by iterative substitution of $g_{3}^{(r_{C}+r_{E})}$ into $g_{3}^{(r_{C}+r_{E}-1)}$, $g_{3}^{(r_{C}+r_{E}-1)}$ into $g_{3}^{(r_{C}+r_{E}-2)}$, etc.
-we produce a LEX Gröbner basis in $x$-shape position.
-∎
-3.2 Bariant’s Attack
-Recently Bariant introduced an attack on Ciminion via univariate greatest common divisor computations [Bar23].
-As in our polynomial model he introduced an auxiliary variable $X$ for the truncated component.
-If one inverts $p_{E}$ and $p_{C}$ using $(c_{1},c_{2},X)$ as output state one then obtains three polynomials, see [Bar23, Fig.
-3],
-$\aleph=f_{1}(X),\qquad K_{1}=f_{2}(X),\qquad K_{2}=f_{3}(X).$ (33)
-Since the first polynomial is univariate in $X$ one can compute its greatest common divisor with $X^{q}-X$ to solve for $X$ and henceforth also solve for $K_{1}$ and $K_{2}$.
-We note that this attack can also be interpreted in terms of term order conversion of our DRL Gröbner basis to LEX, since the univariate polynomial $\aleph=f_{1}(X)$ is simply the univariate polynomial of the Ciminion LEX Gröbner basis in $x$-shape position.
-Moreover, the construction of the univariate LEX polynomial can be effectively expressed in the cost of univariate polynomial multiplication over finite fields [HvdH22] which is quasi-linear in the degree.
-As countermeasure Bariant proposes to add a linear combination of $K_{1}$ and $K_{2}$ to each component after $p_{E}$.
-We note that this modification breaks our Gröbner basis from LABEL:Th:_Ciminion_Gr\"obner_basis.
-As settlement, we propose to add a linear combination of $K_{1}$ and $K_{2}$ only to the non-linear parts of the round functions, e.g.
-$\begin{split}\mathcal{R}^{(l)}:\mathbb{F}_{q}^{3}\times\mathbb{F}_{q}^{2}&\to\mathbb{F}_{q}^{3},\\ \begin{pmatrix}x\\ y\\ z\end{pmatrix},(K_{1},K_{2})&\mapsto\begin{pmatrix}0&0&1\\ 1&c_{4}^{(l)}&c_{4}^{(l)}\\ 0&0&1\end{pmatrix}\begin{pmatrix}x\\ y\\ z+x\cdot y+\alpha\cdot K_{1}+\beta\cdot K_{2}\end{pmatrix}+\begin{pmatrix}c_{1}^{(l)}\\ c_{2}^{(l)}\\ c_{3}^{(l)}\end{pmatrix},\end{split}$ (34)
-where $\alpha,\beta\in\mathbb{F}_{q}$ and $\alpha\cdot\beta\neq 0$.
-This modification does not affect the proof of LABEL:Th:_Ciminion_Gr\"obner_basis, hence we maintain the DRL Gröbner basis for the modified Ciminion polynomial system.
-However, this modification breaks the construction of the univariate LEX polynomial via univariate polynomial multiplication, because the DRL Gröbner basis polynomial for the $r_{C}$th round will now also contain linear terms for the variables $x_{C,3}^{(2)},\dots,x_{C,3}^{(r_{C}-2)}$ instead of only $x_{C,3}^{(r_{C}-1)},x_{E,3}^{(1)},\dots,x_{E,3}^{(r_{E})},x$.
-While the LEX Gröbner basis of the modified Ciminion might still be in $x$-shape position, we now can only construct it by running a FGLM algorithm.
-Note if the modification of Equation 34 is performed in all rounds of $p_{E}$, then the Middle State-Output Relation polynomial system of Bariant et al.
-[BBLP22, §5] becomes underdetermined, so this attack connecting two outputs of two nonce samples would not be applicable anymore.
-3.3 Implementation
-We implemented Ciminion in the computer algebra systems SageMath [Sag24] and OSCAR [OSC24].
-[^3] https://github.com/sca-research/Groebner-Basis-Cryptanalysis-of-Ciminion-and-Hydra.git In principle, the SageMath implementation can handle $128$ bit prime numbers, though we do not recommend using primes of that size for experimental Gröbner basis computations, because SageMath uses Python implementations for primes of that size.
-The OSCAR computer algebra system offers an efficient F4 implementation that can handle primes of size $q<2^{31}$, therefore we recommend to use OSCAR for experimental computations over small prime fields.
-We implemented the Gröbner basis computation from LABEL:Th:_Ciminion_Gr\"obner_basis in SageMath.
-For practical parameters $p=2^{127}+45$, $r_{C}=90$ and $r_{E}=14$ a DRL Gröbner basis can be computed in negligible time.
-3.3.1 Ciminion2
-We also implemented a modified version of Ciminion to protect against Bariant’s attack, see Equation 34, called Ciminion2, which adds $K_{1}+K_{2}$ in the last round of $p_{C}$ and every round of $p_{E}$ to the non-linear component.
-I.e., the round function modifies to
-$\mathcal{R}^{(l)}:\begin{pmatrix}x\\ y\\ z\end{pmatrix},\begin{pmatrix}K_{1}\\ K_{2}\end{pmatrix}\mapsto\begin{pmatrix}0&0&1\\ 1&c_{4}^{(l)}&c_{4}^{(l)}\\ 0&1&1\end{pmatrix}\begin{pmatrix}x\\ y\\ z+x\cdot y+K_{1}+K_{2}\end{pmatrix}+\begin{pmatrix}c_{1}^{(l)}\\ c_{2}^{(l)}\\ c_{3}^{(l)}\end{pmatrix},$ (35)
-where $r_{C}\leq l\leq r_{C}+r_{E}$. For Ciminion2 a DRL Gröbner basis is also computed in negligible time in SageMath for practical parameters.
-4 Analysis of Hydra
-The Hydra heads are based on the Lai–Massey permutation. For the Gröbner basis analysis we recall that the Lai–Massey permutation
-$\mathcal{L}:\begin{pmatrix}x\\ y\end{pmatrix}\mapsto\begin{pmatrix}x+F(x-y)\\ y+F(x-y)\end{pmatrix}$ (36)
-is affine equivalent to the Feistel permutation
-$\mathcal{F}:\begin{pmatrix}x\\ y\end{pmatrix}\mapsto\begin{pmatrix}x\\ y+F(x)\end{pmatrix}$ (37)
-via the matrices $\mathbf{A}_{1}=\begin{pmatrix}1&-1\\ 0&1\end{pmatrix}$ and $\mathbf{A}_{2}=\begin{pmatrix}1&1\\ 0&1\\ \end{pmatrix}$, i.e.
-$\mathcal{L}=\mathbf{A}_{2}\circ\mathcal{F}\circ\mathbf{A}_{1}$.
-Inspired by this equivalence one can reduce the Hydra polynomial system to only one non-linear equation per round, i.e.
-the Hydra polynomial system transforms to a Feistel-like polynomial system.
-In [Ste24, §6.2] the iterated polynomial systems representing block ciphers based on generalized Feistel networks were analyzed.
-For these polynomial systems linear systems were derived, and in case the systems are of full rank the iterated polynomial systems are in generic coordinates, see [Ste24, Theorem 6.5].
-Based on the aforementioned affine equivalence, we are going to replicate this strategy to construct such a linear system for Hydra.
-4.1 Generic Coordinates for Hydra
-Recall that the Hydra iterated polynomial system (Definition 2.2) consists of $816\cdot r_{\mathcal{H}}+$ 8 many polynomials in $416\cdot r_{\mathcal{H}}+$ 4 many variables. Let
-$\mathbf{A}_{n}=\begin{pmatrix}1&0&\ldots&0&-1\\ 0&1&\ldots&0&-1\\ &&\ddots&&\\ 0&0&\ldots&1&-1\\ 0&0&\ldots&0&1\end{pmatrix}\in\operatorname{GL}_{n}\left(\mathbb{F}_{q}\right),$ (38)
-and let $\mathbf{B}=\operatorname{diag}\left(\mathbf{A}_{4},\mathbf{A}_{4}\right)\in\operatorname{GL}_{8}\left(\mathbb{F}_{q}\right)$.
-Then the first seven components of $\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{j}^{(i)}$, where $1\leq i\leq r_{\mathcal{H}}$ and $j\in\{1,2\}$, are affine and the eighth component is quadratic.
-Moreover, the forth and the eighth components of $\mathbf{B}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{f}_{\mathcal{R}}$ are quadratic while the other ones are affine.
-In particular, by extracting the $2\cdot 7\cdot r_{\mathcal{H}}$ affine components of the $\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{j}^{(i)}$’s we can eliminate the variables
-$y_{1},\dots,y_{4},z_{1},\dots,z_{3},x_{j,1}^{(i)},\dots,x_{j,7}^{(i)},$ (39)
-where $10\leq i\leq r_{\mathcal{H}}-$ 1 and $j\in\{1,2\}$, in the quadratic equations.
-(This can be explicitly seen in the proof of Theorem 4.1.) Thus, after the elimination we reduce to $82\cdot r_{\mathcal{H}}+$ 8 polynomials, of whom $6$ are also affine, in $42\cdot r_{\mathcal{H}}+$ 4 many variables.
-We call this polynomial system the downsized Hydra polynomial system.
-Analog to the analysis of Feistel networks in [Ste24, §6.2], we now derive a linear system for Hydra with which we can verify whether the Hydra polynomial system is in generic coordinates.
-Theorem 4.1.
-Let $\mathbb{F}_{q}$ be a finite field, let $\overline{\mathbb{F}_{q}}$ be its algebraic closure, let
-$\mathcal{F}_{\textsf{Hydra}}=\left\{\mathbf{f}_{1}^{(i)},\mathbf{f}_{\mathcal{R}},\mathbf{f}_{2}^{(i)}\right\}_{1\leq i\leq r_{\mathcal{R}}}\subset P=\overline{\mathbb{F}_{q}}\Big{[}\mathbf{y},\mathbf{z},\mathbf{x}_{1}^{(1)},\dots,\mathbf{x}_{1}^{(r_{\mathcal{H}}-1)},\mathbf{x}_{2}^{(0)},\dots,\mathbf{x}_{2}^{(r_{\mathcal{H}}-1)},\mathbf{k}\Big{]}$
-be a Hydra polynomial system, and let
-$\mathcal{G}_{\textsf{Hydra}}=\left\{\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{1}^{(i)},\mathbf{B}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{f}_{\mathcal{R}},\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{2}^{(i)}\right\}_{1\leq i\leq r_{\mathcal{H}}}.$
-Then every homogeneous ideal $I\subset P[x_{0}]$ such that $\mathcal{Z}_{+}(I)\neq\emptyset$ and $\mathcal{G}_{\textsf{Hydra}}^{\text{hom}}\subset I$ is in generic coordinates if the following linear system has rank $416\cdot r_{\mathcal{H}}+$ 4
-$y_{1}-z_{4}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{1}^{(1)}\Big{)}\bigg{)}_{1}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{1}^{(1)}-\mathbf{k}^{\prime}\Big{)}\bigg{)}_{8}$ $=0,$
-$\vdots$
-$z_{3}-z_{4}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{1}^{(1)}\Big{)}\bigg{)}_{7}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{1}^{(1)}-\mathbf{k}^{\prime}\Big{)}\bigg{)}_{8}$ $=0,$
-$\sum_{l=1}^{4}y_{l}-z_{l}$ $=0,$
-$x_{2,k}^{(0)}-x_{2,8}^{(0)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{2}^{(1)}\Big{)}\bigg{)}_{k}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{2}^{(1)}-\mathbf{k}^{\prime}\Big{)}\bigg{)}_{8}$ $=0,$
-$\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{2,l}^{(0)}$ $=0,$
-$x_{j,k}^{(i)}-x_{j,8}^{(i)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{j}^{(i+1)}\Big{)}\bigg{)}_{k}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{j}^{(i+1)}-\mathbf{k}\Big{)}\bigg{)}_{8}$ $=0,$
-$\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{j,l}^{(i)}$ $=0,$
-$x_{j,k}^{(r_{\mathcal{H}}-1)}-x_{j,8}^{(r_{\mathcal{H}}-1)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+(\mathbf{y},\mathbf{z})^{\intercal}\Big{)}\bigg{)}_{k}-\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+(\mathbf{y},\mathbf{z})^{\intercal}\Big{)}\bigg{)}_{8}$ $=0,$
-$\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{j,l}^{(r_{\mathcal{H}}-1)}$ $=0,$
-$x_{j,k}^{(r_{\mathcal{H}}-1)}-x_{j,8}^{(r_{\mathcal{H}}-1)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+\mathbf{x}_{2}^{(0)}\Big{)}\bigg{)}_{k}-\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+\mathbf{x}_{2}^{(0)}\Big{)}\bigg{)}_{8}$ $=0,$
-$\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{j,l}^{(r_{\mathcal{H}}-1)}$ $=0,$
-$y_{1}-y_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{1}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{8}$ $=0,$
-$\vdots$
-$y_{3}-y_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{3}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{8}$ $=0,$
-$z_{1}-z_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{5}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{8}$ $=0,$
-$\vdots$
-$z_{3}-z_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{7}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{8}$ $=0,$
-where $12\leq i\leq r_{\mathcal{H}}-$ 1, $j\in\{1,2\}$ and $1\leq k\leq 7$.
-Proof.
-By [Ste24, Theorem 3.2] we have to show that
-$\sqrt{\mathcal{G}^{\text{top}}}=\left(\mathbf{y},\mathbf{z},\mathbf{x}_{j}^{(i)},\dots,\mathbf{x}_{1}^{(r_{\mathcal{H}}-1)},\mathbf{x}_{2}^{(0)},\dots,\mathbf{x}_{2}^{(r_{\mathcal{H}}-1)},\mathbf{k}\right)$
-To shorten some writing let
-$\mathbf{g}_{j}^{(i)}$ $=\left(\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{j}^{(i)}\right)^{\text{hom}}\mod x_{0},$
-$\mathbf{g}_{\mathcal{R}}$ $=\left(\mathbf{B}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{f}_{\mathcal{R}}\right)^{\text{hom}}\mod x_{0}.$
-Then for $i=1$ we have that
-$\mathbf{g}_{1}^{(1)}$ =($=\begin{pmatrix}\left\{y_{k}-z_{4}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{1}^{(1)}\Big{)}\bigg{)}_{k}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{1}^{(1)}-\mathbf{k}^{\prime}\Big{)}\bigg{)}_{8}\right\}_{1\leq k\leq 7}\\[10.0pt] \left(\sum_{l=1}^{4}y_{l}-z_{l}\right)^{2}\end{pmatrix},$
-$\mathbf{g}_{2}^{(1)}$ =($=\begin{pmatrix}\left\{x_{2,k}^{(0)}-x_{2,8}^{(0)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{2}^{(1)}\Big{)}\bigg{)}_{k}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{2}^{(1)}-\mathbf{k}^{\prime}\Big{)}\bigg{)}_{8}\right\}_{1\leq k\leq 7}\\[10.0pt] \left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{2,l}^{(0)}\right)^{2}\end{pmatrix}.$
-For $12\leq i\leq r_{\mathcal{H}}-$ 1 and $j\in\{1,2\}$ we have that
-$\mathbf{g}_{j}^{(i)}=\begin{pmatrix}\left\{x_{j,k}^{(i)}-x_{j,8}^{(i)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{j}^{(i+1)}\Big{)}\bigg{)}_{k}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{j}^{(i+1)}-\mathbf{k}\Big{)}\bigg{)}_{8}\right\}_{1\leq k\leq 7}\\[10.0pt] \left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{j,l}^{(i)}\right)^{2}\end{pmatrix}.$
-For $i=r_{\mathcal{H}}$ we have that
-$\mathbf{g}_{1}^{(r_{\mathcal{H}})}=$
-($\begin{pmatrix}\left\{x_{1,k}^{(r_{\mathcal{H}}-1)}-x_{1,8}^{(r_{\mathcal{H}}-1)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+(\mathbf{y},\mathbf{z})^{\intercal}\Big{)}\bigg{)}_{k}-\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+(\mathbf{y},\mathbf{z})^{\intercal}\Big{)}\bigg{)}_{8}\right\}_{1\leq k\leq 7}\\[10.0pt] \left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{1,l}^{(r_{\mathcal{H}}-1)}\right)^{2}\end{pmatrix},$
-$\mathbf{g}_{2}^{(r_{\mathcal{H}})}=$
-($\begin{pmatrix}\left\{x_{2,k}^{(r_{\mathcal{H}}-1)}-x_{2,8}^{(r_{\mathcal{H}}-1)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+\mathbf{x}_{2}^{(0)}\Big{)}\bigg{)}_{k}-\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+\mathbf{x}_{2}^{(0)}\Big{)}\bigg{)}_{8}\right\}_{1\leq k\leq 7}\\[10.0pt] \left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{2,l}^{(r_{\mathcal{H}}-1)}\right)^{2}\end{pmatrix}.$
-Finally, for the rolling function we have that
-$\mathbf{B}\mathbf{g}_{\mathcal{R}}=\begin{pmatrix}\left\{y_{k}-y_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{k}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{4}\right\}_{1\leq k\leq 3}\\[10.0pt] g(\mathbf{y},\mathbf{z})\\ \left\{z_{k}-z_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{4+k}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{8}\right\}_{1\leq k\leq 3}\\[10.0pt] g(\mathbf{z},\mathbf{y})\end{pmatrix}.$
-Except for the rolling function, all quadratic functions are powers of affine equations, so by Theorem 2.6 (2) we can replace the powers by the affine equations.
-This yields the linear system from the assertion.
-∎
-Naturally, this theorem can be extended to an arbitrary number of Hydra samples.
-It also follows by a simple contraposition that the downsized Hydra polynomial system in $42\cdot r_{\mathcal{H}}+$ 4 variables is in generic coordinates if the full polynomial system is in generic coordinates.
-Now it is straight-forward to compute the Macaulay bound (Theorem 2.5) for Hydra.
-Corollary 4.2.
-Let $\mathbb{F}_{q}$ be a finite field, let $\overline{\mathbb{F}_{q}}$ be its algebraic closure, and let
-~$\widetilde{\mathcal{F}}_{\textsf{Hydra}}\subset\overline{\mathbb{F}_{q}}\Big{[}z_{4},x_{1,8}^{(1)},\dots,x_{1,8}^{(r_{\mathcal{H}}-1)},x_{2,8}^{(0)},\dots,x_{2,8}^{(r_{\mathcal{H}}-1)},\mathbf{k}\Big{]}$
-be the downsized Hydra polynomial system. If ~$\widetilde{\mathcal{F}}_{\textsf{Hydra}}^{\text{hom}}$ is in generic coordinates, then
-$\operatorname{sd}_{DRL}\left(\widetilde{\mathcal{F}}_{\textsf{Hydra}}\right)\leq 2\cdot r_{\mathcal{H}}+3.$
-Remark 4.3.
-This bound can be further improved by utilizing the $6$ affine equations from the rolling function.
-In case that these $6$ equations are linearly independent of each other and all other affine equations, we can eliminate further $6$ variables.
-Thus, the Hydra polynomial system reduces to $22\cdot r_{\mathcal{H}}+$ 2 quadratic polynomials in $22\cdot r_{\mathcal{H}}-$ 2 many variables.
-Then
-$\operatorname{sd}_{DRL}\left(\widetilde{\mathcal{F}}_{\textsf{Hydra}}\right)\leq 2\cdot r_{\mathcal{H}}.$
-The reason why we did not incorporate this improvement in Theorem 2.5 is that elimination via the affine equations from the rolling function depends non-trivially on $\mathbf{M}_{\mathcal{R}}$ and $\mathbf{M}_{\mathcal{J}}$.
-Thus, a generic treatment is not possible.
-4.2 Extracting a Zero-Dimensional DRL Gröbner Basis for Hydra
-Under the assumption of Remark 4.3, we can extract a zero-dimensional Gröbner basis from the downsized Hydra polynomial system via a linear change of coordinates.
-Then, $2\cdot r_{\mathcal{H}}$ quadratic polynomials stem from Hydra heads and the remaining two from the rolling function.
-Recall that the quadratic terms in the heads are of the form
-$g_{1}^{(1)}$ $=\left(\sum_{l=1}^{4}y_{l}-z_{l}\right)^{2},\qquad$ $g_{2}^{(1)}=\left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{2,l}^{(0)}\right)^{2},$ (40)
-$g_{j}^{(i)}$ $=\left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{j,l}^{(i)}\right)^{2},$ $2\leq i\leq r_{\mathcal{H}}-1,\ j\in\{1,2\}$ (41)
-$g_{1}^{(r_{\mathcal{H}})}$ $=\left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{1,l}^{(r_{\mathcal{H}}-1)}\right)^{2},\qquad$ $g_{2}^{(r_{\mathcal{H}})}=\left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{2,l}^{(r_{\mathcal{H}}-1)}\right)^{2},$ (42)
-i.e.
-they are the squares of linear combinations of the variables.
-Therefore, we collect these linear combinations in a matrix $\mathbf{M}\in\mathbb{F}_{q}^{2\cdot r_{\mathcal{H}}\times(2\cdot r_{\mathcal{H}}-2)}$ analog to the linear system in Theorem 4.1 to verify generic coordinates for Hydra.
-If $\operatorname{rank}\left(\mathbf{M}\right)=2\cdot r_{\mathcal{H}}-2$, then we can perform a linear change of coordinates to new variables $\hat{x}_{1},\dots,\hat{x}_{2\cdot r_{\mathcal{H}}-2}$.
-After the change of coordinates the Hydra polynomial system contains $22\cdot r_{\mathcal{H}}-$ 2 polynomials of the form
-$g_{i}=\hat{x}_{i}^{2}+\sum_{j=1}^{2\cdot r_{\mathcal{H}}-2}\alpha_{i,j}\cdot\hat{x}_{j}+\alpha_{i,0}.$ (43)
-Trivially, $\mathcal{G}_{\textsf{Hydra}}=\{g_{i}\}_{1\leq i\leq 2\cdot r_{\mathcal{H}}-2}\subset\mathbb{F}_{q}\big{[}\hat{x}_{1},\dots,\hat{x}_{2\cdot r_{\mathcal{H}}-2}\big{]}$ forms a zero-dimensional DRL Gröbner basis, see [CLO15, Chapter 2 §9 Theorem 3, Proposition 4].
-In particular, for the $\mathbb{F}_{q}$-vector space dimension of the quotient ring we then have that
-$\dim_{\mathbb{F}_{q}}\left(\mathcal{G}_{\textsf{Hydra}}\right)=2^{2\cdot r_{\mathcal{H}}-2}.$ (44)
-We also note that after an auxiliary division by remainder computation, the additional four polynomials are of the form
-$h_{j}=\sum_{1\leq k<l\leq 2\cdot r_{\mathcal{H}}-2}\alpha_{i,k,l}\cdot\hat{x}_{k}\cdot\hat{x}_{l}+\sum_{k=1}^{2\cdot r_{\mathcal{H}}-2}\beta_{i,k}\cdot\hat{x}_{k}+\beta_{i,0}.$ (45)
-4.3 Implementation
-We have implemented Hydra and its iterated polynomial system in the computer algebra system SageMath [Sag24] and OSCAR [OSC24].
-[^4] https://github.com/sca-research/Groebner-Basis-Cryptanalysis-of-Ciminion-and-Hydra.git In principle, the SageMath implementation can handle $128$ bit prime numbers, though we do not recommend using primes of that size for experimental Gröbner basis computations, because SageMath uses Python implementations for primes of that size.
-The OSCAR computer algebra system offers an efficient F4 implementation that can handle primes of size $q<2^{31}$, therefore we recommend to use OSCAR for experimental computations over small prime fields.
-Moreover, we implemented the linear system from Theorem 4.1 in SageMath to verify whether a Hydra instance is in generic coordinates.
-We stress that our generic coordinates test can handle practical parameters like $128$ bit prime numbers.
-Also, our verification function can handle an arbitrary number of Hydra samples.
-In case a Hydra polynomial system is in generic coordinates, we implemented the change of coordinates from Section 4.2 for two samples to extract a zero-dimensional DRL Gröbner basis from the Hydra polynomial system.
-4.3.1 A Concrete Instance
-In [GØSW22, Appendix C] the Hydra designers propose concrete matrices over the prime field $q=2^{127}+45$
-$\mathbf{M}_{\mathcal{E}}$ $=\operatorname{circ}\left(3,2,1,1\right),$ (46)
-$\mathbf{M}_{\mathcal{I}}$ $=\begin{pmatrix}1&1&1&1\\ 1&4&1&1\\ 3&1&3&1\\ 4&1&1&2\end{pmatrix},$ (47)
-$\mathbf{M}_{\mathcal{J}}$ $=\begin{pmatrix}3&1&1&1&1&1&1&1\\ 7&3&1&1&1&1&1&1\\ 4&1&4&1&1&1&1&1\\ 3&1&1&8&1&1&1&1\\ 7&1&1&1&7&1&1&1\\ 8&1&1&1&1&5&1&1\\ 5&1&1&1&1&1&2&1\\ 4&1&1&1&1&1&1&6\end{pmatrix}.$ (48)
-For security level $\kappa=128$ the designers suggest to use $r_{\mathcal{H}}=39$, see [GØSW23, §5.5].
-For this instance using our SageMath implementation we verified that the Hydra polynomial system is in generic coordinates for $m\in\{2,3,4,5,10\}$ many samples.
-For $m=2$ we can compute a Hydra DRL Gröbner basis with the change of coordinates from Section 4.2.
-In particular, the polynomial system then consists of $22\cdot r_{\mathcal{H}}+$ 2 polynomials in $22\cdot r_{\mathcal{H}}-$ 2 variables, so we can use the slightly improved solving degree bound from Remark 4.3 to estimate the complexity of DRL Gröbner basis computations.
-On an Intel Core i7-10700 with 64 GB RAM verification of generic coordinates for $m=2$ needs $\leq 10\,\text{s}$ and for $m=10$ it needs $\leq 4\,\text{min}$.
-In addition, for $r_{\mathcal{H}}=39$, to extract a Gröbner basis let $\mathbf{M}\in\mathbb{F}_{q}^{2\cdot r_{\mathcal{H}}\times(2\cdot r_{\mathcal{H}}-2)}$ collect the linear combinations in the quadratic terms of the downsized Hydra polynomial system.
-Then, the rows with indices $3,\dots,2\cdot r_{\mathcal{H}}$ are linearly independent.
-We also note that we did not find a round number $r_{\mathcal{H}}\geq 2$ for which the Gröbner basis extraction fails.
-5 Shaving-Off Bits with the Eigenvalue Method
-Let =$\mathcal{F}=\{f_{1},\dots,f_{n}\}\subset\mathbb{F}_{q}[x_{1},\dots,x_{n}]$ be such that $\operatorname{LM}_{DRL}\left(f_{i}\right)=x_{i}^{2}$ for all $i$, then $\mathcal{F}$ is a zero-dimensional Gröbner basis.
-Moreover, for Ciminion and Hydra we found Gröbner bases that satisfy this assumption.
-In Section 2.5 we already discussed that one can solve for the zero-locus of $\mathcal{F}$ by computing the eigenvalues of the multiplication matrices while maintaining the complexity estimate of the probabilistic FGLM algorithm [FGHR14].
-We will now discuss that we can slightly improve the complexity estimate by taking the multiplicative structure of $\mathbb{F}_{q}[x_{1},\dots,x_{n}]/\left(\mathcal{F}\right)$ into account.
-Recall that a $\mathbb{F}_{q}$-vector space basis of $\mathbb{F}_{q}[x_{1},\dots,x_{n}]/\left(\mathcal{F}\right)$ is given by the monomials not contained in $\operatorname{in}_{DRL}\left(\mathcal{F}\right)=\left(x_{1}^{2},\dots,x_{n}^{2}\right)$.
-Let $\mathcal{B}^{\prime}$ be the monomials not contained in $\left(x_{1}^{2},\dots,x_{n-1}^{2}\right)\subset\mathbb{F}_{q}[x_{1},\dots,x_{n-1}]$, then we sort the quotient space basis as =$\mathcal{B}=\left\{\mathcal{B}^{\prime},x_{n}\cdot\mathcal{B}^{\prime}\right\}$.
-The multiplication matrix for $x_{n}$ is then of the form
-[…]
-where $\mathbf{A},\mathbf{B}\in\mathbb{F}_{q}^{2^{n-1}\times 2^{n-1}}$. By definition, the characteristic polynomial is given by the determinant
-$\chi_{x_{n}}(x)=\det\left(\mathbf{I}_{2^{n}\times 2^{n}}\cdot x-\mathbf{M}_{x_{n}}\right)=\det\begin{pmatrix}\mathbf{I}_{2^{n-1}\times 2^{n-1}}\cdot x&-\mathbf{I}_{2^{n-1}\times 2^{n-1}}\\ \mathbf{A}&\mathbf{I}_{2^{n-1}\times 2^{n-1}}\cdot x-\mathbf{B}\end{pmatrix}.$ (50)
-For the LU-decomposition from Equation 24, we lift the determinant computation from the polynomial ring $\mathbb{F}_{q}[x]$ to its fraction field $\mathbb{F}_{q}(x)$, i.e.
-we treat $x$ as invertible.
-Then with Equation 25
-$\chi_{x_{n}}(x)$ $=\det\begin{pmatrix}\mathbf{I}_{2^{n-1}\times 2^{n-1}}\cdot x&-\mathbf{I}_{2^{n-1}\times 2^{n-1}}\\ -\mathbf{A}&\mathbf{I}_{2^{n-1}\times 2^{n-1}}\cdot x-\mathbf{B}\end{pmatrix}$ (51)
-$=\det\left(\mathbf{I}_{2^{n-1}\times 2^{n-1}}\cdot x\right)\cdot\det\left(\mathbf{I}_{2^{n-1}\times 2^{n-1}}\cdot x-\mathbf{B}-\mathbf{A}\cdot x^{-1}\right)$ (52)
-$=x^{2^{n-1}}\cdot\det\left(\mathbf{I}_{2^{n-1}\times 2^{n-1}}\cdot x-\mathbf{B}-\mathbf{A}\cdot x^{-1}\right)$ (53)
-$=\det\left(\mathbf{I}_{2^{n-1}\times 2^{n-1}}\cdot x^{2}-\mathbf{B}\cdot x-\mathbf{A}\right),$ (54)
-where the latter equality follows from the fact that the determinant is a homogeneous function.
-We note that a similar determinant trick has been applied in [BBL+24, §3.2] to improve characteristic polynomial computations for specially shaped Gröbner bases.
-For the next variable $x_{n-1}$, we extract a $\mathbb{F}_{q}$-valued solution for $x_{n}$ via the GCD method, substitute it into $\mathcal{F}\setminus\left\{f_{n}\right\}$, and repeat the trick for the multiplication matrix, see Equation 49, with one variable less.
-Let $D(d,q)$ denote the complexity of the GCD with the field equation, see Equation 1.
-Provided that every characteristic polynomial has at most $N$ roots in $\mathbb{F}_{q}$, then the complexity of the Eigenvalue method with standard determinant complexity is given by
-$\mathcal{O}\Bigg{(}\sum_{i=1}^{n}N^{n-i}\cdot\Big{(}2^{\omega\cdot i}+D\big{(}2^{i},q\big{)}\Big{)}\Bigg{)}=\mathcal{O}\Bigg{(}2^{\omega}\cdot\frac{2^{\omega\cdot n}-N^{n}}{2^{\omega}-N}+\sum_{i=1}^{n}N^{n-i}\cdot D\big{(}2^{i},q\big{)}\bigg{)}\Bigg{)}.$ (55)
-With the determinant trick from Equation 54 the complexity for the Eigenvalue Method then improves to
-$\mathcal{O}\Bigg{(}\sum_{i=1}^{n}N^{n-i}\cdot\Big{(}2^{\omega\cdot(i-1)}+D\big{(}2^{i},q\big{)}\Big{)}\Bigg{)}=\mathcal{O}\Bigg{(}\frac{2^{\omega\cdot n}-N^{n}}{2^{\omega}-N}+\sum_{i=1}^{n}N^{n-i}\cdot D\big{(}2^{i},q\big{)}\bigg{)}\Bigg{)}.$ (56)
-In addition, for $N=1$, this implies by $\omega\geq 2$ that
-$\log_{2}\left(\frac{2^{\omega\cdot n}-1}{2^{\omega}-1}\right)\leq\omega\cdot n-\log_{2}\left(3\right).$ (57)
-I.e., we expect an improvement of at least $\log_{2}\left(3\right)\approx 1.5849$ compared to generic probabilistic term order conversion for an ideal in shape position.
-By imposing more structure on $\mathcal{F}$ we can refine Equation 56 even further. Suppose that every $f_{i}\in\mathcal{F}$ is of the form
-$f_{i}=x_{i}^{2}+\mathcal{A}_{i}(x_{1},\dots,x_{n}),$ (58)
-where $\mathcal{A}_{i}$ is an affine polynomial. Assume we have found a solution for $x_{n}$, say $x_{n}=\alpha_{n}$, then we can rearrange
-$\beta_{n-1}\cdot x_{n-1}=\alpha_{n}^{2}+\tilde{\mathcal{A}}_{i}(x_{1},\dots,x_{n-2},\alpha_{n}).$ (59)
-(If $\beta_{n-1}=0$, then we simply rename the variables.) We can then further eliminate the variable $x_{n-1}$ in the polynomial system $\mathcal{F}\setminus\{f_{n-1},f_{n}\}\subset\mathbb{F}_{q}[x_{1},\dots,x_{n}]$ by substituting the previous affine equation for $x_{n-1}$ and $x_{n}=\alpha_{n}$.
-I.e., by finding one eigenvalue we can eliminate two variables.
-We also note that the Hydra DRL Gröbner basis (Section 4.2) satisfies the shape of Equation 58.
-Then, by iterating this variable elimination Equation 56 improves to
-$\mathcal{O}\left(\begin{rcases}\begin{dcases}\sum_{i=1}^{\frac{n}{2}}N^{n-2\cdot i}\cdot\Big{(}2^{\omega\cdot(2\cdot i-1)}+D\big{(}2^{2\cdot i},q\big{)}\Big{)},&2\mid n,\\ \sum_{i=1}^{\frac{n-1}{2}}N^{n-1-2\cdot i}\cdot\Big{(}2^{\omega\cdot 2\cdot i}+D\big{(}2^{2\cdot i+1},q\big{)}\Big{)},&2\nmid n\end{dcases}\end{rcases}\right)$ (60)
-$=\mathcal{O}\left(\begin{rcases}\begin{dcases}2^{\omega}\cdot\frac{2^{\omega\cdot n}-N^{n}}{2^{2\cdot\omega}-N^{2}}+\sum_{i=1}^{\frac{n}{2}}N^{n-2\cdot i}\cdot D\big{(}2^{2\cdot i},q\big{)},&2\mid n,\\ 2^{2\cdot\omega}\cdot\frac{2^{\omega\cdot(n-1)}-N^{n}}{2^{2\cdot\omega}-N^{2}}+\sum_{i=1}^{\frac{n-1}{2}}N^{n-1-2\cdot i}\cdot D\big{(}2^{2\cdot i+1},q\big{)},&2\nmid n\end{dcases}\end{rcases}\right),$ (61)
-where we again assume that every characteristic polynomial has at most $N$ roots in $\mathbb{F}_{q}$.
-Again for $N=1$, this implies that
-$\log_{2}\left(\begin{rcases}\begin{dcases}2^{\omega}\cdot\frac{2^{\omega\cdot n}-1}{2^{2\cdot\omega}-1},&2\mid n,\\ 2^{2\cdot\omega}\cdot\frac{2^{\omega\cdot(n-1)}-1}{2^{2\cdot\omega}-1},&2\nmid n\end{dcases}\end{rcases}\right)$ $\leq\omega\left(n+1\right)-\log_{2}\left(2^{2\cdot\omega}-1\right)$ (62)
-$\approx\begin{dcases}2\cdot n-1.9069,&\omega=2,\\ 3\cdot n-2.9773,&\omega=3.\end{dcases}$ (63)
-I.e., we expect an improvement of $\approx\omega$ compared to generic probabilistic term order conversion.
-6 Boolean Macaulay Matrices
+As already mentioned, polynomial systems in generic coordinates have a proven DRL solving degree upper bound.
+Theorem 2.6 ([CG21, Theorem 9, 10, Corollary 2]).
+Let $K$ be an algebraically closed field, and let =$\mathcal{F}=\{f_{1},\dots,f_{m}\}\subset K[x_{1},\dots,x_{n}]$ be an inhomogeneous polynomial system such that $\left(\mathcal{F}^{\text{hom}}\right)$ is in generic coordinates.
+Let $d_{i}=\deg\left(f_{i}\right)$ and assume that $d_{1}\geq\ldots\geq d_{m}$, then for $l\in\min\{n+1,m\}$
+$\operatorname{sd}_{DRL}\left(\mathcal{F}\right)\leq d_{1}+\ldots+d_{l}-l+1.$
+In the Gröbner basis and algebraic cryptanalysis literature, the latter inequality is often referred to as the Macaulay bound.
+2.3.3 Complexity Estimates via the Solving Degree
+It is well-known that the complexity of Gaussian elimination of a matrix $\mathbf{A}\in K^{n\times m}$ with $r=\operatorname{rank}\left(\mathbf{A}\right)$ is given by $\mathcal{O}\left(n\cdot m\cdot r^{\omega-2}\right)$ field operations [Sto00, §2.2], where $2.3715522\leq\omega<$ 2.371552 is a linear algebra constant [WXXZ24].
+Given an inhomogeneous polynomial system $\mathcal{F}\subset P=K[x_{1},\dots,x_{n}]$, it is well-known that the number of monomials in $P$ of degree $d$ is $\binom{n+d-1}{d}$.
+Thus, the number of columns of $M_{\leq d}$ is bounded by $\sum_{i=0}^{d}\binom{n+i-1}{i}=\binom{n+d}{d}$, and the number of rows by $\left|\mathcal{F}\right|\cdot\binom{n+d}{d}$.
+So, the complexity of Gaussian elimination on the inhomogeneous Macaulay matrices $M_{\leq d}$ is bounded by
+$\mathcal{O}\left(\left|\mathcal{F}\right|\cdot\binom{n+d}{d}^{\omega}\right).$ (14)
+2.4 Boolean Macaulay Matrices
 Let $\mathcal{F}\subset\mathbb{F}_{2}[x_{1},\dots,x_{n}]$ be a polynomial system.
-If we are only interested in $\mathbb{F}_{2}$-valued solutions, then it is sufficient to compute a Gröbner basis of the ideal $\left(\mathcal{F},x_{1}^{2}-x_{1},\dots,x_{n}^{2}-x_{n}\right)$.
+If we are only interested in $\mathbb{F}_{2}$-valued solutions, then it is sufficient to compute a Gröbner basis of the ideal $\big{(}\mathcal{F},x_{1}^{2}-x_{1},\dots,x_{n}^{2}-x_{n}\big{)}$.
 On the other hand, via the field equations $F=\left(x_{1}^{2}-x_{1},\dots,x_{n}^{2}-x_{n}\right)$ we can reduce any monomial in the Macaulay matrix to a square-free monomial.
 Effectively, this reduces the full Macaulay matrix to a square-free submatrix, the so-called Boolean Macaulay matrix.
-Gaussian elimination for Boolean DRL Macaulay matrices over $\mathbb{F}_{2}$ was analyzed in [BFSS13] by Bardet et al.
-Interestingly, for Ciminion & Hydra we found zero-dimensional DRL Gröbner bases with leading monomials $x_{i}^{2}$.
+Gaussian elimination for Boolean Macaulay matrices with respect to DRL was analyzed in [BFSS13] by Bardet et al.
+Interestingly, for Ciminion and Hydra we will construct zero-dimensional DRL Gröbner bases in Sections 3.1 and 4.2 with leading monomials $x_{i}^{2}$.
 Hence, for Gröbner basis computations we can use these bases to reduce the Macaulay matrix to a “Boolean” one over an arbitrary finite field $\mathbb{F}_{q}$.
@@
     $\mathcal{F}$ is a $>$-Gröbner basis.
+Note that Condition (ii) is introduced so that the remainder of multivariate polynomial division does not depend on the ordering of the elements of $\mathcal{F}$, see [CLO15, Chapter 2 §6 Proposition 1].
 Definition 2.8 (cf. [BFSS13, Definition 1]).
@@
 Assume that $\mathcal{F}_{\text{Bool}}$ is $K$-Boolean with respect to $>$.
-For the (inhomogeneous) $K$-Boolean Macaulay matrix $M_{\leq d}^{\text{Bool}}$:
-  • Columns are indexed by square-free monomials $s\in P\setminus\operatorname{in}_{>}\left(\mathcal{F}_{\text{Bool}}\right)$ with $\deg\left(m\right)\leq d$ sorted via $>$.
-  • Rows are index by polynomials $t\cdot f$ where $t\in P\setminus\operatorname{in}_{>}\left(\mathcal{F}_{\text{Bool}}\right)$ is a square-free monomial and $f\in\mathcal{F}$ such that $\deg\left(t\cdot f\right)\leq d$.
+For the (inhomogeneous) $K$-Boolean Macaulay matrix $M_{\leq d}^{\text{Bool}}$ of $\mathcal{F}\cup\mathcal{F}_{\text{Bool}}$:
+  • Columns are indexed by square-free monomials $s\in P\setminus\operatorname{LM}_{>}\left(\mathcal{F}_{\text{Bool}}\right)$ with $\deg\left(m\right)\leq d$ sorted via $>$.
+  • Rows are indexed by polynomials $t\cdot f$ where $t\in P\setminus\operatorname{LM}_{>}\left(\mathcal{F}_{\text{Bool}}\right)$ is a square-free monomial and $f\in\mathcal{F}$ such that $\deg\left(t\cdot f\right)\leq d$.
   • In the row $t\cdot f$ one writes the coefficients of the remainder $r_{t\cdot f}=t\cdot f\mod\left(\mathcal{F}_{\text{Bool}}\right)$ with respect to $>$.
-Note that any monomial $t\in P$ can be reduced to a square-free monomial $\hat{t}\equiv t\mod\left(\mathcal{F}_{\text{Bool}}\right)$ with $\deg\left(\hat{t}\right)\leq\deg\left(t\right)$ due to degree compatibility of $>$.
+Note that any monomial $t\in P$ can be reduced to a square-free polynomial $\hat{t}\equiv t\mod\left(\mathcal{F}_{\text{Bool}}\right)$ with $\deg\left(\hat{t}\right)\leq\deg\left(t\right)$ due to degree compatibility of $>$.
 Since remainders with respect to a Gröbner basis are unique, see [CLO15, Chapter 2 §6 Proposition 1], for the remainder of $t\cdot f$ it does not matter whether we reduce $t$ first and then $\hat{t}\cdot f$ or $t\cdot f$ at once.
 So it is indeed sufficient to consider only products with square-free monomials.
-Obviously, the remainder $r_{t\cdot f}=t\cdot f\mod\left(\mathcal{F}_{\text{Bool}}\right)$ is polynomial consisting of square-free terms.
-Moreover, by elementary properties of multivariate division by remainder with respect to a degree compatible term order, see [CLO15, Chapter 2 §3 Theorem 3], we have that
+Obviously, the remainder $r_{t\cdot f}=t\cdot f\mod\left(\mathcal{F}_{\text{Bool}}\right)$ is a polynomial consisting of only square-free terms.
+Moreover, by elementary properties of multivariate polynomial division with respect to a degree compatible term order, see [CLO15, Chapter 2 §3 Theorem 3], we have that
 $t\cdot f=\sum_{i=1}^{m}g_{i}\cdot f_{i}^{\text{Bool}}+r_{t\cdot f},$ (15)
@@
 I.e., the $K$-Boolean Macaulay matrix can be constructed via row operations from the Macaulay matrix $M_{\leq d}$ of $\mathcal{F}\cup\mathcal{F}_{\text{Bool}}$.
-Alternatively, we can obtain the $K$-Boolean Macaulay matrix by pivoting non-square-free monomials in $M_{\leq d}$ via rows coming from $\mathcal{F}_{\text{Bool}}$, and then performing an auxiliary Gaussian elimination step that stops after all redundant pivots coming from $\mathcal{F}_{\text{Bool}}$ have been eliminated.
+Alternatively, we can obtain the $K$-Boolean Macaulay matrix by pivoting non-square-free monomials in $M_{\leq d}$ via rows coming from $\mathcal{F}_{\text{Bool}}$, and then performing an auxiliary Gaussian elimination step that stops after eliminating all non-square-free monomials in the rows coming from $\mathcal{F}$.
 Let $M_{\leq d}$ be a Macaulay matrix, we denote with $\operatorname{rowsp}\left(M_{\leq d}\right)$ the row space basis, i.e.
 the polynomial system obtained via Gaussian elimination on $M_{\leq d}$.
-Analog, we use this notation for Boolean Macaulay matrices.
-Definition 6.3 (Boolean Solving Degree).
+Analogously, we use this notation for Boolean Macaulay matrices.
+Definition 2.9.
 Let $K$ be a field, let $\mathcal{F}_{\text{Bool}},\mathcal{F}\subset P=K[x_{1},\dots,\allowbreak x_{n}]$ be polynomial systems, and let $>$ be a degree compatible term order on $P$.
 Assume that $\mathcal{F}_{\text{Bool}}$ is $K$-Boolean with respect to $>$.
-The Boolean solving degree of $\mathcal{F}\cup\mathcal{F}_{\text{Bool}}$ is the least degree $d$ such that $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)\cup\mathcal{F}_{\text{Bool}}$ contains a Gröbner basis with respect to $>$.
+The $K$-Boolean solving degree of $\mathcal{F}\cup\mathcal{F}_{\text{Bool}}$ is the least degree $d$ such that $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)\cup\mathcal{F}_{\text{Bool}}$ contains a Gröbner basis with respect to $>$.
 We denote it by $\operatorname{sd}^{\text{Bool}}_{>}\left(\mathcal{F}\cup\mathcal{F}_{\text{Bool}}\right)$.
-Note that $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)$ might not contain zero-dimensional Gröbner basis of the full system.
+Note that $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)$ might not contain a zero-dimensional Gröbner basis of the full system.
 A trivial example is =$\mathcal{F}=\left\{x_{1}^{2},x_{1}\cdot x_{2},x_{2}^{2}\right\}$.
 Therefore, to extract the Gröbner basis from $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)\cup\mathcal{F}_{\text{Bool}}$ we set up the ideal of $>$-leading terms and then compute its minimal generating set.
-With the minimal generating set we can then filter the Gröbner basis.
-(It is also obvious that all non-square-free monomials are generated by the leading terms of $\mathcal{F}_{\text{Bool}}$, hence it is indeed sufficient to just consider $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)\cup\mathcal{F}_{\text{Bool}}$.)
+With the minimal generating set we can then filter the $>$-Gröbner basis.444 It is obvious that all leading terms of $\operatorname{rowsp}\left(M_{\leq d}\right)$ not coming from $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)$ are already contained in $\operatorname{LM}_{>}\left(\mathcal{F}_{\text{Bool}}\right)$.
+Since $\mathcal{F}_{\text{Bool}}$ is a $>$-Gröbner basis it is indeed sufficient to just consider $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)\cup\mathcal{F}_{\text{Bool}}$.
 The next theorem follows straight-forward from $M_{\leq d}^{\text{Bool}}\subset M_{\leq d}$ (as $K$-vector spaces).
@@
 Let $K$ be a field, let $\mathcal{F}_{\text{Bool}},\mathcal{F}\subset P=K[x_{1},\dots,x_{n}]$ be polynomial systems, and let $>$ be a degree compatible term order on $P$.
+Let $D=\min_{f\in\mathcal{F}}\deg\left(f\right)$, and let $d=\operatorname{sd}^{\text{Bool}}_{>}\left(\mathcal{F}\cup\mathcal{F}_{\text{Bool}}\right)$.
 Assume that $\mathcal{F}_{\text{Bool}}$ is $K$-Boolean with respect to $>$.
@@
   2. (2)
-    If $d=\operatorname{sd}^{\text{Bool}}_{>}\left(\mathcal{F}\cup\mathcal{F}_{\text{Bool}}\right)<n$, then Gaussian elimination on $M_{\leq d}^{\text{Bool}}$ requires
-    $\mathcal{O}\left(\left|\mathcal{F}\right|\cdot\left(\sum_{i=1}^{d}\binom{n}{i}\right)^{\omega}\right)$
+    If $d-D<n$, then Gaussian elimination on $M_{\leq d}^{\text{Bool}}$ requires
+    $\mathcal{O}\left(\left|\mathcal{F}\right|\cdot\left(\sum_{i=0}^{d-D}\binom{n}{i}\right)\cdot\left(\sum_{i=0}^{d}\binom{n}{i}\right)^{\omega-1}\right)$
     field operations.
   3. (3)
-    If $d=\operatorname{sd}_{>}\left(\mathcal{F}\cup\mathcal{F}_{\text{Bool}}\right)\geq n$, then Gaussian elimination on $M_{\leq d}^{\text{Bool}}$ requires at most
+    If $d-D\geq n$, then Gaussian elimination on $M_{\leq d}^{\text{Bool}}$ requires at most
     $\mathcal{O}\left(\left|\mathcal{F}\right|\cdot 2^{\omega\cdot n}\right)$
-    field operations to produce a $>$-Gröbner basis.
+    field operations.
 Proof.
-For (1), we pivot non-square-free monomials via polynomials from $\mathcal{F}_{\text{Bool}}$, then we perform the auxiliary Gaussian elimination step that produces $M_{\leq d}^{\text{Bool}}$ as submatrix, and finally we perform Gaussian elimination on $M_{\leq d}^{\text{Bool}}$.
+For (1), in $M_{\leq d}$ we pivot non-square-free columns via rows coming from $\mathcal{F}_{\text{Bool}}$, then we perform the auxiliary Gaussian elimination step that produces $M_{\leq d}^{\text{Bool}}$ as submatrix, and finally we perform Gaussian elimination on $M_{\leq d}^{\text{Bool}}$.
 Since Gaussian elimination does not depend on the order of reductions $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)\cup\mathcal{F}_{\text{Bool}}$ has to contain the $>$-Gröbner basis.
-For (2), trivially in degree $i$ there exist $\binom{n}{i}$ square-free monomials, and every $f\in\mathcal{F}$ can contribute to at most $\binom{n}{i}$ rows.
-I.e., the number of columns is bounded by $D=\sum_{i=1}^{d}\binom{n}{i}$, and the number of rows by $\left|\mathcal{F}\right|\cdot D$.
-The complexity of Gaussian elimination then follows from [Sto00, §2.2].
-For (3), let $s\in P$ be a non-square-free monomial with $\deg\left(s\right)>n$, then $\hat{s}\equiv s\mod\left(\mathcal{F}_{\text{Bool}}\right)$ where $\hat{s}\in P$ is square-free, in particular $\deg\left(\hat{s}\right)\leq n$.
-Therefore, is suffices to perform Gaussian elimination on the full $K$-Boolean Macaulay matrix.
-Moreover, it is well known that $\sum_{i=0}^{n}\binom{n}{i}=2^{n}$.
+For (2), it is well-known that in degree $i$ there exist $\binom{n}{i}$ square-free monomials in $P$.
+The polynomial $f\in\mathcal{F}$ contributes $\sum_{i=0}^{d-\deg\left(f\right)}\binom{n}{i}$ many rows to the $K$-Boolean Macaulay matrix.
+Then, the total number of rows in $M_{\leq d}^{\text{Bool}}$ is bounded by
+$\sum_{f\in\mathcal{F}}\sum_{i=0}^{d-\deg\left(f\right)}\binom{n}{i}\leq\sum_{f\in\mathcal{F}}\sum_{i=0}^{d-D}\binom{n}{i}=\left|F\right|\cdot\sum_{i=0}^{d-D}\binom{n}{i}.$
+The number of columns is bounded by $\sum_{i=0}^{d}\binom{n}{i}$, so we also have that $\operatorname{rank}\left(M_{\leq d}^{\text{Bool}}\right)\leq\sum_{i=0}^{d}\binom{n}{i}$, and the claim follows from the complexity of Gaussian elimination [Sto00, §2.2].
+For (3), if $d-D\geq n$, then there exists at least one polynomial $f\in\mathcal{F}$ which has to be multiplied with any square-free monomial $s\in P$.
+So, $f$ contributes $\sum_{i=0}^{n}\binom{n}{i}=2^{n}$ many rows.
+Obviously, any other $g\in\mathcal{F}$ contributes at most $\leq 2^{n}$ many rows.
+Hence, the total number of rows is bounded by $\left|\mathcal{F}\right|\cdot 2^{n}$, and the number of columns as well as the rank are bounded by $2^{n}$.
+Again, the claim follows by [Sto00, §2.2].
 ∎
 Remark 2.11.
-This theorem can be refined by replacing the binomial coefficient $\binom{n}{i}$ by the number of monomials in $P\setminus\operatorname{in}_{>}\left(\mathcal{F}_{\text{Bool}}\right)$ of degree $i$.
+This theorem can be refined by replacing the binomial coefficients $\binom{n}{i}$ with the number of monomials in $P\setminus\operatorname{LM}_{>}\left(\mathcal{F}_{\text{Bool}}\right)$ of degree $i$.
 So far our analysis does not differ from the one over $\mathbb{F}_{2}$, however the binary field has a special feature.
-Division by remainder with respect to $\left(x_{1}^{2}-x_{1},\dots,x_{n}^{2}-x_{n}\right)$ is “for free”, i.e.
+Division with respect to the field equations $\left(x_{1}^{2}-x_{1},\dots,x_{n}^{2}-x_{n}\right)$ is “for free”, i.e.
 its complexity is negligible compared to Gaussian elimination.
 For general $K$-Boolean polynomial systems this is not necessarily the case.
-Luckily, the complexity of general division by remainder with respect to quadratic $K$-Boolean polynomial systems follows as corollary to [Ten19, Proposition 3.10].
+The complexity of general polynomial division with respect to quadratic $K$-Boolean polynomial systems follows as corollary to [Ten19, Proposition 3.10].
 Corollary 2.12.
-Let $K$ be a field, let $\mathcal{F}_{\text{Bool}}\subset P=K[x_{1},\dots,x_{n}]$ be a quadratic polynomial system, let $f\in P$ be a polynomial, and let $>$ be a degree compatible term order on $P$.
+Let $K$ be a field, let $\mathcal{F}_{\text{Bool}}\subset P=K[x_{1},\dots,x_{n}]$ be a quadratic polynomial system, and let $>$ be a degree compatible term order on $P$.
 Assume that $\mathcal{F}_{\text{Bool}}$ is $K$-Boolean with respect to $>$.
-Then computation of $r_{f}=f\mod\left(\mathcal{F}_{\text{Bool}}\right)$ with respect to $>$ requires
-$\mathcal{O}\left(\frac{n^{2}+3\cdot n+2}{2}\cdot\sum_{i=0}^{\deg\left(f\right)}\binom{n+i-1}{i}\right)$
-$\in\mathcal{O}\left(\frac{n^{2}+3\cdot n+2}{2}\cdot\deg\left(f\right)\cdot\binom{n+\deg\left(f\right)-1}{\deg\left(f\right)}\right)$
+  1. (1)
+    Let $f\in P$, then computation of $r_{f}=f\mod\left(\mathcal{F}_{\text{Bool}}\right)$ with respect to $>$ requires
+    $\mathcal{O}\left(\frac{n^{2}+3\cdot n+2}{2}\cdot\binom{n+\deg\left(f\right)}{\deg\left(f\right)}\right)$
+    field operations.
+  2. (2)
+    Let $\mathcal{F}\subset P$, $D_{\min}=\min_{f\in\mathcal{F}}\deg\left(f\right)$ and $D_{\max}=\max_{f\in\mathcal{F}}\deg\left(f\right)$, then construction of $M_{\leq d}^{\text{Bool}}$ for $\mathcal{F}\cup\mathcal{F}_{\text{Bool}}$ requires
+    $\mathcal{O}\left(\left|\mathcal{F}\right|\cdot\frac{n^{2}+3\cdot n+2}{2}\cdot\sum_{i=0}^{d-D_{\min}}\binom{n}{i}\cdot\binom{n+D_{\max}+i}{D_{\max}+i}\right)$
     field operations.
 Proof.
-There exist $\sum_{i=0}^{\deg\left(f\right)}\binom{n+i-1}{i}$ monomials of degree $\leq\deg\left(f\right)$ and $\binom{n+1}{2}+\binom{n}{1}+1$ monomials of degree $\leq 2$.
+For (1), in $P$ there exist $\sum_{i=0}^{\deg\left(f\right)}\binom{n+i-1}{i}=\binom{n+\deg\left(f\right)}{\deg\left(f\right)}$ monomials of degree $\leq\deg\left(f\right)$ and $\binom{n+1}{2}+\binom{n}{1}+1$ monomials of degree $\leq 2$.
 Now the claim follows from [Ten19, Proposition 3.10].
+For (2), with (1) we can estimate the complexity of the construction as
+$\sum_{f\in\mathcal{F}}\sum_{i=0}^{d-\deg\left(f\right)}\binom{n}{i}\cdot\frac{n^{2}+3\cdot n+2}{2}\cdot\binom{n+\deg\left(f\right)+i}{\deg\left(f\right)+i}$
+$\stackrel{{\scriptstyle(\ast)}}{{\leq}}\frac{n^{2}+3\cdot n+2}{2}\cdot\sum_{f\in\mathcal{F}}\sum_{i=0}^{d-D_{\min}}\binom{n}{i}\cdot\binom{n+D_{\max}+i}{D_{\max}+i}$
+$=\left|\mathcal{F}\right|\cdot\frac{n^{2}+3\cdot n+2}{2}\cdot\sum_{i=0}^{d-D_{\min}}\binom{n}{i}\cdot\binom{n+D_{\max}+i}{D_{\max}+i},$
+where ()($\ast$)( ∗ ) follows from $d-\deg\left(f\right)\leq d-D_{\min}$ and the fact that $\binom{n+d}{d}$ is strictly increasing in $d$, hence $\binom{n+\deg\left(f\right)+i}{\deg\left(f\right)+i}\leq\binom{n+D_{\max}+i}{D_{\max}+i}$.
 ∎
-Thus, the complexity of constructing the $K$-Boolean Macaulay matrix $M_{\leq d}^{\text{Bool}}$ with respect to quadratic $K$-Boolean system $\mathcal{F}_{\text{Bool}}$ is bounded by
-$\mathcal{O}\left(\left|\mathcal{F}\right|\cdot\frac{n^{2}+3\cdot n+2}{2}\cdot\left(\sum_{i=0}^{d}\binom{n+i-1}{i}\right)\cdot\left(\sum_{i=0}^{d}\binom{n}{i}\right)\right).$ (65)
-Moreover, we have the inequality
+Of course, the $K$-Boolean Macaulay matrix generalizes to more advanced linear algebra-based Gröbner basis algorithms like F4/5 [Fau99, Fau02].
+2.5 The Eigenvalue Method
+Now we discuss how the zeros of a polynomial system can be found via linear algebra once a zero-dimensional DRL Gröbner basis of the input system is known [KR16, Chapter 6].
+Let $I\subset P=K[x_{1},\dots,x_{n}]$ be a zero-dimensional ideal, i.e.
+$P/I$ is a finite dimensional $K$-vector space [KR00, Proposition 3.7.1], and let $d=\dim_{K}\left(P/I\right)$.
+If $>$ is a term order on $P$ and $\mathcal{G}\subset I$ is $>$-Gröbner basis, then a $K$-vector space basis $\mathcal{B}$ of $P/I$ is given by the monomials not contained in $\operatorname{LM}_{>}\left(\mathcal{G}\right)=\operatorname{LM}_{>}\left(I\right)$.
+Now let $f\in P$, then the multiplication map
+$\theta_{f}:P/I\to P/I,\qquad\qquad x\mapsto x\cdot f,$ (16)
+is $K$-linear, so it can be represented by a matrix.
+This matrix is called the multiplication matrix $\mathbf{M}_{f}$ for $f$.
+Given a $>$-Gröbner basis its computation is straight-forward: Index the columns of $\mathbf{M}_{f}$ by the $K$-vector space basis $\mathcal{B}$, and the rows by $b\cdot f$, where $b\in\mathcal{B}$.
+The row $b\cdot f$ is then simply the coefficient vector of $b\cdot f$ in $P/I$, which can be computed via $b\cdot f\mod(\mathcal{G})$ with respect to $>$.
+Now assume in addition that $K$ is algebraically closed, then for every point in the variety $\mathbf{x}\in\mathcal{V}\left(I\right)=\left\{\mathbf{x}\in K^{n}\mid\forall g\in I\!:g(\mathbf{x})=0\right\}$ the point $f(\mathbf{x})$ is an eigenvalue of $\mathbf{M}_{f}$, see [KR16, Proposition 6.2.1].
+Therefore, we can compute the variety $\mathcal{V}\left(I\right)$ as follows: Compute the eigenvalues of $\mathbf{M}_{x_{1}},\dots,\mathbf{M}_{x_{n}}$, iterate over all possible combinations, and verify whether a combination is indeed a solution to $I$.
+This approach is known as Eigenvalue Method [KR16, §6.2.A].
+It is well-known that the eigenvalues of a matrix $\mathbf{M}\in K^{N\times N}$ are given by the roots of its characteristic polynomial
+$\chi_{\mathbf{M}}(x)=\det\left(\mathbf{I}_{N\times N}\cdot x-\mathbf{M}\right)\in K[x].$ (17)
+So the complexity of the Eigenvalue Method boils down to $n$ determinant computations with polynomial entries.
+2.5.1 Solving Structured Polynomial Systems Over Finite Fields
+Let =$\mathcal{F}=\{f_{1},\dots,f_{n}\}\subset P=K[x_{1},\dots,x_{n}]$ be such that $\operatorname{LM}_{DRL}\left(f_{i}\right)=x_{i}^{d_{i}}$ for all $1\leq i\leq n$.
+Then, trivially $\mathcal{F}$ is a DRL Gröbner basis, see e.g.
+[CLO15, Chapter 2 §9 Theorem 3, Proposition 4] with $\dim_{K}\left(\mathcal{F}\right)=\prod_{i=1}^{n}d_{i}$.
+Let $D=\prod_{i=1}^{n-1}d_{i}$, and let $\mathcal{B}^{\prime}\subset K[x_{1},\dots,x_{n-1}]/\left(x_{1}^{d_{1}},\dots,x_{n-1}^{d_{n-1}}\right)$ be some $K$-vector space basis, then $\mathcal{B}=\bigcup_{i=0}^{d_{n}-1}x_{n}^{i}\cdot\mathcal{B}^{\prime}$ is a $K$-vector space basis for $P/(\mathcal{F})$.
+In particular, the multiplication matrix $\mathbf{M}_{x_{n}}$ is then of the form
 […]
-on M_≤d^Bool
-[…]
-on M_≤d
-[…]
-on M_≤d^Bool
-[…]
-[…]
-on M_≤d
-[…]
-I.e.
-for $\omega=2$, if we ignore implied constants and prefactors, then we expect that construction of the $K$-Boolean Macaulay matrix dominates the Gröbner basis computation for general quadratic $K$-Boolean polynomial systems.
-Of course, the $K$-Boolean Macaulay matrix generalizes to more advanced linear algebra-based Gröbner basis algorithms.
-Current state-of-the-art are Faugère’s F4 [Fau99] and Matrix-F5 [Fau02] algorithms.
-These algorithms perform Gaussian elimination on the Macaulay matrix $M_{\leq d}$ for increasing values of $d$.
-Moreover, F4 and F5 avoid redundant rows via efficient selection criteria.
-We illustrate a simplified variant of F4.
-Let $\mathcal{F}\subset P=K[x_{1},\dots,x_{n}]$ be a polynomial system, let $>$ be a term order, and let $M_{\leq d}$ be the Macaulay matrix for $\mathcal{F}$ in degree $d$.
-Then, we perform Gaussian elimination on $M_{\leq d}$ which yields the row space basis $\mathcal{G}_{1}$, now we set up a new Macaulay matrix $M_{\leq d}$ for $\mathcal{G}_{1}$ and again compute the row space basis $\mathcal{G}_{2}$ via Gaussian elimination.
-We iterate this procedure until the row space basis becomes stationary, i.e.
-$\mathcal{G}_{i}=\mathcal{G}_{i+1}$, then we verify whether $\mathcal{G}_{i}$ is already a $>$-Gröbner basis.
-If not then we pass to degree $d+1$ and repeat the procedure.
-Obviously, if the input system contains a $K$-Boolean polynomial system, then we can replace every full Macaulay matrix by its $K$-Boolean submatrix.
-7 Cryptanalysis of Ciminion & Hydra
-In this section we discuss how our results from the previous sections can be applied to cryptanalysis of Ciminion and Hydra.
-7.1 Ciminion
+where $\mathbf{0}\in K^{D\times D}$ denotes the zero matrix, $\mathbf{I}\in K^{D\times D}$ the identity matrix, and $\mathbf{A}_{0},\dots,\allowbreak\mathbf{A}_{d_{n}-1}\in K^{D\times D}$ some matrices.
+By [BBL+24, Lemma 2] we then have that
+$\chi_{\mathbf{M}_{x_{n}}}(x)=\det\left(x_{n}^{d_{n}}\cdot\mathbf{I}-\sum_{i=0}^{d_{n}-1}x_{n}^{i}\cdot\mathbf{A}_{i}\right),$ (19)
+and computation of this determinant requires [BBL+24, §3.2]
+$\mathcal{O}\bigg{(}d_{n}\cdot\log_{2}(d_{n})^{2}\cdot\Big{(}1+\log_{2}\big{(}\log_{2}(d_{n})\big{)}\Big{)}\cdot\Big{(}\frac{D}{d_{n}}\Big{)}^{\omega}\bigg{)}$ (20)
+field operations, where $2.3715522\leq\omega<$ 2.371552 is again the linear algebra constant [WXXZ24].
+In this paper we solely work over finite fields $\mathbb{F}_{q}$, and we are only interested in the $\mathbb{F}_{q}$-valued roots of $\mathcal{F}$.
+Given a univariate polynomial $f\in\mathbb{F}_{q}$ with $\deg\left(f\right)=d$ we can always achieve this restriction by passing to the GCD with the field equation $g=\gcd\left(f,x^{q}-x\right)$.
+By [BBLP22, §3.1] computation of this GCD requires
+$\mathcal{O}\left(\begin{rcases}\begin{dcases}d\cdot\log_{2}(d)\cdot\log\big{(}\log(d)\big{)}\cdot\big{(}\log(d)+\log_{2}(q)\big{)},&d<q,\\ d\cdot\log(d)^{2}\cdot\log\big{(}\log(d)\big{)},&d\geq q\end{dcases}\end{rcases}\right)$ (21)
+field operations. Provided that only a few roots of $f$ come from $\mathbb{F}_{q}$ the final factoring step can be neglected.
+In this paper we will encounter quadratic DRL Gröbner bases =$\mathcal{F}=\{f_{1},\dots,f_{n}\}\subset P=\mathbb{F}_{q}[x_{1},\dots,x_{n}]$ of the form
+$f_{i}=\mathcal{Q}\left(x_{i},\dots,x_{n}\right)+\mathcal{A}_{i}\left(x_{1},\dots,x_{n}\right),$ (22)
+where $\mathcal{Q}_{i}\in P$ is a quadratic polynomial which has the term $x_{i}^{2}$ and $\mathcal{A}_{i}\in P$ is an affine polynomial which has the term $x_{n-i+1}$ for $1\leq i\leq n$.
+It is easy to see that $\mathcal{F}$ is a DRL Gröbner basis for $x_{1}>\ldots>x_{n}$.
+For such systems we can give an optimized Eigenvalue Method to find the $\mathbb{F}_{q}$-valued solutions.
+We use a $\mathbb{F}_{q}$-vector space basis such that the multiplication matrix $\mathbf{M}_{x_{n}}$ has the shape from Equation 18.
+Then, the characteristic polynomial is constructed in $\mathcal{O}\left(2^{\omega\cdot(n-1)+1}\right)$ field operations, and its $\mathbb{F}_{q}$-valued roots are extracted via the GCD with the field equation.
+Let $\alpha\in\mathbb{F}_{q}$ be such a solution for $x_{n}$, then we can substitute $x_{n}=\alpha$ back into $\mathcal{F}$.
+In particular, from $f_{n}$ we obtain the affine equation
+$\alpha^{2}+\tilde{\mathcal{A}_{n}}=0.$ (23)
+Now we rearrange this equation to have $x_{1}$ on the right-hand side.
+Then, we can additionally eliminate the variable $x_{1}$ in the equations $f_{1}=\ldots=f_{n-1}=0$.
+Since $x_{1}$ is not present in the quadratic terms of $f_{2},\dots,f_{n-1}\in\mathbb{F}_{q}[x_{2},\dots,x_{n-1}]$, they are again a zero-dimensional DRL Gröbner basis.
+The polynomial $f_{1}$ is an additional overdetermined polynomial which we ignore.
+Now we iterate this procedure to solve for $x_{n-1},x_{n-2},\dots,x_{\left\lfloor{\frac{n}{2}}\right\rfloor}$, then we lift to a global solution guess for $x_{1},\dots,x_{n}$, and finally verify it on $\mathcal{F}$.
+Let $\mathcal{D}(d,q)$ denote the complexity of the GCD between a degree $d$ polynomial and the field equation of $\mathbb{F}_{q}$, moreover let $N$ be a bound on the number of $\mathbb{F}_{q}$-valued solutions at each iteration in our solving algorithm.
+Then, this dedicated Eigenvalue Method requires
+$\mathcal{O}\left(\begin{rcases}\begin{dcases}\sum_{i=1}^{\frac{n}{2}}N^{\frac{n}{2}-i}\cdot\Big{(}2^{\omega\cdot(2\cdot i-1)+1}+\mathcal{D}\big{(}2^{2\cdot i},q\big{)}\Big{)},&2\mid n,\\ \sum_{i=1}^{\frac{n-1}{2}}N^{\frac{n-1}{2}-i}\cdot\Big{(}2^{\omega\cdot 2\cdot i+1}+\mathcal{D}\big{(}2^{2\cdot i+1},q\big{)}\Big{)},&2\nmid n\end{dcases}\end{rcases}\right)$ (24)
+$=\mathcal{O}\left(\begin{rcases}\begin{dcases}2^{\omega+1}\cdot\frac{2^{\omega\cdot n}-N^{\frac{n}{2}}}{2^{2\cdot\omega}-N}+\sum_{i=1}^{\frac{n}{2}}N^{\frac{n}{2}-i}\cdot\mathcal{D}\big{(}2^{2\cdot i},q\big{)},&2\mid n,\\ 2^{2\cdot\omega+1}\cdot\frac{2^{\omega\cdot(n-1)}-N^{\frac{n-1}{2}}}{2^{2\cdot\omega}-N}+\sum_{i=1}^{\frac{n-1}{2}}N^{\frac{n-1}{2}-i}\cdot\mathcal{D}\big{(}2^{2\cdot i+1},q\big{)},&2\nmid n\end{dcases}\end{rcases}\right)$ (25)
+field operations.
+3 Analysis of Ciminion
+At round level Ciminion is a three branch Feistel network.
+In [Ste24, §4.1] a DRL Gröbner basis for Feistel-$2n/n$ polynomial systems has been computed by substituting the linear branches into the non-linear ones.
+In principle, we are going to replicate this strategy to construct a DRL Gröbner basis for Ciminion.
+3.1 A DRL Gröbner Basis for Ciminion
+For simplicity let us denote the matrix in the $i$th Ciminion round by $\mathbf{A}_{i}\in\mathbb{F}_{q}^{3\times 3}$, see Equation 2, then
+$\mathbf{A}_{i}=\begin{pmatrix}0&0&1\\ 1&c_{4}^{(i)}&c_{4}^{(i)}\\ 0&1&1\end{pmatrix},\qquad\mathbf{A}_{i}^{-1}=\begin{pmatrix}0&1&-c_{4}^{(i)}\\ -1&0&1\\ 1&0&0\end{pmatrix},$ (26)
+where $c_{4}^{(i)}\in\mathbb{F}_{q}\setminus\{0,1\}$.
+Theorem 3.1.
+Let $\mathbb{F}_{q}$ be a finite field, and let $\mathcal{F}_{\texttt{Ciminion}}=\left\{\mathbf{f}^{(1)},\dots,\mathbf{f}^{(r)}\right\}\subset\mathbb{F}_{q}\Big{[}y_{1},y_{2},\mathbf{x}^{(1)},\allowbreak\dots,\mathbf{x}^{(r-1)},x\Big{]}$ be a Ciminion polynomial system.
+For the DRL term order $y_{1}>y_{2}>\mathbf{x}^{(1)}>\ldots>\mathbf{x}^{(r-1)}>x$, the following procedure computes a DRL Gröbner basis for Ciminion:
+  1. (1)
+    For every round $1\leq i\leq r$ compute $\mathbf{g}^{(i)}=\mathbf{A}_{i}^{-1}\mathbf{f}_{i}$.
+  2. (2)
+    Collect the linear polynomials from the $\mathbf{g}^{(i)}$’s in $\mathcal{L}$, and perform Gaussian elimination on $\mathcal{L}$.
+  3. (3)
+    Reduce the non-linear polynomials in the $\mathbf{g}^{(i)}$’s modulo $\mathcal{L}$.
+Proof.
+By inspection of the round function (Equation 2) we have that
+$\mathbf{g}^{(1)}$ =($\ y_{1}\\ \aleph\cdot y_{1}+y_{2}\end{pmatrix}+\mathbf{A}_{1}^{-1}\left(\mathbf{c}^{(1)}-\mathbf{x}^{(1)}\right),$ (27)
+$\mathbf{g}^{(i)}$ $=\begin{pmatrix}x_{1}^{(i-1)}\\ x_{2}^{(i-1)}\\ x_{3}^{(i-1)}+x_{1}^{(i-1)}\cdot x_{2}^{(i-1)}\end{pmatrix}+\mathbf{A}_{i}^{-1}\left(\mathbf{c}^{(i)}-\mathbf{x}^{(i)}\right),$ (28)
+$\mathbf{g}^{(r)}$ $=\begin{pmatrix}x_{1}^{(r-1)}\\ x_{2}^{(r-1)}\\ x_{3}^{(r-1)}+x_{1}^{(r-1)}\cdot x_{2}^{(r-1)}\end{pmatrix}+\mathbf{A}_{r}^{-1}\left(\mathbf{c}^{(r)}-\begin{pmatrix}c_{1}-p_{1}\\ c_{2}-p_{2}\\ x\end{pmatrix}\right).$ (29)
+Hence, $\mathcal{L}$ contains $\mathbf{g}^{(1)}$ and the first two components of all remaining $\mathbf{g}^{(i)}$’s.
+For Gaussian elimination on $\mathcal{L}$ let us start in the last round $i=r$.
+By Equations 26 and 29 the linear polynomials with leading monomials $x_{1}^{(r-1)}$, $x_{2}^{(r-1)}$ contain a linear term in $x$ and a constant term.
+Moving to $i=r-1$ next the linear polynomials from $\mathbf{g}^{(r-1)}$ depend on $\mathbf{x}^{(r-2)}$ and $\mathbf{x}^{(r-1)}$, and via Gaussian elimination the variables $x_{1}^{(r-1)}$ and $x_{2}^{(r-1)}$ get eliminated.
+Thus, the linear polynomials with leading monomials in $x_{1}^{(r-2)}$ and $x_{2}^{(r-2)}$ can only depend on $x_{3}^{(r-1)}$ and $x$.
+Moreover, by Equation 26 the coefficients for $x_{3}^{(r-1)}$ are non-zero.
+Via a downwards induction we can then conclude for $12\leq i\leq r-$ 1 that after Gaussian elimination
+$\operatorname{LM}_{DRL}\left(g_{1}^{(i)}\right)$ $=x_{1}^{(i-1)},$
+$\operatorname{LM}_{DRL}\left(g_{2}^{(i)}\right)$ $=x_{2}^{(i-1)},$
+$g_{1}^{(i)},g_{2}^{(i)}$ $\in\mathbb{F}_{q}\big{[}x_{3}^{(i)},\dots,x_{3}^{(r-1)},x\big{]},$
+and that the coefficients of $x_{3}^{(i)}$ in $g_{1}^{(i)}$ and $g_{2}^{(i)}$ are non-zero.
+The first round needs special care since the first row of $\mathbf{A}_{1}^{-1}$ is non-zero on the variable $x_{2}^{(1)}$.
+(If it were zero on $x_{1}^{(1)}$ and $x_{2}^{(1)}$, the first polynomial in $\mathbf{g}^{(1)}$ would have leading monomial $x_{3}^{(1)}$, and we could proceed.) Expansion of Equation 27 yields
+$\mathbf{g}^{(1)}=\begin{pmatrix}\aleph\\ y_{1}\\ \aleph\cdot y_{1}+y_{2}\end{pmatrix}+\mathbf{A}_{1}^{-1}\mathbf{c}^{(1)}-\begin{pmatrix}x_{2}^{(1)}-c_{4}^{(1)}\cdot x_{3}^{(1)}\\ -x_{1}^{(1)}+x_{3}^{(1)}\\ x_{1}^{(1)}\end{pmatrix}.$
+Producing a polynomial with leading monomial $y_{2}$ is straight-forward, we simply have to take $g_{3}^{(1)}-\aleph\cdot g_{2}^{(1)}$.
+On the other hand, there is second polynomial with leading monomial $x_{2}^{(1)}$ after inversion of the matrices
+$g_{2}^{(2)}=x_{2}^{(1)}+\left(\mathbf{A}_{2}^{-1}\mathbf{c}^{(2)}\right)_{2}-\left(-x_{1}^{(2)}+x_{3}^{(2)}\right).$
+Then
+$g_{1}^{(1)}+g_{2}^{(2)}=\aleph+\left(\mathbf{A}_{1}^{-1}\mathbf{c}^{(1)}\right)_{1}+c_{4}^{(1)}\cdot x_{3}^{(1)}+\left(\mathbf{A}_{2}^{-1}\mathbf{c}^{(2)}\right)_{2}+x_{1}^{(2)}-x_{3}^{(2)},$
+and by the choice of our DRL term order and $c_{4}^{(1)}\neq 0$ this polynomial has leading monomial $x_{3}^{(1)}$.
+Obviously, after Gaussian elimination of $\mathcal{L}$ the terms of $g_{1}^{(1)}+g_{2}^{(2)},g_{2}^{(1)},g_{3}^{(1)}-\aleph\cdot g_{2}^{(1)}$ after the leading monomials will also only depend on the variables $x_{3}^{(2)},\dots,x_{3}^{(r-1)},x$.
+Lastly, let us consider the reduction modulo $\mathcal{L}$. By our previous observations all linear equations are of the form
+$y_{1}+\mathcal{A}_{1}\left(x_{3}^{(2)},\dots,x_{3}^{(r-1)},x\right)$ $=0,$
+$y_{2}+\mathcal{A}_{2}\left(x_{3}^{(2)},\dots,x_{3}^{(r-1)},x\right)$ $=0,$
+$x_{3}^{(1)}+\mathcal{A}_{3}\left(x_{3}^{(2)},\dots,x_{3}^{(r-1)},x\right)$ $=0,$
+$x_{1}^{(i-1)}+c_{4}^{(i)}\cdot x_{3}^{(i)}+\ldots$ $=0,$
+$x_{2}^{(i-1)}-x_{3}^{(i)}+\ldots$ $=0,$
+$x_{1}^{(r-1)}+c_{4}^{(r)}\cdot x+\ldots$ $=0,$
+$x_{2}^{(r-1)}-x+\ldots$ $=0,$
+where $\mathcal{A}_{1}$, $\mathcal{A}_{2}$ and $\mathcal{A}_{3}$ are affine polynomials and $12\leq i\leq r-$ 1. So, for $2\leq i\leq r$ we can directly conclude from Equations 28 and 29 that
+$\hat{g}_{3}^{(i)}=\left(g_{3}^{(i)}\mod(\mathcal{L})\right)\in\mathbb{F}_{q}\Big{[}x_{3}^{(2)},\dots,x_{3}^{(r-1)},x\Big{]},$
+has leading monomial ${x_{3}^{(i)}}^{2}$ and respectively $x^{2}$ for the last round.
+Thus, we have constructed a polynomial system with pairwise coprime leading monomials, so by [CLO15, Chapter 2 §9 Theorem 3, Proposition 4] we have constructed a Gröbner basis. ∎
+Note that if we only keep the $r-1$ quadratic polynomials of $\mathcal{G}$, then we still have a fully determined polynomial system in $r-1$ variables.
+Therefore, we call the quadratic polynomials of $\mathcal{G}$ the downsized Ciminion polynomial system.
+For the next corollary, recall that a LEX Gröbner basis =$\mathcal{F}=\{f_{1},\dots,f_{n}\}\subset K[x_{1},\dots,\allowbreak x_{n}]$ is said to be in $x_{n}$-shape position if it is of the form
+$f_{i}$ $=x_{i}-\tilde{f}_{n}(x_{n}),\qquad 1\leq i\leq n-1,$ (30)
+$f_{n}$ $=\tilde{f}_{n}(x_{n})$ (31)
+It is well-known that the LEX Gröbner basis of a radical ideal can be brought into shape position for some ordering of the variables, see [KR00, Theore 3.7.25].
+Corollary 3.2.
+In the situation of Theorem 3.1, let $\mathcal{G}_{\texttt{Ciminion}}$ denote the Ciminion DRL Gröbner basis.
+  1. (1)
+    $\dim_{\mathbb{F}_{q}}\left(\mathcal{F}_{\texttt{Ciminion}}\right)=2^{r-1}=2^{r_{C}+r_{E}-1}$.
+  2. (2)
+    Let $I\subset\overline{\mathbb{F}_{q}}\Big{[}y_{1},y_{2},\mathbf{x}^{(1)},\dots,\mathbf{x}^{(r-1)},x\Big{]}[x_{0}]$ be a homogeneous ideal such that $\mathcal{Z}_{+}(I)\neq\emptyset$ and $\mathcal{G}_{\texttt{Ciminion}}^{\text{hom}}\subset I$.
+Then $I$ is in generic coordinates.
+  3. (3)
+    $\operatorname{sd}_{DRL}\left(\mathcal{G}_{\texttt{Ciminion}}\right)\leq r=r_{C}+r_{E}$.
+  4. (4)
+    The LEX Gröbner basis for Ciminion under $y_{1}>y_{2}>\mathbf{x}^{(1)}>\ldots>\mathbf{x}^{(r-1)}>x$ is in $x$-shape position.
+Proof.
+(1), (2) and (3) are consequences of the DRL Gröbner basis from Theorem 3.1 and application of [Ste24, Theorem 3.2] and Theorem 2.6.
+For the last claim we take a look at the shape of the quadratic polynomials in the DRL Gröbner basis
+$g_{3}^{(2)}$ $=\mathcal{Q}_{2}\left(x_{3}^{(2)},\dots,x_{3}^{(r-1)},x\right)+\mathcal{A}_{2}\left(x_{3}^{(2)},\dots,x_{3}^{(r-1)},x\right),$
+$g_{3}^{(i)}$ $=\mathcal{Q}_{i}\left(x_{3}^{(i)},\dots,x_{3}^{(r-1)},x\right)+\mathcal{A}_{i}\left(x_{3}^{(i-1)},\dots,x_{3}^{(r-1)},x\right),\qquad 3\leq i\leq r$
+$g_{3}^{(r)}$ $=\mathcal{Q}\left(x\right)+\mathcal{A}_{r}\left(x_{3}^{(r-1)},x\right),$
+where the $\mathcal{Q}_{i}$’s are quadratic polynomials which contain the monomial ${x_{3}^{(i)}}^{2}$ and respectively $x^{2}$ for the last round, and the $\mathcal{A}_{i}$’s are affine polynomials that contain the monomial $x_{3}^{(i-1)}$ for all $i\geq 3$ and without any restrictions on $\mathcal{A}_{2}$.
+Then, by iterative substitution of $g_{3}^{(r)}$ into $g_{3}^{(r-1)}$ to eliminate $x_{3}^{(r-1)}$, $g_{3}^{(r-1)}$ into $g_{3}^{(r-2)}$ to eliminate $x_{3}^{(r-2)}$, etc.
+we produce a LEX Gröbner basis in $x$-shape position.
+∎
+3.2 Bariant’s Attack
+Recently Bariant introduced an attack on Ciminion via univariate greatest common divisor computations [Bar23].
+As in our polynomial model, he introduced an auxiliary variable $X$ for the truncated component.
+If one inverts $p_{C}\circ p_{E}$ using $(c_{1}-p_{1},c_{2}-p_{2},X)^{\intercal}$ as output state, then one obtains three polynomials, see [Bar23, Fig.
+4],
+$\aleph=f_{1}(X),\qquad K_{1}=f_{2}(X),\qquad K_{2}=f_{3}(X).$ (32)
+Since the first polynomial is univariate in $X$, one can compute its greatest common divisor with $X^{q}-X$ to solve for $X$ and henceforth also solve for $K_{1}$ and $K_{2}$.
+Observe that Bariant’s univariate polynomial is exactly the univariate polynomial of the LEX Gröbner basis.
+Lemma 3.3.
+Let $\mathbb{F}_{q}$ be a finite field, let $I_{\texttt{Ciminion}}\subset\mathbb{F}_{q}\big{[}y_{1},y_{2},\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(r-1)},x\big{]}$ be a Ciminion ideal, let $\tilde{f}\in\mathbb{F}_{q}[x]$ be the univariate LEX polynomial from Corollary 3.2, and let $f\in\mathbb{F}_{q}[x]$ be Bariant’s univariate polynomial.
+Then $f=\alpha\cdot\tilde{f}$ for some $\alpha\in\mathbb{F}_{q}^{\times}$.
+Proof.
+Ciminion is an iterated Feistel permutation therefore we have that
+$\mathcal{R}^{(i)}\left(\mathbf{x}^{(i-1)}\right)\equiv\mathbf{x}^{(i)}\mod I_{\texttt{Ciminion}}\qquad\Longleftrightarrow\qquad\mathbf{x}^{(i-1)}\equiv{\mathcal{R}^{(i)}}^{-1}\left(\mathbf{x}^{(i)}\right)\mod I_{\texttt{Ciminion}},$
+where $12\leq i\leq r-$ 1, and analogously for $i=1,r$.
+Since the inverse round functions are in $I_{\texttt{Ciminion}}$, we can iteratively substitute them into each other.
+This yields exactly the polynomials from Equation 32, so $f\in I_{\texttt{Ciminion}}$.
+By Corollary 3.2 we can construct a Ciminion LEX Gröbner basis in $x$-shape position, then for any $g\in\mathbb{F}_{q}[x]$ we have $g\in I_{\texttt{Ciminion}}\Leftrightarrow\tilde{f}\mid g$ [CLO15, Chapter 4 §5 Exercise 13].
+Since $\deg\left(\tilde{f}\right)=\deg\left(f\right)$ they can only differ up to a multiplicative constant.
+∎
+Moreover, the construction of the univariate LEX polynomial can be effectively expressed in the cost of univariate polynomial multiplication over finite fields [CK91], which is quasi-linear in the degree.
+As a countermeasure, Bariant proposes to add a linear combination of $K_{1}$ and $K_{2}$ to each component after $p_{C}$ [Bar23, §4].
+We note that in general this modification breaks our Gröbner basis from Theorem 3.1.
+As a settlement, we propose to add a linear combination of $K_{1}$ and $K_{2}$ only to the non-linear parts of the round functions, i.e.
+$\begin{split}\mathcal{R}^{(i)}:\mathbb{F}_{q}^{3}\times\mathbb{F}_{q}^{2}&\to\mathbb{F}_{q}^{3},\\ \begin{pmatrix}x\\ y\\ z\end{pmatrix},\begin{pmatrix}K_{1}\\ K_{2}\end{pmatrix}&\mapsto\begin{pmatrix}0&0&1\\ 1&c_{4}^{(i)}&c_{4}^{(i)}\\ 0&0&1\end{pmatrix}\begin{pmatrix}x\\ y\\ z+x\cdot y+\alpha\cdot K_{1}+\beta\cdot K_{2}\end{pmatrix}+\mathbf{c}^{(i)},\end{split}$ (33)
+where $\alpha,\beta\in\mathbb{F}_{q}^{\times}$.
+This modification does not affect the proof of Theorem 3.1, hence we maintain the DRL Gröbner basis for the modified Ciminion polynomial system.
+However, this modification breaks the construction of the univariate LEX polynomial via univariate polynomial multiplication, because the quadratic DRL Gröbner basis polynomial for the $r_{C}$th round will now also contain linear terms for the variables $x_{3}^{(2)},\dots,x_{3}^{(r-1)},x$ instead of only $x_{3}^{(r_{C}-1)},\dots,x_{3}^{(r-1)},x$.
+While the LEX Gröbner basis of the modified Ciminion might still be in $x$-shape position, we can now construct it only via an FGLM algorithm.
+Note if the modification of Equation 33 is performed in all rounds of $p_{E}$, then the Middle State-Output Relation polynomial system of Bariant et al.
+[BBLP22, §5] becomes underdetermined, so this attack also becomes more difficult.
+3.3 Implementation
+We implemented Ciminion in the computer algebra systems SageMath [Sag24] and OSCAR [OSC24].
+[^5] https://github.com/sca-research/Groebner-Basis-Cryptanalysis-of-Ciminion-and-Hydra.git The SageMath implementation can handle $128$ bit prime numbers, in particular the Gröbner basis from Theorem 3.1 can be computed for primes of that size.
+For $p=2^{127}+45$, $r_{C}=90$ and $r_{E}=14$ it is computed in negligible time.
+3.3.1 Ciminion2
+We also implemented a modified version of Ciminion to protect against Bariant’s attack, called Ciminion2, which adds $K_{1}+K_{2}$ in every non-linear round of $p_{E}$ and $p_{C}$, i.e.
+only in the first round of $p_{E}$ the key addition is omitted.
+The round function for the first sample then modifies to
+$\mathcal{R}^{(i)}:\begin{pmatrix}x\\ y\\ z\end{pmatrix},\begin{pmatrix}K_{1}\\ K_{2}\end{pmatrix}\mapsto\begin{pmatrix}0&0&1\\ 1&c_{4}^{(i)}&c_{4}^{(i)}\\ 0&1&1\end{pmatrix}\begin{pmatrix}x\\ y\\ z+x\cdot y+K_{1}+K_{2}\end{pmatrix}+\mathbf{c}^{(i)},$ (34)
+where $2\leq i\leq r_{C}+r_{E}$.
+For higher samples one then adds $K_{2\cdot j-1}+K_{2\cdot j}$ in $p_{E}$.
+For Ciminion2 a DRL Gröbner basis for the first sample is also computed in negligible time in SageMath for practical parameters.
+Due to the additional key addition the quadratic polynomials in the DRL Gröbner basis modify to
+$g_{3}^{(i)}$ $=\mathcal{Q}_{i}\left(x_{3}^{(i)},\dots,x_{3}^{(r-1)},x\right)+\mathcal{A}_{i}\left(x_{3}^{(2)},\dots,x_{3}^{(r-1)},x\right),\qquad 2\leq i\leq r-1,$ (35)
+$g_{3}^{(r)}$ $=\mathcal{Q}_{r}(x)+\mathcal{A}_{i}\left(x_{3}^{(2)},\dots,x_{3}^{(r-1)},x\right),$ (36)
+where $\mathcal{Q}_{i}$ is quadratic and has the term ${x_{3}^{(i)}}^{2}$ and respectively $x^{2}$ for the last round, and $\mathcal{A}_{i}$ is affine and has non-zero coefficients for $x_{3}^{(2)},\dots,x_{3}^{(r-1)},x$.
+3.4 Cryptanalysis
+We now analyze the complexity of a Gröbner basis attack on Ciminion and Ciminion2 via our Gröbner basis, and compare it to the designer’s estimation and Bariant’s attack.
+3.4.1 Estimation of the Designers
 For Gröbner basis cryptanalysis of Ciminion [DGGK21, §4.4] the designers considered a weaker scheme where the keys are added after the permutations $p_{E}$ instead of before the rolling functions, see [DGGK21, Fig.
 7].
-The designer claim that the weaker scheme is simpler to analyze than Ciminion.
-We reject this claim for the iterated Ciminion polynomial system, after all we have computed a DRL Gröbner basis for Ciminion with rather simple affine variable eliminations, see LABEL:Th:_Ciminion_Gr\"obner_basis.
-The designers assumed that weakened Ciminion polynomial systems are regular, moreover for regular systems it is well-known that the degree of regularity is equal to the Macaulay bound.
+The designers claimed that the weaker scheme is simpler to analyze than Ciminion.
+We reject this claim, after all we have computed a DRL Gröbner basis for Ciminion with rather simple linear variable eliminations, see Theorem 3.1.
+The designers assumed that weakened Ciminion polynomial systems are regular, moreover for regular systems it is well-known that the degree of regularity $d_{\operatorname{reg}}$ is equal to the Macaulay bound (Theorem 2.6).
 Then, the DRL Gröbner basis complexity is bounded by [BFS04, Theorem 7]
-$\mathcal{O}\left(m\cdot d_{\operatorname{reg}}\cdot\binom{n+d_{\operatorname{reg}}}{d_{\operatorname{reg}}}^{\omega}\right).$ (67)
-For Ciminion cryptanalysis, it is then assumed that the DRL Gröbner basis computation is the most costly part in an attack.
-The Ciminion designers did not provide a dedicated analysis for the iterated polynomial model from Definition 2.1.
-They mostly analyze polynomial models where $p_{E}$ and $p_{C}$ or respectively $p_{E}\circ p_{C}$ are modeled without intermediate variables, i.e.
-they used a single polynomial vector for these permutations.
-The designers further claim that polynomial models with many intermediate variables do not improve upon the few variables but high degree polynomial models.
-On the other hand, let $I,J\subset P$ be zero-dimensional ideals, then elementary we have that
-$I\subset J\qquad\Rightarrow\qquad\mathcal{Z}\left(J\right)\subset\mathcal{Z}\left(I\right),$ (68)
-and by the isomorphism theorem for rings
-$I\subset J\qquad\Rightarrow\qquad\big{(}P/I\big{)}\big{/}(J/I)\cong P/J.$ (69)
-In particular, the latter isomorphism also induces a surjection of finite dimensional $K$-vector spaces $P/I\twoheadrightarrow P/J$ which implies that $\dim_{K}\left(I\right)\geq\dim_{K}\left(J\right)$.
-In cryptographic terms this means that if $>$-Gröbner bases for $I$ and $J$ are known, then solving for the zero locus of $I$ is at least as hard as solving for the zero locus of $J$ if one uses only generic algorithms.
-Back to Ciminion, the designers proposed to study the model in two variables for the keys for the full function, we call this model the fully substituted model. Trivially, we have that
-(fully substituted Ciminion model)(iterated Ciminion $(\text{fully substituted {Ciminion} model})\subset(\text{iterated {Ciminion} model}).$ (70)
-Since we have a DRL Gröbner basis for the iterated model at hand, we reject the claim by the Ciminion designers that introducing intermediate state variables does not improve upon the fully substituted model.
-So, the only non-trivial step in a fully determined Ciminion polynomial system is the construction of a univariate polynomial.
-As outlined in Corollary 3.2 and Section 3.2, the univariate LEX polynomial can be constructed by iterated substitution of a univariate polynomial into a quadratic polynomial.
-By [HvdH22] the complexity of univariate polynomial multiplication over a finite field is
-$\mathcal{O}\Big{(}d\cdot\log(q)\cdot\log\big{(}d\cdot\log(q)\big{)}\Big{)},$ (71)
-where $d$ is the maximum of the degrees of the polynomials. Then, the complexity of constructing the univariate LEX polynomial of Ciminion is bounded by
-$\mathcal{O}\Big{(}2^{r_{C}+r_{E}-2}\cdot\log(q)\cdot\log\big{(}2^{r_{C}+r_{E}-2}\cdot\log(q)\big{)}\Big{)}.$ (72)
-In Table 3 we present the complexity of term order conversion via fast polynomial multiplication and the GCD with the field equation (Equation 1) for Ciminion.
-Since the complexities of polynomial multiplication and the GCD are quasi-linear in the degree, one requires more than $110$ rounds for standard Ciminion to achieve $128$ bits of security.
-Recall that we proposed a slightly modified variant Ciminion2 (Section 3.3.1) with a key addition to the non-linear component in the last rounds of $p_{C}$ every round of $p_{E}$.
-As discussed in Section 3.2, this modification breaks the construction of the univariate LEX polynomial via fast polynomial multiplication.
-For Ciminion2 we therefore either have to do term order conversion to LEX or use the Eigenvalue method.
-For term order conversion via the probabilistic FGLM algorithm [FGHR14], since we cannot prove $x$-shape position we have to compute the LEX Gröbner basis for the radical ideal which has complexity
-$\tilde{\mathcal{O}}\left((r_{C}+r_{E}-1)\cdot 2^{\omega\cdot(r_{C}+r_{E}-1)}\right).$
-On the other hand, for Ciminion2 it suffices to solve for the key variables via the Eigenvalue method.
-Since the Ciminion2 DRL Gröbner basis satisfies the assumption of Section 5, the complexity of characteristic polynomial construction is
-$\mathcal{O}\left(2\cdot 2^{\omega\cdot(r_{C}+r_{E}-2)}\right).$ (73)
-In Table 3 we present the complexity of the Eigenvalue Method and the GCD with the field equation for Ciminion2.
-Since the Eigenvalue Method adds a factor $\omega$ (in logarithmic terms), more than $66$ rounds for Ciminion2 already achieve $128$ bits of security.
-If we add additional polynomials to the Ciminion or Ciminion2 polynomial systems, then we can recompute the DRL Gröbner bases.
-It is clear from the Macaulay bound (Theorem 2.5) that a polynomial $f\in P$ with $\deg\left(f\right)\geq 2$ can only increase the Macaulay bound.
-On the other hand, if $\deg\left(f\right)=1$, then we can eliminate one variable for free.
-Note that this does not affect the property of being in generic coordinates, but we would have to recompute the Ciminion Macaulay bound (Corollary 3.2).
-For the former case, we use
-$\mathcal{O}\left(\binom{2\cdot(r_{C}+r_{E}-1)}{r_{C}+r_{E}}^{\omega}\right)$ (74)
-as baseline for the complexity of DRL Gröbner basis computations.
-Based on our proven DRL complexity estimate, more than $36$ rounds for Ciminion and Ciminion2 polynomial systems achieve $128$ bits of security.
-Lastly, in Table 3 we provide an additional column to compare our estimations to the rational of the Ciminion designers.
-In [DGGK21, §4.4] it is proposed to model the Ciminion function via two polynomials in two variables, i.e.
-the fully substituted model, of degree $2^{r_{C}+r_{E}-1}$.
-Under the assumption that these polynomials are regular we can then use the binomial term in Equation 67 as baseline for DRL Gröbner basis computations
-$\mathcal{O}\left(\binom{2+2\cdot\left(2^{r_{C}+r_{E}-1}-1\right)+1-1}{2\cdot\left(2^{r_{C}+r_{E}-1}-1\right)+1}^{\omega}\right)=\mathcal{O}\left(2^{\omega\cdot(r_{C}+r_{E})}\right),$ (75)
-For completeness, this complexity should not be compared to Equation 74.
-Since our DRL Gröbner basis complexity assumes that we have a Ciminion DRL Gröbner basis at hand plus some additional information that can be modeled via a non-linear polynomial.
-The complexity of Equation 75 quantifies the complexity of finding a DRL Gröbner basis in the first place without any additional information.
-Now ignorance of the iterated polynomial model in [DGGK21] becomes even more damaging.
-After all, we have constructed an iterated DRL Gröbner basis in $\mathcal{O}\left(r_{C}+r_{E}\right)$ while for the fully substituted model we obtain an exponential complexity estimate for finding the DRL basis.
-Moreover, the strategies for direct polynomial system solving outperform the complexity of finding the DRL basis for the fully substituted model.
+$\mathcal{O}\left(\binom{n+d_{\operatorname{reg}}}{d_{\operatorname{reg}}}^{\omega}\right)$ (37)
+field operations. For Ciminion cryptanalysis, it is then conjectured that the DRL Gröbner basis computation is the most costly part in an attack.
+The designers did not analyze the iterated model, they only studied a model of two equations in two key variables, we call this model the fully substituted model.
+By construction the polynomials in the fully substituted model have degree $2^{r_{C}+r_{E}-1}$.
+Under the regularity assumption we then have that $d_{\operatorname{reg}}=2\cdot\left(2^{r_{C}+r_{E}-1}-1\right)+1=2^{r_{C}+r_{E}}-1$, so Equation 37 evaluates to
+$\mathcal{O}\left(\binom{2^{r_{C}+r_{E}}+1}{2^{r_{C}+r_{E}}-1}^{\omega}\right)=\mathcal{O}\left(\left(\frac{2^{r_{C}+r_{E}}\cdot\left(2^{r_{C}+r_{E}}+1\right)}{2}\right)^{\omega}\right).$ (38)
+In Table 3 this estimate is evaluated for sample we round numbers.
+3.4.2 Complexity of Bariant’s Attack
+Bariant’s attack is split into two parts: construction of the univariate polynomial via an iterated sequence of squarings and the GCD with the field equation.
+Multiplication of two univariate polynomials $f,g\in\mathbb{F}_{q}[x]$ with $\deg\left(f\right),\deg\left(g\right)\leq d$ can be performed in $\mathcal{O}\Big{(}d\cdot\log_{2}(d)\cdot\log_{2}\big{(}\log_{2}(d)\big{)}\Big{)}$ field operations [CK91], and the complexity for the GCD was given in Equation 21.
+The construction step then requires [Bar23, §3.2]
+$\mathcal{O}\big{(}2^{r_{C}+r_{E}}\cdot(r_{C}+r_{E})\cdot\log_{2}(r_{C}+r_{E})\big{)}$ (39)
+field operations, and the GCD requires
+$\mathcal{O}\Big{(}2^{r_{C}+r_{E}-1}\cdot(r_{C}+r_{E}-1)\cdot\big{(}r_{C}+r_{E}-1+\log_{2}(q)\big{)}\cdot\log_{2}(r_{C}+r_{E})\Big{)}.$ (40)
+Therefore, in general the GCD step dominates the complexity. In Table 3 Bariant’s attack is evaluated for sample we round numbers.
+Note that the GCD step implicitly assumes that $2^{r_{C}+r_{E}-1}<q$, though if the degree of the Ciminion function exceeds $q$, then an exhaustive search over the truncated output always performs better than the GCD computation.
+3.4.3 Eigenvalue Method
+For standard Ciminion it is sufficient to find a $\mathbb{F}_{q}$-valued root for $x$, because we can then invert $p_{C}\circ p_{E}$ to obtain a key guess.
+Obviously, our Ciminion DRL Gröbner basis (Theorem 3.1) satisfies the structural assumption from Section 2.5.1, so the complexity of constructing a univariate polynomial in $x$ is, see Equation 20,
+$\mathcal{O}\left(2^{\omega\cdot(r_{C}+r_{E}-2)+1}\right).$ (41)
+However, for standard Ciminion this complexity is always dominated by Bariant’s attack, see Section 3.4.2.
+For Ciminion2 on the other hand solving for $x$ is not sufficient, because we cannot invert $p_{C}\circ p_{E}$ anymore due to additional key additions.
+Therefore, for Ciminion2 we have to find a $\mathbb{F}_{q}$-valued solution for all quadratic variables to derive a key guess.
+Since the Ciminion2 DRL Gröbner basis satisfies the structure of Equation 22, the complexity of the Eigenvalue Method is given by Equation 25
+$\mathcal{O}\left(\begin{rcases}\begin{dcases}2^{\omega+1}\cdot\frac{2^{\omega\cdot(r_{C}+r_{E}-1)}-N^{\frac{r_{C}+r_{E}-1}{2}}}{2^{2\cdot\omega}-N},&2\mid(r_{C}+r_{E}-1),\\ 2^{2\cdot\omega+1}\cdot\frac{2^{\omega\cdot(r_{C}+r_{E}-2)}-N^{\frac{r_{C}+r_{E}-2}{2}}}{2^{2\cdot\omega}-N},&2\nmid(r_{C}+r_{E}-1)\end{dcases}\end{rcases}\right),$ (42)
+where $N$ is a bound on the number of $\mathbb{F}_{q}$-valued solutions in every iteration.
+In addition, we absorbed the terms for the GCDs with the field equations into the implied constant, because these steps are always dominated by the construction of the respective characteristic polynomial.
+Since a bound on $N$ is in general not known we consider a worst-case scenario approach.
+We assume that the first $\mathbb{F}_{q}$-valued solution that an adversary recovers in every iteration leads to the true key, i.e.
+we assume that $N=1$ in Equation 42.
+In Table 3 we evaluated the discussed Ciminion complexity estimations.
+All our estimations are from a designer’s point of view, which assume an ideal adversary who has access to a hypothetical matrix multiplication algorithm which achieves $\omega=2$.
+As already mentioned, for standard Ciminion Bariant’s attack is the most performative attack, and at least $r_{C}+r_{E}\geq 112$ is required to achieve $128$ bits of security.
+For Ciminion2 on the other hand this attack is not feasible, so we have to fall back to the Eigenvalue Method.
+Then, we require at least $r_{C}+r_{E}\geq 66$ to achieve $128$ bits of security.
+Finally, the estimation of the Ciminion designers grossly underestimates the capabilities of Gröbner basis attacks.
+After all, we can construct an iterated Ciminion DRL Gröbner basis in negligible time, whereas the designers estimated this complexity to be in $\mathcal{O}\left(2^{2\cdot\omega\cdot(r_{C}+r_{E})}\right)$.
 Table 3: Complexity estimations for various Gröbner basis attacks on Ciminion and Ciminion2 with $q=2^{127}+45$ and $\omega=2$.
 Complexity (bits)
-$r_{C}$ $r_{E}$ Fast Polynomial Multiplication Eigenvalue Method DRL Computation Iterated Model DRL Computation Fully Substituted Model [DGGK21, §4.4]
-$30$ $6$ $49$ $68$ $134$ $72$
-$60$ $6$ $80$ $128$ $253$ $132$
-$90$ $6$ $111$ $188$ $372$ $192$
-$90$ $14$ $119$ $204$ $404$ $208$
-$90$ $23$ $128$ $222$ $440$ $226$
-Also, recall that $r_{C}\in\left\{s+6,\left\lceil\frac{2\cdot(s+6)}{3}\right\rceil\right\}$, where $s$ is the security level, for two Ciminion instantiations, see Table 2.
-In particular, for Ciminion2 this already implies that for the quotient space dimension $\dim_{\mathbb{F}_{q}}\left(\mathcal{F}_{\texttt{Ciminion}}\right)\geq 2^{\frac{s}{2}}$.
-Therefore, round numbers based on Table 2 should be sufficient to achieve the security level in Gröbner basis attacks via the iterated polynomial system, provided of course a fix against Bariant’s attack [Bar23] is implemented.
-7.2 Hydra
-For Gröbner basis cryptanalysis of Hydra the designers considered the downsized iterated Hydra polynomial system, see [GØSW23, §7.4] and [GØSW22, Appendix I.2].
-The designers assumed that the downsized system of $22\cdot r_{\mathcal{H}}+$ 2 equations in $22\cdot r_{\mathcal{H}}-$ 2 variables is semi-regular [Frö85, Par10].
-Under the semi-regularity assumption one can extract the degree of regularity ([BFS04, Definition 4]) $d_{\operatorname{reg}}$ of the Hydra polynomial system from the Hilbert series [KR05, Chapter 5] which then yields the complexity estimate ([BFS04, Theorem 7])
-$\mathcal{O}\left((2\cdot r_{\mathcal{H}}+2)\cdot d_{\operatorname{reg}}\cdot\binom{2\cdot r_{\mathcal{H}}-2+d_{\operatorname{reg}}}{d_{\operatorname{reg}}}^{\omega}\right)$ (76)
-for DRL Gröbner basis computations.
+$r_{C}+r_{E}$ Bariant’s Attack [Bar23] Eigenvalue Method Fully Substituted Model [DGGK21, §4.4]
+$32$ $45.60$ $61.09$ $126$
+$33$ $46.67$ $63.09$ $130$
+$65$ $80.18$ $127.09$ $258$
+$66$ $81.22$ $129.09$ $262$
+$111$ $127.45$ $219.09$ $442$
+$112$ $128.47$ $221.09$ $446$
+Recall from [DGGK21, Table 1] that $r_{C}\in\left\{s+6,\left\lceil\frac{2\cdot(s+6)}{3}\right\rceil\right\}$, where $s$ is the required security level.
+For Ciminion2 this already implies that $\dim_{\mathbb{F}_{q}}\left(\mathcal{F}_{\texttt{Ciminion2}}\right)\geq 2^{\frac{s}{2}}$, and the complexity of the Eigenvalue Method will scale approximately with $\gtrsim 2^{\omega\cdot\frac{s}{2}}\geq 2^{s}$.
+So, for now we conclude that Ciminion2 resists Gröbner basis attacks within the security level, even under ideal assumptions on the adversary.
+4 Analysis of Hydra
+The Hydra heads are based on the Lai–Massey permutation. For the Gröbner basis analysis we recall that the Lai–Massey permutation
+$\mathcal{L}:\begin{pmatrix}x\\ y\end{pmatrix}\mapsto\begin{pmatrix}x+F(x-y)\\ y+F(x-y)\end{pmatrix}$ (43)
+is linearly equivalent to the Feistel permutation
+$\mathcal{F}:\begin{pmatrix}x\\ y\end{pmatrix}\mapsto\begin{pmatrix}x\\ y+F(x)\end{pmatrix}$ (44)
+via the matrices $\mathbf{A}_{1}=\begin{pmatrix}1&-1\\ 0&1\end{pmatrix}$ and $\mathbf{A}_{2}=\begin{pmatrix}1&1\\ 0&1\\ \end{pmatrix}$, i.e.
+$\mathcal{L}=\mathbf{A}_{2}\circ\mathcal{F}\circ\mathbf{A}_{1}$.
+Inspired by this equivalence we can reduce the Hydra polynomial system to only one non-linear equation per round, i.e.
+the Hydra polynomial system transforms to a Feistel-like polynomial system.
+In [Ste24, §6.2] the iterated polynomial systems representing generalized Feistel ciphers were analyzed.
+For these polynomial systems linear systems were derived, and in case the systems are of full rank the iterated polynomial systems are in generic coordinates, see [Ste24, Theorem 6.5].
+For a system in generic coordinates we then have a proven upper bound on the solving degree, see Theorem 2.6.
+Based on the aforementioned linear equivalence, we are going to replicate this strategy to construct a linear system for Hydra.
+Further, if the Hydra linear system has full rank, then we can transform the system into a DRL Gröbner basis via a linear change of coordinates.
+4.1 Generic Coordinates for Hydra
+Recall that the Hydra iterated polynomial system (Definition 2.2) consists of $816\cdot r_{\mathcal{H}}+$ 8 polynomials in $416\cdot r_{\mathcal{H}}+$ 4 variables. Let
+$\mathbf{A}_{n}=\begin{pmatrix}1&0&\ldots&0&-1\\ 0&1&\ldots&0&-1\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\ldots&1&-1\\ 0&0&\ldots&0&1\end{pmatrix}\in\operatorname{GL}_{n}\left(\mathbb{F}_{q}\right),$ (45)
+and let $\mathbf{B}=\operatorname{diag}\left(\mathbf{A}_{4},\mathbf{A}_{4}\right)\in\operatorname{GL}_{8}\left(\mathbb{F}_{q}\right)$.
+Then the first seven components of $\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{j}^{(i)}$, where $1\leq i\leq r_{\mathcal{H}}$ and $j\in\{1,2\}$, are affine and the eighth component is quadratic.
+Moreover, the fourth and the eighth components of $\mathbf{B}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{f}_{\mathcal{R}}$ are quadratic while the other ones are affine.
+In particular, by extracting the $2\cdot 7\cdot r_{\mathcal{H}}$ affine components of the $\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{j}^{(i)}$’s we can eliminate the variables
+$y_{1},\dots,y_{4},z_{1},\dots,z_{3},x_{j,1}^{(i)},\dots,x_{j,7}^{(i)},$ (46)
+where $10\leq i\leq r_{\mathcal{H}}-$ 1 and $j\in\{1,2\}$, in the quadratic equations.
+(This can be explicitly seen in the proof of Theorem 4.1.) Thus, after the elimination we reduce to $82\cdot r_{\mathcal{H}}+$ 8 polynomials, of which $6$ are also affine, in $42\cdot r_{\mathcal{H}}+$ 4 variables.
+Next we derive a linear system to verify being in generic coordinates for Hydra.
+Theorem 4.1.
+Let $\mathbb{F}_{q}$ be a finite field, let $\overline{\mathbb{F}_{q}}$ be its algebraic closure, let
+$\mathcal{F}_{\textsf{Hydra}}=\left\{\mathbf{f}_{1}^{(i)},\mathbf{f}_{\mathcal{R}},\mathbf{f}_{2}^{(i)}\right\}_{1\leq i\leq r_{\mathcal{H}}}\subset P=\overline{\mathbb{F}_{q}}\Big{[}\mathbf{y},\mathbf{z},\mathbf{x}_{1}^{(1)},\dots,\mathbf{x}_{1}^{(r_{\mathcal{H}}-1)},\mathbf{x}_{2}^{(0)},\dots,\mathbf{x}_{2}^{(r_{\mathcal{H}}-1)},\mathbf{k}\Big{]}$
+be a Hydra polynomial system, and let
+$\mathcal{G}_{\textsf{Hydra}}=\left\{\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{1}^{(i)},\mathbf{B}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{f}_{\mathcal{R}},\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{2}^{(i)}\right\}_{1\leq i\leq r_{\mathcal{H}}}.$
+Then every homogeneous ideal $I\subset P[x_{0}]$ such that $x_{0}\notin\sqrt{I}$ and $\mathcal{G}_{\textsf{Hydra}}^{\text{hom}}\subset I$ is in generic coordinates if the following linear system has rank $416\cdot r_{\mathcal{H}}+$ 4
+$y_{1}-z_{4}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{1}^{(1)}\Big{)}\bigg{)}_{1}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{1}^{(1)}-\mathbf{k}^{\prime}\Big{)}\bigg{)}_{8}$ $=0,$
+$\vdots$
+$z_{3}-z_{4}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{1}^{(1)}\Big{)}\bigg{)}_{7}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{1}^{(1)}-\mathbf{k}^{\prime}\Big{)}\bigg{)}_{8}$ $=0,$
+$\sum_{l=1}^{4}y_{l}-z_{l}$ $=0,$
+$x_{2,k}^{(0)}-x_{2,8}^{(0)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{2}^{(1)}\Big{)}\bigg{)}_{k}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{2}^{(1)}-\mathbf{k}^{\prime}\Big{)}\bigg{)}_{8}$ $=0,$
+$\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{2,l}^{(0)}$ $=0,$
+$x_{j,k}^{(i-1)}-x_{j,8}^{(i-1)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{j}^{(i)}\Big{)}\bigg{)}_{k}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{j}^{(i)}-\mathbf{k}^{\prime}\Big{)}\bigg{)}_{8}$ $=0,$
+$\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{j,l}^{(i-1)}$ $=0,$
+$x_{1,k}^{(r_{\mathcal{H}}-1)}-x_{1,8}^{(r_{\mathcal{H}}-1)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+(\mathbf{y},\mathbf{z})^{\intercal}\Big{)}\bigg{)}_{k}-\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+(\mathbf{y},\mathbf{z})^{\intercal}\Big{)}\bigg{)}_{8}$ $=0,$
+$\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{1,l}^{(r_{\mathcal{H}}-1)}$ $=0,$
+$x_{2,k}^{(r_{\mathcal{H}}-1)}-x_{2,8}^{(r_{\mathcal{H}}-1)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+\mathbf{x}_{2}^{(0)}\Big{)}\bigg{)}_{k}-\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+\mathbf{x}_{2}^{(0)}\Big{)}\bigg{)}_{8}$ $=0,$
+$\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{2,l}^{(r_{\mathcal{H}}-1)}$ $=0,$
+$y_{1}-y_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{1}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{4}$ $=0,$
+$\vdots$
+$y_{3}-y_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{3}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{4}$ $=0,$
+$z_{1}-z_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{5}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{8}$ $=0,$
+$\vdots$
+$z_{3}-z_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{7}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{8}$ $=0,$
+where $12\leq i\leq r_{\mathcal{H}}-$ 1, $j\in\{1,2\}$ and $1\leq k\leq 7$.
+Proof.
+By [Ste24, Theorem 3.2] $\left(\mathcal{G}_{\textsf{Hydra}}^{\text{hom}}\right)$ is in generic coordinates if
+$\sqrt{\mathcal{G}_{\textsf{Hydra}}^{\text{top}}}=\left(\mathbf{y},\mathbf{z},\mathbf{x}_{1}^{(1)},\dots,\mathbf{x}_{1}^{(r_{\mathcal{H}}-1)},\mathbf{x}_{2}^{(0)},\dots,\mathbf{x}_{2}^{(r_{\mathcal{H}}-1)},\mathbf{k}\right).$ (47)
+To shorten some writing let
+$\mathbf{g}_{j}^{(i)}$ $=\left(\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{j}^{(i)}\right)^{\text{hom}}\mod(x_{0}),$
+$\mathbf{g}_{\mathcal{R}}$ $=\left(\mathbf{B}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{f}_{\mathcal{R}}\right)^{\text{hom}}\mod(x_{0}).$
+Then for $i=1$ we have that
+$\mathbf{g}_{1}^{(1)}$ =($=\begin{pmatrix}\left\{y_{k}-z_{4}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{1}^{(1)}\Big{)}\bigg{)}_{k}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{1}^{(1)}-\mathbf{k}^{\prime}\Big{)}\bigg{)}_{8}\right\}_{1\leq k\leq 7}\\[10.0pt] \left(\sum_{l=1}^{4}y_{l}-z_{l}\right)^{2}\end{pmatrix},$
+$\mathbf{g}_{2}^{(1)}$ =($=\begin{pmatrix}\left\{x_{2,k}^{(0)}-x_{2,8}^{(0)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{2}^{(1)}\Big{)}\bigg{)}_{k}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{2}^{(1)}-\mathbf{k}^{\prime}\Big{)}\bigg{)}_{8}\right\}_{1\leq k\leq 7}\\[10.0pt] \left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{2,l}^{(0)}\right)^{2}\end{pmatrix}.$
+For $12\leq i\leq r_{\mathcal{H}}-$ 1 and $j\in\{1,2\}$ we have that
+$\mathbf{g}_{j}^{(i)}=\begin{pmatrix}\left\{x_{j,k}^{(i-1)}-x_{j,8}^{(i-1)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{j}^{(i)}\Big{)}\bigg{)}_{k}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{j}^{(i)}-\mathbf{k}^{\prime}\Big{)}\bigg{)}_{8}\right\}_{1\leq k\leq 7}\\[10.0pt] \left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{j,l}^{(i-1)}\right)^{2}\end{pmatrix}.$
+For $i=r_{\mathcal{H}}$ we have that
+$\mathbf{g}_{1}^{(r_{\mathcal{H}})}=$
+($\begin{pmatrix}\left\{x_{1,k}^{(r_{\mathcal{H}}-1)}-x_{1,8}^{(r_{\mathcal{H}}-1)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+(\mathbf{y},\mathbf{z})^{\intercal}\Big{)}\bigg{)}_{k}-\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+(\mathbf{y},\mathbf{z})^{\intercal}\Big{)}\bigg{)}_{8}\right\}_{1\leq k\leq 7}\\[10.0pt] \left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{1,l}^{(r_{\mathcal{H}}-1)}\right)^{2}\end{pmatrix},$
+$\mathbf{g}_{2}^{(r_{\mathcal{H}})}=$
+($\begin{pmatrix}\left\{x_{2,k}^{(r_{\mathcal{H}}-1)}-x_{2,8}^{(r_{\mathcal{H}}-1)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+\mathbf{x}_{2}^{(0)}\Big{)}\bigg{)}_{k}-\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+\mathbf{x}_{2}^{(0)}\Big{)}\bigg{)}_{8}\right\}_{1\leq k\leq 7}\\[10.0pt] \left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{2,l}^{(r_{\mathcal{H}}-1)}\right)^{2}\end{pmatrix}.$
+Finally, for the rolling function we have that
+$\mathbf{g}_{\mathcal{R}}=\begin{pmatrix}\left\{y_{k}-y_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{k}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{4}\right\}_{1\leq k\leq 3}\\[10.0pt] g(\mathbf{y},\mathbf{z})\\ \left\{z_{k}-z_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{4+k}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{8}\right\}_{1\leq k\leq 3}\\[10.0pt] g(\mathbf{z},\mathbf{y})\end{pmatrix}.$
+Except for the rolling function, all quadratic polynomials are squares of linear polynomials.
+Collecting the linear polynomials from $\mathcal{G}_{\textsf{Hydra}}^{\text{top}}$ together with the linear polynomials from the squares, we yield the linear system from the assertion.
+It is clear that all these linear polynomials are elements of $\sqrt{\mathcal{G}_{\textsf{Hydra}}^{\text{top}}}$.
+Since the highest degree components are homogeneous, we always have that $\sqrt{\mathcal{G}_{\textsf{Hydra}}^{\text{top}}}\subset\left(\mathbf{y},\mathbf{z},\mathbf{x}_{1}^{(i)},\dots,\mathbf{x}_{1}^{(r_{\mathcal{H}}-1)},\mathbf{x}_{2}^{(0)},\dots,\mathbf{x}_{2}^{(r_{\mathcal{H}}-1)},\mathbf{k}\right)$.
+Thus, if the linear system from the assertion has full rank, then we have established the equality from Equation 47.
+∎
+Naturally, this theorem can be extended to an arbitrary number of Hydra samples. Now it is straight-forward to compute the Macaulay bound (Theorem 2.6) for Hydra.
+Corollary 4.2.
+Let $\mathbb{F}_{q}$ be a finite field, let $\overline{\mathbb{F}_{q}}$ be its algebraic closure, let
+$\mathcal{F}_{\textsf{Hydra}}=\left\{\mathbf{f}_{1}^{(i)},\mathbf{f}_{\mathcal{R}},\mathbf{f}_{2}^{(i)}\right\}_{1\leq i\leq r_{\mathcal{H}}}\subset P=\overline{\mathbb{F}_{q}}\Big{[}\mathbf{y},\mathbf{z},\mathbf{x}_{1}^{(1)},\dots,\mathbf{x}_{1}^{(r_{\mathcal{H}}-1)},\mathbf{x}_{2}^{(0)},\dots,\mathbf{x}_{2}^{(r_{\mathcal{H}}-1)},\mathbf{k}\Big{]}$
+be a Hydra polynomial system. Let
+$\mathcal{G}_{\textsf{Hydra}}=\left\{\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{1}^{(i)},\mathbf{B}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{f}_{\mathcal{R}},\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{2}^{(i)}\right\}_{1\leq i\leq r_{\mathcal{H}}},$
+and assume that $\left(\mathcal{G}_{\textsf{Hydra}}^{\text{hom}}\right)$ is in generic coordinates.
+If the affine polynomials in $\mathcal{G}_{\textsf{Hydra}}$ have rank $62\cdot 7\cdot r_{\mathcal{H}}+$ 6, then
+$\operatorname{sd}_{DRL}\left(\mathcal{G}_{\textsf{Hydra}}\right)\leq 2\cdot r_{\mathcal{H}}.$
+Proof.
+If the affine polynomials have rank $62\cdot 7\cdot r_{\mathcal{H}}+$ 6, then we can reduce to $22\cdot r_{\mathcal{H}}+$ 2 quadratic polynomials in $22\cdot r_{\mathcal{H}}-$ 2 variables.
+This variable elimination does not affect being in generic coordinates, so the Macaulay bound (Theorem 2.6) evaluates to
+$\operatorname{sd}_{DRL}\left(\mathcal{G}_{\textsf{Hydra}}\right)\leq 2\cdot(2\cdot r_{\mathcal{H}}-1)-\left(2\cdot r_{\mathcal{H}}-1\right)+1=2\cdot r_{\mathcal{H}}.\qed$
+4.2 Extracting a Zero-Dimensional DRL Gröbner Basis for Hydra
+Due to the structure of the Hydra heads we can actually produce a DRL Gröbner basis via a simple linear change of coordinates.
+Again, let $\mathcal{G}_{\textsf{Hydra}}$ be as in Theorem 4.1, we assume that
+  1. (i)
+    the linear system from Theorem 4.1 has full rank, and
+  2. (ii)
+    the linear polynomials in $\mathcal{G}_{\textsf{Hydra}}^{\text{top}}$ have rank $62\cdot 7\cdot r_{\mathcal{H}}+$ 6.
+I.e., $\left(\mathcal{G}_{\textsf{Hydra}}^{\text{hom}}\right)$ is in generic coordinates, and we can use the linear polynomials in $\mathcal{G}_{\textsf{Hydra}}$ to reduce the system to $22\cdot r_{\mathcal{H}}+$ 2 quadratic polynomials in $22\cdot r_{\mathcal{H}}-$ 2 variables.
+By construction $2\cdot r_{\mathcal{H}}$ of these quadratic polynomials come from the Hydra heads and their quadratic terms are squares of linear equations.
+In particular, before the variable elimination the quadratic terms are, see Definition 2.2 and the proof of Theorem 4.1,
+$g_{1,8}^{(1)}$ $=\left(\sum_{l=1}^{4}y_{l}-z_{l}\right)^{2}$ $=\mathcal{L}_{1,1}^{2},$ (48)
+$g_{2,8}^{(1)}$ $=\left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{2,l}^{(0)}\right)^{2}$ $=\mathcal{L}_{2,1}^{2},$ (49)
+$g_{j,8}^{(i)}$ $=\left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{j,l}^{(i-1)}\right)^{2}$ $=\mathcal{L}_{j,i}^{2},\qquad 2\leq i\leq r_{\mathcal{H}},\ j\in\{1,2\},$ (50)
+Recall that these linear equations =$\mathcal{L}=\{\mathcal{L}_{1,i},\mathcal{L}_{2,i}\}_{1\leq i\leq 2\cdot r_{\mathcal{H}}}$ also appeared in the linear system from Theorem 4.1.
+By Assumption (i), $\mathcal{L}$ together with the linear equations from $\mathcal{G}_{\textsf{Hydra}}^{\text{top}}$ has rank $416\cdot r_{\mathcal{H}}+$ 4.
+After eliminating $62\cdot 7\cdot r_{\mathcal{H}}+$ 6 variables with the affine equations from $\mathcal{G}_{\textsf{Hydra}}$, the polynomials from Equations 48, 49 and 50 modify to
+$g_{j,8}^{(i)}=\left(\tilde{\mathcal{L}}_{j,i}+a_{j,i}\right)^{2},\qquad 1\leq i\leq r_{\mathcal{H}},\ j\in\{1,2\},$ (51)
+where ~$\tilde{\mathcal{L}}_{j,i}$ is a linear polynomial in the remaining $22\cdot r_{\mathcal{H}}-$ 2 variables and $a_{j,i}\in\mathbb{F}_{q}$.
+Moreover, the rank of a matrix does not depend on the order of Gaussian elimination, so by Assumption (i) the new linear polynomials ~={~$\tilde{\mathcal{L}}=\left\{\tilde{\mathcal{L}}_{1,i},\tilde{\mathcal{L}}_{2,i}\right\}_{1\leq i\leq 2\cdot r_{\mathcal{H}}}$ still must contain a maximal linearly independent subset of rank $22\cdot r_{\mathcal{H}}-$ 2.
+Let $\mathbf{x}_{\text{nl}}$ denote the remaining $22\cdot r_{\mathcal{H}}-$ 2 non-linear variables.
+We can now collect this maximal linearly independent subset of $\tilde{\mathcal{L}}$ in a matrix $\mathbf{M}\in\mathbb{F}_{q}^{(2\cdot r_{\mathcal{H}}-2)\times(2\cdot r_{\mathcal{H}}-2)}$.
+Columns of $\mathbf{M}$ are indexed by the elements of $\mathbf{x}_{\text{nl}}$, and in the rows of $\mathbf{M}$ we write the coefficients of the maximal linearly independent equations.
+Next we pick new variables $\mathbf{\hat{x}}=\left(\hat{x}_{1},\dots,\hat{x}_{2\cdot r_{\mathcal{H}}-2}\right)^{\intercal}$, and finally we perform a linear change of variables $\mathbf{\hat{x}}=\mathbf{M}\mathbf{x}_{\text{nl}}$.
+By construction $2\cdot r_{\mathcal{H}}$ polynomials in $\mathcal{G}_{\textsf{Hydra}}$ are of the form
+~$\tilde{\mathcal{L}}_{j,i}(\mathbf{x}_{\text{nl}})^{2}+\mathcal{A}_{j,i}(\mathbf{x}_{\text{nl}})=0,$ (52)
+where $\mathcal{A}_{j,i}\in\mathbb{F}_{q}[\mathbf{x}_{\text{nl}}]$ is affine, $1\leq i\leq 2\cdot r_{\mathcal{H}}$ and $j\in\{1,2\}$.
+Therefore, after the change of coordinates we produce $22\cdot r_{\mathcal{H}}-$ 2 polynomials of the form
+$g_{i}=\hat{x}_{i}^{2}+\sum_{j=1}^{2\cdot r_{\mathcal{H}}-2}\alpha_{i,j}\cdot\hat{x}_{j}+\alpha_{i,0},$ (53)
+where $\alpha_{i,j},\alpha_{i,0}\in\mathbb{F}_{q}$, together with four additional quadratic polynomials.
+In particular, $\hat{\mathcal{G}}=\left\{g_{i}\right\}_{1\leq i\leq 2\cdot r_{\mathcal{H}}-2}\subset\mathbb{F}_{q}\big{[}\mathbf{\hat{x}}\big{]}$ is a DRL Gröbner basis with $\dim_{\mathbb{F}_{q}}\left(\hat{\mathcal{G}}\right)=2^{2\cdot r_{\mathcal{H}}-2}$.
+We summarize this construction in a more formal manner in the next corollary.
+Corollary 4.3.
+Let $\mathbb{F}_{q}$ be a finite field, let
+$\mathcal{F}_{\textsf{Hydra}}=\left\{\mathbf{f}_{1}^{(i)},\mathbf{f}_{\mathcal{R}},\mathbf{f}_{2}^{(i)}\right\}_{1\leq i\leq r_{\mathcal{H}}}\subset P=\mathbb{F}_{q}\Big{[}\mathbf{y},\mathbf{z},\mathbf{x}_{1}^{(1)},\dots,\mathbf{x}_{1}^{(r_{\mathcal{H}}-1)},\mathbf{x}_{2}^{(0)},\dots,\mathbf{x}_{2}^{(r_{\mathcal{H}}-1)},\mathbf{k}\Big{]}$
+be a Hydra polynomial system, and let
+$\mathcal{G}_{\textsf{Hydra}}$ =$=\left\{\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{1}^{(i)},\mathbf{B}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{f}_{\mathcal{R}},\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{2}^{(i)}\right\}_{1\leq i\leq r_{\mathcal{H}}},$
+$\mathcal{G}_{lin}$ =$=\left\{g\in\mathcal{G}_{\textsf{Hydra}}\mid\deg\left(g\right)=1\right\},$
+$\mathcal{G}_{squ}$ =$=\left\{g\in\mathcal{G}_{\textsf{Hydra}}\mid\deg\left(g\right)=2\right\}.$
+Assume that
+  1. (i)
+    the linear system from Theorem 4.1 has full rank, and
+  2. (ii)
+    $\operatorname{rank}\left(\mathcal{G}_{lin}^{\text{top}}\right)=2\cdot 7\cdot r_{\mathcal{H}}+6$.
+[…]
+  1. (1)
+[…]
+  2. (2)
+    For every $21\leq i\leq 2\cdot r_{\mathcal{H}}-$ 2 there exists $g_{i}\in\hat{\mathcal{G}}_{squ}$ such that
+    $g_{i}=\hat{x}_{i}^{2}+\mathcal{A}_{i},$
+[…]
+  3. (3)
+    $\dim_{\mathbb{F}_{q}}\left(\mathcal{G}_{lin}\cup\mathcal{G}_{sub}\cup\hat{\mathcal{G}}_{squ}\right)\leq 2^{2\cdot r_{\mathcal{H}}-2}$.
+Besides being tedious to do by hand, success of the change of coordinates depends on the matrices used in the Hydra heads.
+Therefore, we did not attempt a full proof by hand, instead for specific instances we outsourced it to a computer algebra program where the construction can be performed in a matter of seconds.
+4.3 Implementation
+We have implemented Hydra and its iterated polynomial system in the computer algebra systems SageMath [Sag24] and OSCAR [OSC24].
+[^6] https://github.com/sca-research/Groebner-Basis-Cryptanalysis-of-Ciminion-and-Hydra.git In both systems we implemented the linear system from Theorem 4.1 and the change of coordinates from Section 4.2.
+With the SageMath implementation we can compute the DRL Gröbner basis for full rounds $r_{\mathcal{H}}=39$ over $128$ bit prime fields in negligible time.
+The OSCAR computer algebra system offers an efficient F4 implementation that can handle primes of size $q<2^{31}$, therefore we recommend to use OSCAR for experimental computations over small prime fields.
+For completeness, we mention that we implemented the rolling function without a constant addition, i.e.
+we set $\mathbf{c}_{\mathcal{R}}^{(i)}=\mathbf{0}$ for all $i\geq 1$.
+However, we emphasize that the constant addition does not affect either the linear system from Theorem 4.1 or the change of coordinates from Section 4.2.
+Example 4.4 (A Concrete Instance).
+In [GØSW22, Appendix C] the Hydra designers propose concrete matrices over the prime field $q=2^{127}+45$
+$\mathbf{M}_{\mathcal{E}}$ $=\operatorname{circ}\left(3,2,1,1\right),\qquad\qquad\mathbf{M}_{\mathcal{I}}=\begin{pmatrix}1&1&1&1\\ 1&4&1&1\\ 3&1&3&1\\ 4&1&1&2\end{pmatrix},$ (54)
+$\mathbf{M}_{\mathcal{J}}$ $=\begin{pmatrix}3&1&1&1&1&1&1&1\\ 7&3&1&1&1&1&1&1\\ 4&1&4&1&1&1&1&1\\ 3&1&1&8&1&1&1&1\\ 7&1&1&1&7&1&1&1\\ 8&1&1&1&1&5&1&1\\ 5&1&1&1&1&1&2&1\\ 4&1&1&1&1&1&1&6\end{pmatrix}.$ (55)
+For security level $\kappa=128$ bits the designers suggest to use $r_{\mathcal{H}}=39$, see [GØSW23, §5.5].
+For this instance using our SageMath implementation we verified that the Hydra polynomial system is in generic coordinates for $m\in\{2,3,4,5,10\}$ many samples.
+For $m=2$ we can compute a Hydra DRL Gröbner basis with the change of coordinates from Section 4.2.
+Recall the quadratic terms from the Hydra heads, see Equations 48, 49 and 50.
+For $r_{\mathcal{H}}\geq 2$ and after the elimination of $62\cdot 7\cdot r_{\mathcal{H}}+$ 6 variables, see Equation 51, we observed that
+$\operatorname{rank}\left(\tilde{\mathcal{L}}_{1,8}^{(3)},\dots,\tilde{\mathcal{L}}_{1,8}^{(r_{\mathcal{H}})},\tilde{\mathcal{L}}_{2,8}^{(1)},\dots,\tilde{\mathcal{L}}_{2,8}^{(r_{\mathcal{H}})}\right)=2\cdot r_{\mathcal{H}}-2.$ (56)
+So, before the elimination of $62\cdot 7\cdot r_{\mathcal{H}}+$ 6 variables our change of coordinates corresponds to, see Equations 48, 49 and 50,
+$\hat{x}_{i}$ $=\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{1,l}^{(i-1)},\qquad$ $3\leq i\leq r_{\mathcal{H}},$ (57)
+$\hat{x}_{i+r_{\mathcal{H}}-2}$ $=\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{2,l}^{(j-1)},\qquad$ $1\leq i\leq r_{\mathcal{H}}.$ (58)
+We also note that we did not find a round number $r_{\mathcal{H}}\geq 2$ for which the Gröbner basis extraction fails.
+4.4 Cryptanalysis
+Next we discuss the cryptanalytic impact of our theoretical analysis.
+Throughout, this section we will always assume that we can perform the linear change of coordinates from Section 4.2.
+In particular, we stress that for the Hydra instance from Example 4.4 we are always able to construct a DRL Gröbner basis.
+Therefore, we have a quadratic Hydra DRL Gröbner basis in $22\cdot r_{\mathcal{H}}-$ 2 equations and variables together with four additional quadratic polynomials at hand.
+This opens the door for two different attack strategies:
+  1. (A)
+    Recomputation of the DRL Gröbner basis for the overdetermined system of $22\cdot r_{\mathcal{H}}+$ 2 quadratic polynomials.
+  2. (B)
+    Polynomial system solving techniques for the DRL Gröbner basis of $22\cdot r_{\mathcal{H}}-$ 2 quadratic polynomials.
+For comparison, we also review the complexity estimation of the designers.
+4.4.1 Estimation of the Designers
+The Hydra designers considered the iterated polynomial model, see [GØSW23, §7.4] and [GØSW22, Appendix I.2], and reduced it to $22\cdot r_{\mathcal{H}}+$ 2 quadratic polynomials in $22\cdot r_{\mathcal{H}}-$ 2 variables.
+The designers assumed that this downsized system is semi-regular [BFS04].
+Then, the degree of regularity $d_{\operatorname{reg}}$ is given by the index of the first negative coefficient in the power series
+$H(t)=\frac{(1-t^{2})^{2\cdot r_{\mathcal{H}}+2}}{(1-t)^{2\cdot r_{\mathcal{H}}-2}},$ (59)
+and the complexity of computing the DRL Gröbner basis is then bounded by [BFS04, Theorem 7]
+$\mathcal{O}\left(\binom{2\cdot r_{\mathcal{H}}-2+d_{\operatorname{reg}}}{d_{\operatorname{reg}}}^{\omega}\right).$ (60)
 For $r_{\mathcal{H}}^{\ast}=29$ the designers estimate $d_{\operatorname{reg}}=23$, with $\omega=2$ this yields a complexity estimate of $\approx 2^{130.8}$ bits.
 Moreover, it is claimed that $r_{\mathcal{H}}=r_{\mathcal{H}}^{\ast}+2=31$ rounds are sufficient to provide $128$ bits of security against Gröbner basis attacks.
-7.2.1 Term Order Conversion & Eigenvalue Method Complexity Estimate
-In Section 4.2 we have outlined how one can extract a zero-dimensional DRL Gröbner basis from the Hydra polynomial system via a linear change of coordinates.
-We can directly use this Gröbner basis to perform term order conversion to LEX, or apply the dedicated Eigenvalue Method from Section 5 combined with $\mathbb{F}_{q}$-valued root extraction via the GCD.
-Since we could not prove that a LEX Hydra Gröbner basis is in shape position, we have to use the complexity for the probabilistic FGLM [FGHR14] of radical ideals which becomes
-$\tilde{\mathcal{O}}\left((r_{\mathcal{H}}-1)\cdot 2^{\omega\cdot 2\cdot(r_{\mathcal{H}}-1)+1}\right).$ (77)
-For the Eigenvalue Method, the Hydra DRL Gröbner basis satisfies the assumptions of Section 5, and in addition it satisfies the shape condition of Equation 58.
-Therefore, we can use the complexity estimate from Equation 61.
-Since we do not know the number of $\mathbb{F}_{q}$-valued eigenvalues for the Hydra multiplication matrices we opt for a better safe than sorry approach.
-We assume that the first eigenvalue that an adversary recovers for each variable leads to the true key, i.e.
-we assume $N=1$.
-Then, the Eigenvalue Method complexity becomes
-$\mathcal{O}\left(2^{\omega}\cdot\frac{2^{\omega\cdot(2\cdot r_{\mathcal{H}}-2)}-1}{2^{2\cdot\omega}-1}+\sum_{i=1}^{r_{\mathcal{H}}-1}D\big{(}2^{2\cdot i},q\big{)}\right),$ (78)
-where $D\big{(}2^{i},q\big{)}$ again denotes the complexity of the GCD with the field equation (Equation 1).
-In Table 4 we have computed the complexities for term order conversion and efficient $\mathbb{F}_{q}$-valued root extraction via the GCD.
-From the formulae, we can directly see that for $\omega=2$ and $r_{\mathcal{H}}\geq 34$ the complexities will always exceed $128$ bits.
-On the other, hand for $r_{\mathcal{H}}^{\ast}=29$ term order conversion requires only $118$ bits, and for $r_{\mathcal{H}}=r_{\mathcal{H}}^{\ast}+2=31$ it requires $126$ bits.
-This invalidates the claim by the Hydra designers that $31$ rounds are sufficient to protect against Gröbner basis attacks.
-Moreover, term order conversion always outperforms DRL Gröbner basis computation under the semi-regularity assumption.
-Lastly, for the round numbers in Table 4 the Eigenvalue Method shaves-off at least $6$ additional bits compared to term order conversion.
-Table 4: Complexity estimate of the Eigenvalue Method and term order conversion to LEX with efficient $\mathbb{F}_{q}$-valued root extraction for Hydra with $q=2^{127}+45$ and $\omega=2$.
+4.4.2 Recomputing the DRL Gröbner Basis
+After constructing the Hydra DRL Gröbner basis we have one big advantage compared to the designer estimation: We yield a $\mathbb{F}_{q}$-Boolean polynomial system, see Section 2.4.
+Thus, in the Macaulay matrix we only have to consider square-free monomials.
+Effectively, this saves time and space.
+Proven Solving Degree Bound.
+Recall from Corollary 4.2 that $\operatorname{sd}_{DRL}\left(\mathcal{G}_{\textsf{Hydra}}\right)\leq 2\cdot r_{\mathcal{H}}$.
+So, in the worst case we have to construct the full $\mathbb{F}_{q}$-Boolean Macaulay matrix, see Theorem 2.10.
+The complexity for constructing the $\mathbb{F}_{q}$-Boolean Macaulay matrix was given in Corollary 2.12, for the Hydra parameters $\left|F\right|=4$, $n,d=2\cdot r_{\mathcal{H}}-2$ and $D=2$ we yield the complexity
+$\mathcal{O}\left(4\cdot r_{\mathcal{H}}\cdot(2\cdot r_{\mathcal{H}}-1)\cdot\sum_{i=0}^{2\cdot r_{\mathcal{H}}-2}\binom{2\cdot r_{\mathcal{H}}+i}{i+2}\cdot\binom{2\cdot r_{\mathcal{H}}-2}{i}\right).$ (61)
+Further, by Theorem 2.10 Gaussian elimination on the $\mathbb{F}_{q}$-Boolean Macaulay matrix can be performed in $\mathcal{O}\left(4\cdot 2^{\omega\cdot(2\cdot r_{\mathcal{H}}-2)}\right)$ field operations.
+In Table 4 we evaluated these complexities for sample round numbers.
+Semi-Regularity Assumption.
+If we assume that the Hydra polynomial system is semi-regular after the change of variables (Section 4.2), then Equation 61 improves to
+$\mathcal{O}\left(4\cdot r_{\mathcal{H}}\cdot(2\cdot r_{\mathcal{H}}-1)\cdot\sum_{i=0}^{d_{\operatorname{reg}}-2}\binom{2\cdot r_{\mathcal{H}}-2}{i}\cdot\binom{2\cdot r_{\mathcal{H}}+i}{i+2}\right),$ (62)
+and the complexity for Gaussian elimination improves to
+$\mathcal{O}\left(4\cdot\left(\sum_{i=0}^{d_{\operatorname{reg}}-2}\binom{2\cdot r_{\mathcal{H}}-2}{i}\right)\cdot\left(\sum_{i=0}^{d_{\operatorname{reg}}}\binom{2\cdot r_{\mathcal{H}}-2}{i}\right)^{\omega-1}\right).$ (63)
+In Table 4 we evaluated the complexities together with the one from the designers (Section 4.4.1).
+Construction of the $\mathbb{F}_{q}$-Boolean Macaulay matrix dominates this approach.
+As expected, under the semi-regularity assumption the Boolean approach is slightly more performative than the designers’ estimation.
+In particular, an ideal adversary with $\omega=2$ could break up to $r_{\mathcal{H}}=29$ rounds below $128$ bits.
+Moreover, if the $\mathbb{F}_{q}$-Boolean Macaulay matrix can be constructed with negligible cost, then an ideal adversary could break up to $r_{\mathcal{H}}=34$ rounds.
+Table 4: Hydra complexity estimations for recomputing the DRL Gröbner basis.
+The column $d_{\operatorname{reg}}$ shows the degree of regularity under the semi-regularity assumption.
+All estimations use $\omega=2$.
 Complexity (bits)
-$r_{\mathcal{H}}$ Eigenvalue Method Term Order Conversion Semi-Regular Estimate [GØSW23]
-$20$ $75$ $82$ $89$
-$29$ $111$ $118$ $131$
-$31$ $119$ $126$ $142$
-$33$ $127$ $135$ $153$
-$34$ $131$ $139$ $159$
-$39$ $151$ $159$ $183$
-7.2.2 Proven DRL Complexity Estimate
-Via Theorem 4.1 and our SageMath implementation (Section 4.3) we obtain the proven DRL solving degree estimate for the Hydra polynomial instance from Section 4.3.1
-$\operatorname{sd}_{DRL}\left(\mathcal{F}_{\textsf{Hydra}}\right)\leq 2\cdot r_{\mathcal{H}}.$ (79)
-Since the Hydra polynomial system contains a zero-dimensional DRL Gröbner basis after a linear change of coordinates (Section 4.2), we can reduce Gaussian elimination on the full Macaulay matrix to Gaussian elimination on the $\mathbb{F}_{q}$-Boolean Macaulay matrix.
-With Theorem 6.4 and Equation 65 the complexity of constructing the $\mathbb{F}_{q}$-Boolean Macaulay matrix and Gaussian elimination for Hydra becomes
-$\mathcal{O}\left(\left(8\cdot r_{\mathcal{H}}^{2}-4\cdot r_{\mathcal{H}}\right)\cdot\left(\sum_{i=0}^{2\cdot r_{\mathcal{H}}-2}\binom{2\cdot r_{\mathcal{H}}+i-3}{i}\cdot\binom{2\cdot r_{\mathcal{H}}-2}{i}\right)+2^{\omega\cdot 2\cdot(r_{\mathcal{H}}-1)+2}\right).$ (80)
-In Table 5 we evaluate our complexity bound.
-Table 5: DRL Gröbner basis complexity estimate for Hydra via Equation 80 with $\omega=2$.
-Complexity (Bits)
-$r_{\mathcal{H}}$ Proven Estimate Semi-Regular Estimate [GØSW23]
-$20$ $105$ $89$
-$29$ $151$ $131$
-$31$ $161$ $142$
-$33$ $172$ $153$
-$34$ $177$ $159$
-$39$ $203$ $183$
-Therefore, we have derived a computer aided proof for the complexity of Gröbner basis attacks on Hydra without resorting to unproven assumptions like semi-regularity.
-Our proven complexity estimate is significantly larger than the hypothesized one for Hydra, since the degree of regularity estimate is always less than the Macaulay bound.
-7.2.3 Semi-Regular Complexity Estimate for the Boolean Macaulay Matrices
-As already mentioned, in the Hydra cryptanalysis it is assumed that the polynomial system of $22\cdot r_{\mathcal{H}}+$ 2 equations in $22\cdot r_{\mathcal{H}}-$ 2 variables is semi-regular.
-If we assume that the substituted Hydra polynomial system that contains the DRL Gröbner basis (Section 4.2) is also semi-regular, then we can again improve from the generic semi-regular complexity estimate (Equation 76) by passing to the $\mathbb{F}_{q}$-Boolean Macaulay matrix.
-Again, we can then use Equation 65 to obtain the complexity of construction and Gaussian elimination on the $\mathbb{F}_{q}$-Boolean Macaulay matrices
-$\mathcal{O}\left(\begin{aligned} \left(8\cdot r_{\mathcal{H}}^{2}-4\cdot r_{\mathcal{H}}\right)&\cdot\left(\sum_{i=0}^{d_{\operatorname{reg}}}\binom{2\cdot r_{\mathcal{H}}+i-3}{i}\cdot\binom{2\cdot r_{\mathcal{H}}-2}{i}\right)\\ &+4\cdot\left(\sum_{i=0}^{d_{\operatorname{reg}}}\binom{2\cdot r_{\mathcal{H}}-2}{i}\right)^{\omega}\end{aligned}\right).$ (81)
-In Table 6 we evaluate this complexity estimate.
-In addition, we show the complexity just for Gaussian elimination on the $\mathbb{F}_{q}$-Boolean Macaulay matrices, i.e.
-the complexity under the assumption that the Boolean Macaulay matrices are constructed “for free”.
-If we account for the cost of constructing the Boolean Macaulay matrices, then this method performs worse than the optimized Eigenvalue Method.
-However, it still beats the cryptanalysis of the Hydra designers for larger round numbers.
-Though, if we just consider the Gaussian elimination step, then this method slightly improves upon the Eigenvalue method.
-Table 6: DRL Gröbner basis complexity estimate for Hydra via Equation 80 with $\omega=2$.
-Complexity (Bits)
-$r_{\mathcal{H}}$ $d_{\operatorname{reg}}$ Boolean Gaussian Elimination Boolean Semi-Regular Estimate Semi-Regular Estimate [GØSW23]
-$20$ $16$ $74$ $91$ $89$
-$29$ $23$ $108$ $130$ $131$
-$31$ $25$ $116$ $140$ $142$
-$33$ $27$ $125$ $149$ $153$
-$34$ $28$ $129$ $154$ $159$
-$39$ $32$ $148$ $176$ $183$
+Proven Solving Degree Semi-Regular
+$r_{\mathcal{H}}$ $d_{\operatorname{reg}}$ Boolean Matrix Construction Boolean Gaussian Elimination Boolean Matrix Construction Boolean Gaussian Elimination Semi-Regular Estimate [GØSW23]
+$28$ $22$ $148.72$ $110$ $118.28$ $102.10$ $125.31$
+$29$ $23$ $153.89$ $114$ $123.29$ $106.25$ $130.80$
+$30$ $24$ $159.05$ $118$ $128.27$ $110.40$ $136.29$
+$31$ $25$ $164.21$ $122$ $133.25$ $114.54$ $141.77$
+$32$ $26$ $169.37$ $126$ $138.21$ $118.66$ $147.26$
+$33$ $27$ $174.52$ $130$ $143.16$ $122.79$ $152.75$
+$34$ $28$ $179.67$ $134$ $148.10$ $126.90$ $158.23$
+$35$ $28$ $184.82$ $138$ $150.37$ $129.66$ $160.24$
+$39$ $32$ $205.41$ $154$ $170.12$ $146.16$ $182.22$
+$45$ $37$ $236.24$ $178$ $197.03$ $169.55$ $211.72$
 Empirical F4 Working Degree.
 The Hydra designers performed small scale Gröbner basis experiments for Hydra with the F4 [Fau99] algorithm, see [GØSW22, Table 4].
-We replicated their experiment with our OSCAR implementation over the prime field $\mathbb{F}_{7741}$.
-We report our results in Table 7, but as input system we use the substituted system from Section 4.2 which contains a zero-dimensional DRL Gröbner basis.
+With our OSCAR implementation we replicated these experiments to verify that F4’s highest working degree matches the semi-regular degree of regularity before and after the change of coordinates.
+Our results are recorded in Table 5.
 Our experiments have been performed on an AMD EPYC-Rome (48) CPU with 94 GB RAM.
-Like the Hydra designers we observed that the empirical F4 working degree meets the semi-regular degree of regularity estimate.
-Table 7: Highest working degree and running times for F4 in small scale Hydra polynomial systems over the prime field $\mathbb{F}_{7741}$.
-Our data Grassi et al.’s data [GØSW22, Table 4]
-$r_{\mathcal{H}}$ $3$ $4$ $5$ $6$ $7$ $8$ $9$ $10$ $6$ $7$ $8$ $9$ $10$ $11$
-Highest F4 working degree $3$ $3$ $4$ $5$ $5$ $6$ $7$ $8$ $5$ $5$ $6$ $7$ $8$ $8$
-Total time (s) $0.01$ $0.01$ $0.01$ $0.03$ $0.21$ $2.65$ $66.85$ $2295.7$ $0.1$ $0.9$ $13.4$ $293.8$ $8799.5$ $229606.8$
-Overall, we conclude that currently the most competitive algebraic attacks against Hydra directly apply polynomial system solving techniques to the zero-dimensional DRL Gröbner basis, see Table 4.
+In all our experiments F4’s highest working degree indeed matches the semi-regular degree of regularity.
+As expected, after the change of coordinates the computation is faster, because F4 will implicitly eliminate all monomials which are divisible by a square.
+Table 5: Hydra highest working degree and total running time of F4 over the prime field $\mathbb{F}_{7741}$.
+The column $d_{\operatorname{reg}}$ shows the degree of regularity under the semi-regularity assumption.
+Before Change of Coordinates After Change of Coordinates Grassi et al.’s data [GØSW22, Table 4]
+$r_{\mathcal{H}}$ $d_{\operatorname{reg}}$ Highest F4 working degree Total time (s) Highest F4 working degree Total time (s) Highest F4 working degree Total time (s)
+$3$ $3$ $3$ $0.01$ $3$ $0.01$
+$4$ $3$ $3$ $0.01$ $3$ $0.01$
+$5$ $4$ $4$ $0.02$ $4$ $0.01$
+$6$ $5$ $5$ $0.16$ $5$ $0.03$ $5$ $0.1$
+$7$ $5$ $5$ $1.35$ $5$ $0.25$ $5$ $0.9$
+$8$ $6$ $6$ $14.88$ $6$ $2.68$ $6$ $13.4$
+$9$ $7$ $7$ $384.13$ $7$ $67.19$ $7$ $293.8$
+$10$ $8$ $8$ $8676.30$ $8$ $3194.95$ $8$ $8799.5$
+$11$ $8$ $8$ $229606.8$
+4.4.3 Term Order Conversion & Eigenvalue Method
+With a Hydra DRL Gröbner basis at hand, see Section 4.2, we can proceed to direct system solving techniques.
+For example, we can use a generic algorithm like the state-of-the-art probabilistic FGLM algorithm [FGHR14] to convert to a LEX Gröbner basis, and then factor the univariate polynomial.
+Assuming that the Hydra DRL Gröbner basis is in shape position the complexity of term order conversion is bounded by [FGHR14, Proposition 3]
+$\mathcal{O}\Big{(}(2\cdot r_{\mathcal{H}}-2)\cdot 2^{2\cdot r_{\mathcal{H}}-2}\cdot\big{(}2^{(\omega-1)\cdot(2\cdot r_{\mathcal{H}}-2)}+(2\cdot r_{\mathcal{H}}-2)^{2}\cdot\log_{2}(2\cdot r_{\mathcal{H}}-2)\big{)}\Big{)}$ (64)
+field operations.
+Next we have to factor the univariate polynomial of degree $2^{2\cdot r_{\mathcal{H}}-2}$ via the GCD with the field equation.
+However, even for full rounds $r_{\mathcal{H}}=39$ this degree is well below $2^{128}$.
+Therefore, the GCD complexity from Equation 21 will not exceed $128$ bits, and henceforth can be ignored.
+On the other hand, the Hydra DRL Gröbner basis satisfies the structure assumption from Equation 22, so we can pass to the dedicated Eigenvalue Method to find a $\mathbb{F}_{q}$-valued solution.
+For Hydra, the complexity from Equation 25 then evaluates to
+$\mathcal{O}\left(2^{\omega+1}\cdot\frac{2^{\omega\cdot(2\cdot r_{\mathcal{H}}-2)}-N^{r_{\mathcal{H}}-1}}{2^{2\cdot\omega}-N}\right),$ (65)
+where $N$ is a bound on the number of $\mathbb{F}_{q}$-valued solutions in every iteration.
+Again, the GCD complexity can be considered negligible.
+Since a bound on $N$ is in general not available, we consider a worst-case scenario approach.
+We assume that the first root which an adversary recovers in every iteration leads to the true Hydra key, i.e.
+we use $N=1$ in Equation 65.
+In Table 6 we evaluated the complexities for direct system solving techniques and compared them to the other strategies.
+Our estimations are again from a designer’s point of view which assumes an ideal adversary who can achieve the linear algebra constant $\omega=2$.
+In this scenario, term order conversion breaks up to $r_{\mathcal{H}}=31$ rounds below $128$ bits.
+This invalidates the claim of the Hydra designers [GØSW23, §7.4] that $r_{\mathcal{H}}=31$ rounds are sufficient to provide $128$ bits of security against Gröbner basis attacks for an ideal adversary.
+Moreover, via the dedicated Eigenvalue Method up to $r_{\mathcal{H}}=33$ do not achieve $128$ bits of security.
+For full rounds $r_{\mathcal{H}}=39$ the security margin is reduced by $\approx 30$ bits, however the security claim is not affected.
+Let $r_{\mathcal{H}}^{\ast}$ be the minimum round number which achieves $128$ bits of security in a Gröbner basis attack, i.e.
+according to our analysis we have $r_{\mathcal{H}}^{\ast}=34$.
+Recall from Equation 12 that the head round number is set as $r_{\mathcal{H}}=\left\lceil 1.25\cdot\max\left\{24,2+r_{\mathcal{H}}^{\ast}\right\}\right\rceil$.
+Thus, our analysis suggests that this round number is increased to $r_{\mathcal{H}}=45$ to recover the security margin originally intended by the designers.
+In case one wants to be extra conservative, one can consider the construction of the $\mathbb{F}_{q}$-Boolean Macaulay matrix to be for free.
+Then, we have $r_{\mathcal{H}}^{\ast}=35$, see Table 4, which implies $r_{\mathcal{H}}=47$.
+Table 6: Hydra complexity estimations for term order conversion and the Eigenvalue Method.
+The column $d_{\operatorname{reg}}$ shows the degree of regularity under the semi-regularity assumption.
+All estimations use $\omega=2$.
+Complexity (Bits)
+$r_{\mathcal{H}}$ $d_{\operatorname{reg}}$ Term Order Conversion Eigenvalue Method Boolean Semi-Regular Estimate Semi-Regular Estimate [GØSW23]
+$28$ $22$ $113.75$ $107.09$ $118.28$ $125.31$
+$29$ $23$ $117.81$ $111.09$ $123.29$ $130.80$
+$30$ $24$ $121.86$ $115.09$ $128.27$ $136.29$
+$31$ $25$ $125.91$ $119.09$ $133.25$ $141.77$
+$32$ $26$ $129.95$ $123.09$ $138.21$ $147.26$
+$33$ $27$ $134.00$ $127.09$ $143.16$ $152.75$
+$34$ $28$ $138.04$ $131.09$ $148.10$ $158.23$
+$35$ $28$ $142.09$ $135.09$ $150.37$ $160.24$
+$39$ $32$ $158.25$ $151.09$ $170.12$ $182.22$
+$45$ $37$ $182.46$ $175.09$ $197.03$ $211.72$
+Small Scale Experiments.
+With our OSCAR implementation we computed the characteristic polynomial of the Hydra multiplication matrix $\mathbf{M}_{\hat{x}_{2\cdot r_{\mathcal{H}}-2}}$, see Equation 19, for small $r_{\mathcal{H}}$ and measured the time of the overall computation.
+In addition, we computed the polynomial matrix determinant from Equation 19 for uniformly random matrices $\mathbf{A}_{0},\mathbf{A}_{1}\in\mathbb{F}_{q}^{(2\cdot r_{\mathcal{H}}-3)\times(2\cdot r_{\mathcal{H}}-3)}$.
+For construction and determinant of the matrix we used OSCAR’s functions for multivariate polynomial division and determinant respectively.
+Our results are recorded in Table 7.
+In addition, we performed term order conversion to LEX for our Hydra DRL Gröbner basis with OSCAR’s fglm function.
+For comparison, we also generated a random multivariate quadratic polynomial system in the same number of variables as Hydra, computed the DRL Gröbner basis, and finally converted to LEX via FGLM.
+Our FGLM results are recorded in Table 8.
+All experiments in this section have been performed on an AMD EPYC-Rome (48) CPU with 94 GB RAM.
+For the characteristic polynomial, we observed that the construction time for the multiplication matrix is negligible compared to the one for the characteristic polynomial.
+Moreover, Hydra characteristic polynomials are computed faster than for random matrices $\mathbf{A}_{0}$ and $\mathbf{A}_{1}$.
+For term order conversion, we observed that FGLM on Hydra systems requires less time than for random multivariate quadratic systems.
+We also note that in all our experiments the Hydra LEX Gröbner basis was in $\hat{x}_{2\cdot r_{\mathcal{H}}-2}$-shape position.
+Table 7: Hydra running times for construction of the multiplication matrix and its characteristic polynomial over the prime field $\mathbb{F}_{7741}$.
+The column “Determinant Random Matrix” shows the running time of the determinant from Equation 19 with uniformly random $\mathbf{A}_{0},\mathbf{A}_{1}\in\mathbb{F}_{7741}^{(2\cdot r_{\mathcal{H}}-3)\times(2\cdot r_{\mathcal{H}}-3)}$.
+Time (s)
+$r_{\mathcal{H}}$ Construction Hydra Matrix Characteristic Polynomial Hydra Matrix Determinant Random Matrix
+$5$ $<0.1$ $0.3$ $0.4$
+$6$ $0.7$ $73.9$ $82.8$
+$7$ $11.1$ $8336.7$ $11722.9$
+Table 8: Hydra FGLM running times over the prime field $\mathbb{F}_{7741}$.
+The column “Random Quadratic F4” shows the running time for computing a DRL Gröbner basis with F4 for a random multivariate quadratic polynomial system in the same number of variables as the Hydra system.
+The column “Random Quadratic FGLM” shows the running time of FGLM for the random DRL Gröbner basis.
+Time (s)
+$r_{\mathcal{H}}$ Hydra FGLM Random Quadratic F4 Random Quadratic FGLM
+$5$ $0.3$ $0.1$ $0.4$
+$6$ $20.0$ $0.8$ $26.1$
+$7$ $1265.7$ $15.6$ $1598.5$
+$8$ $78345.1$ $1123.9$ $99492.9$
 Acknowledgments
-Matthias Steiner has been supported in part by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No.
-725042) and the enCRYPTON project (grant agreement No.
-101079319).
+Matthias Steiner has been supported in part by the enCRYPTON project (grant agreement No. 101079319).